### v1

# Conditioning of Banach Space Valued Gaussian Random Variables: An Approximation Approach Based on Martingales

Abstract
In this paper we investigate the conditional distributions of two Banach space valued, jointly Gaussian random variables. These conditional distributions are again Gaussian and their means and covariances are determined by a general approximation scheme based upon a martingale idea. We then apply our general results to the case of Gaussian processes with continuous paths conditioned to partial observations of their paths.
1 Introduction
Gaussian processes for machine learning (GP4ML), see e.g. [36, 5, 24], are a standard and widespread non-parametric Bayesian technique that allows for both prediction and uncertainty quantification in a single approach. In addition, their Bayesian update rule also plays, an important role in other fields such as information-based complexity, see e.g. [44], probabilistic numerics and uncertainty quantification, see e.g. [42, 9, 20], and Bayesian methods for inverse problems, see e.g. [41].
In the simplest setup, GP4ML assume that a Gaussian process $X:=(X_{t})_{t\in T}$ is given, where $T$ is some index set. Moreover, it is assumed that both the mean function $m:T\to\mathbb{R}$ and the covariance function $k:T\times T\to\mathbb{R}$ defined by
$m(t)$ $:=\mathbb{E}X_{t}\,,$ (1)
$k(t_{1},t_{2})$ $:=\operatorname{cov}(X_{t_{1}},X_{t_{2}})$ (2)
are available. For some fixed, finite set $S:=\{s_{1},\dots,s_{n}\}\subset T$ it is further assumed that we can observe the Gaussian vector $Y:=(X_{s})_{s\in S}$, that is, we can observe the Gaussian process $X$ at the points $s_{1},\dots,s_{n}$. Given an observation $y=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ of $Y$, the “conditional Gaussian process” of $X$ given the observations $Y=y$ is then also a Gaussian process, whose mean and covariance can be computed by
$m_{X|Y=y}(t)$ $=m(t)+K_{t,S}K_{S,S}^{\dagger}(y-\mathbb{E}Y)\,,$ (3)
$k_{X|Y}(t_{1},t_{2})$ $=k(t_{1},t_{2})-K_{t_{1},S}K_{S,S}^{\dagger}K_{S,t_{2}}\,.$ (4)
Here, $K_{t_{1},S}$, $K_{S,S}$, and $K_{S,t_{2}}$ denote covariance matrices with respect to the indicated indices and $K_{S,S}^{\dagger}$ is the Moore-Penrose inverse of the $n\times n$-matrix $K_{S,S}$. Here we note that in the GP4ML literature, it is often assumed that we have noisy observations of the form $Y+\varepsilon$, where $\varepsilon$ is a centered Gaussian vector of i.i.d. entries that is independent of $Y$. In this case, the formulas slightly change, see the discussion following (29).
Usually, the Gaussian process $X$ is viewed as a mechanism to generate random functions, e.g. random continuous functions on a compact metric space $T$. In this case, the “conditional Gaussian process” of $X$ given the observations $Y=y$ is also viewed as such a mechanism. To make this mathematically rigorous, one needs to view $X$ as a $C(T)$-valued Gaussian random variable, where $C(T)$ denotes the Banach space of continuous $\mathbb{R}$-valued functions on $T$, see Sections 2 and 4 for these and the following notions. Moreover, the above $Y$ is an $\mathbb{R}^{n}$-valued Gaussian random variable such that $X$ and $Y$ are jointly Gaussian. Given their joint distribution $\mathrm{P}_{(X,Y)}$ on $C(T)\times\mathbb{R}^{n}$, one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$, since $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is the conditional distribution of $X$ given the observation $Y=y$. Note that by definition, each $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $C(T)$ and therefore we can also ensure that the “conditional Gaussian process” again produces continuous functions.
The references mentioned at the beginning clearly indicate that for the conditioning problem $C(T)$ is by no means the only Banach space one is interested in. Therefore, let us now consider the general case. To this end, let $E$ and $F$ be Banach spaces, which due to technical simplicity are assumed to be separable. Moreover, let $X$ be an $E$-valued random variable and $Y$ be an $F$-valued random variable such that $X$ and $Y$ are jointly Gaussian. In the abstract setting one then seeks a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$. Again, $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a distribution on $E$, which equals the conditional distribution of $X$ given the observation $Y=y$. Compared to the $C(T)$-case described earlier, this generalization opens several additional possibilities. First, of course, we can consider Banach spaces of functions $E$ other than $C(T)$, e.g. spaces of differentiable functions. Second, we no longer need to assume that we can observe the random functions on some finite number of observational points. Instead, we could assume, e.g., that we observe derivatives of the random functions at some observational points, cf. [35]. More generally, $Y$ can, for example, be composed by any other finite set of bounded linear functionals as considered in e.g. [45], or even by an infinite set of such functions. In particular, for the $C(T)$-example above, the abstract setting makes it possible to consider arbitrary subsets $S\subset T$ for the observational points, which is important for PDE-related applications, see e.g. [35]. To abstract from this, let us assume that $Y$ is given by
$Y=\mathsf{S}\circ X\,,$ (5)
where $\mathsf{S}:E\to F$ is a known bounded and linear operator. Clearly, in the $C(T)$-example, $\mathsf{S}$ would be the restriction operator $C(T)\to C(S)$ given by $f\mapsto f_{|S}$, and hence (5) indeed generalizes this example.
Although working with infinite dimensional $Y$ has clear conceptual advantages, there is also a drawback: Indeed, in many situations, it is practically impossible to fully observe or process infinite dimensional $Y$. In these cases, one often uses a sequence $(Y_{n})$ of observable, finite dimensional Gaussian random variables, for which the condition step is feasible. For example, in the $C(T)$-case with infinite $S\subset T$, one could choose an increasing sequence $(S_{n})$ of finite subsets $S_{n}\subset S$ whose union is dense in $S$ and set $Y_{n}:=(X_{s})_{s\in S_{n}}$. The hope of this approach is, of course, that $\mathrm{P}_{X|Y_{n}}$ provides a convergent approximation of $\mathrm{P}_{X|Y}$, e.g. in terms of the means and covariances, or as distributions.
Both the fully infinite dimensional conditioning and its possible approximations by finite dimensional observations have various applications as described in e.g. [44, 9, 20, 35]. Here, we only mention Bayesian quadrature, Bayesian optimization, and Bayesian solvers for differential equations. In addition, they can help to better understand GP4ML and their limiting behavior for $n\to\infty$. For this reason, a better understanding of the regular conditional probabilities of $X$ given $Y$ for jointly Gaussian, Banach space valued random variables $X$ and $Y$ as well as of possible approximations is of high practical interest.
Let us therefore have a look what is known so far. To begin with, in the case of finite dimensional observations, that is $F=\mathbb{R}^{n}$, [6, Corollary 3.10.3] shows that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. However, no formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are provided. Moreover, the observational model (5) for centered $X$ and with arbitrary $E$ and $F$ has been investigated in [43], where the authors show that the regular conditional probabilities are Gaussian measures, see their Theorem 3.11. Moreover, in the case $F=\mathbb{R}^{n}$, their Corollary 3.13 present explicit formulas for the conditional mean and covariance with the help of a so-called representing sequence of the covariance operator $\operatorname{cov}Y$. In addition, a closer look at their proofs reveals, that analogous formulas hold true for general, separable $F$. The latter formulas have been made explicit and generalized to non-centered $X$ in [45]. Moreover, both papers establish the linearity of the conditional mean mapping, and [45] also describes formulas for sequential updates and the case of $C(T)$-valued $X$. In particular, for $E=C(T)$ and $F=\mathbb{R}^{n}$, the update rules (3) and (4) are generalized to arbitrary, bounded linear observational operators $\mathsf{S}:C(T)\to\mathbb{R}^{n}$. Unfortunately, however, the mean and covariance formulas for infinite dimensional $F$ become less handy, as they require to have a representing sequence for $\operatorname{cov}Y$. Although such sequences as well as constructions for them exist, see e.g. [43, Lemma 3.5] or [40, Lemma 8.2.3], it seems fair to say that they are rarely at hand in situations in which $F$ is not a Hilbert space. As a consequence, it seems fair to say that the derived formulas have some limited practical value.
The case of $C(T)$-valued $X$ is also considered in [29]. There, $T=[0,t^{*}]$ is an interval and $Y$ is given by a partial observation of $X$ up to some earlier time $s\in[0,t^{*}]$, that is $Y:=X_{|[0,s]}$. Under an additional assumption, the author then proves that the conditional distribution $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure on $C(T)$ and that the map $y\mapsto\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is even continuous if the set of probability measures on $C(T)$ is equipped with a metric describing the weak convergence of probability measures. In addition, it is shown that this additional assumption is also necessary for the existence of such continuous conditional probabilities. In fact it is worth to mention that both results are actually first shown for the general observational model (5) and then applied to the $C(T)$-case, but the motivation of [29] clearly focuses on the $C(T)$-case.
A different dependence of the conditional distributions is considered in [1]. There, the general case of joint Gaussian random variables $X$ and $Y$ is considered and, in addition, it is assumed that their joint distribution $\mathrm{P}_{(X,Y),\alpha}$ depends measurably on a parameter $\alpha$. The main result of [1] then shows that there exists a (regular) conditional probability $\mathrm{P}_{(X|Y),\alpha}$ such that $(y,\alpha)\mapsto\mathrm{P}_{(X|Y),\alpha}(\,\cdot\,|y)$ is measurable if the set of probability measures is equipped with the Borel $\sigma$-algebra generated by the topology of weak convergence of probability measures.
If both $E$ and $F$ are Hilbert spaces, the conditioning problem has also been investigated by different techniques. For example, in [33] formulas for the mean and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ are established in the full infinite dimensional setting in terms of both a) so-called shorted operators and b) suitable sequences of oblique projections. Moreover, [26] shows, besides many other results, that some of the formulas derived in [33] can be significantly simplified. Independently, [31, 17] established simple formulas for the the mean and covariance, which essentially resemble (3) and (4) in an abstract sense, under the additional assumption that $\operatorname{cov}Y$ is injective.
The approximation in the $C(T)$-case with $S=T$ has also been considered in the literature. For example, convergence rates for the conditional mean functions, that is for $m_{X|Y_{n}=y_{n}}\to m_{X|Y=y}$, have been derived in [47] for specific Gaussian processes and the results of [15] can also be used to obtain such rates for random observational points and additional, vanishing observational noise. We refer to [24, Section 5.1] for a detailed discussion, where we note that [15] provides the better results compared to the source [39] mentioned in [24], since a) a clipping step for the mean is avoided and b) also stronger norms for the convergence are considered. In addition, convergence rates for Sobolev type covariance functions can be derived using the fill-distance and results from kernel interpolation [49]. For a detailed discussion, we again refer to [24] and the references mentioned therein. More recently, [28] established a uniform convergence result for the conditional mean and covariance functions under rather general assumptions on the underlying space $T$ and the prior kernel $k$. In addition, this paper provides a small list of earlier, less general convergence results in the $C(T)$-case. Besides the $C(T)$-case, however, we are not aware of any convergence results. We view this gap as substantial since conditioning on observations that are not direct partial observations of the paths of $X$ are interesting for various practical reasons, see e.g. the discussions in [45, 35].
In this work, we address the general, infinite dimensional conditioning and the described approximation problem simultaneously. Namely, we show that given a sufficiently rich sequence $(y_{n}^{\prime})\subset F^{\prime}$ in the dual of $F$, see Definition 3.2 and the examples in Section 4, and considering the $\mathbb{R}^{n}$-valued Gaussian random variables $Y_{n}:=(\langle y_{1}^{\prime},Y\rangle_{F^{\prime},F},\dots,\langle y_{n}^{\prime},Y\rangle_{F^{\prime},F}$, we have the weak convergence
$\mathrm{P}_{X|Y_{n}}(\,\cdot\,|y_{n})$ $\to\mathrm{P}_{X|Y}(\,\cdot\,|y)$
for almost all observations $y$ of $Y$, where $y_{n}:=(\langle y_{1}^{\prime},y\rangle_{F^{\prime},F},\dots,\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}$. In addition, we show that the means and covariance operators also converge, see Theorem 3.4 for these three statements. Moreover, we prove that the means and covariances of $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|y_{n})$ can be computed analogously to the GP4ML case (3) and (4), see Theorems 3.1 and 3.4. Furthermore, we show that the set of $y\in F$, for which the means converge, is a measurable, affine linear subspace of $F$ with full measure and this can be used to construct a particular version of $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, see Theorem 3.6. As the title of this paper indicates, the major work horse in our proofs is a Banach space-valued martingale approach, which is coupled with the relationship between regular conditional probabilities and conditional expectations in the Banach space case, see Theorem 5.10. In this respect we note that martingale arguments have already been used by e.g. [31, 33] in the Hilbert space setting, but as far as we know, not in the general Banach space setting. In addition, the resulting martingale approximations in [31, 33] are restricted to rather specific filtrations, whereas our results only require a natural richness assumption on the filtration. Finally, we apply our general theory to the $C(T)$-valued case, where we establish uniformly convergent limit versions of the update rules (3) and (4) for general $S\subset T$.
The rest of this paper is organized as follows: In Section 2 we recall all notions and notations from probability theory and functional analysis that are necessary to understand our main results. These main results are then presented in Section 3 and the $C(T)$-valued case is subsequently investigated in Section 4. Sections 5 and 6 contain the proofs of the results of Section 3, respectively Section 4. Finally, we have attached four detailed appendices on A) Banach space valued integration, B) covariance operators, C) characteristic functions for Banach space valued random variables, and D) regular conditional probabilities and conditional expectations. Most material in these appendices is simply collected and unified from the existing literature, but a couple of mostly simple results are also added since we could not find them in the literature.
2 Preliminaries
Throughout this paper, $E$ and $F$ denote Banach spaces. Moreover, we write $B_{E}$ for the closed unit ball of $E$ and $E^{\prime}$ for the dual of $E$. In addition, $\iota_{E}:E\to E^{\prime\prime}$ denotes the canonical embedding of $E$ into its bi-dual $E^{\prime\prime}$. Furthermore, we write $\langle\,\cdot\,,\,\cdot\,\rangle_{E^{\prime},E}:E^{\prime}\times E\to\mathbb{R}$ for the standard dual pairing defined by
$\langle x^{\prime},x\rangle_{E^{\prime},E}:=x^{\prime}(x)\,,\qquad\qquad x^{\prime}\in E^{\prime},x\in E.$
In addition, recall that the $w^{*}$-topology on $E^{\prime}$ is the smallest topology, such that all maps $\langle\,\cdot\,,x\rangle_{E^{\prime},E}:E^{\prime}\to\mathbb{R}$ are continuous. The $w^{*}$-topology has some highly useful properties, which are summarized in the following result, whose first two assertions can be found in in [32, Theorem 2.6.18 and Theorem 2.6.23] or [11, Theorem V.3.1 and Theorem V.5.1]. The last assertion follows from the fact that compact metric spaces are separable.
Theorem 2.1.
Let $E$ be a separable Banach space. Then $B_{E^{\prime}}$ is $w^{*}$-compact and the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable. Finally, there exists a countable $\mathfrak{D}\subset B_{E^{\prime}}$ that is $w^{*}$-dense in $B_{E^{\prime}}$.
Note that for the proofs of our main results we need the existence of such sets $\mathfrak{D}$, and another key result, namely Proposition 4.1, also need such sets in its formulation. In addition, the fact that the $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable is often used in the the proofs as it allows to deal with sequences rather than nets and this in turn opens the door for limit theorems of integrals.
In this work, we need two $\sigma$-algebras on $E$: The first one is the usual Borel $\sigma$-algebra $\mathfrak{B}(E)$ that is generated by the norm topology on $E$ and the second one is $\sigma(E^{\prime})$, that is, the smallest $\sigma$-algebra on $E$, for which all bounded, linear $x^{\prime}:E\to\mathbb{R}$ are measurable, where, as usual, $\mathbb{R}$ is equipped with $\mathfrak{B}:=\mathfrak{B}(\mathbb{R})$. Similarly, we write $\mathfrak{B}^{d}:=\mathfrak{B}(\mathbb{R}^{d})$. Throughout this work, we assume that $\mathbb{R}^{d}$ is equipped with $\mathfrak{B}^{d}$.
In general, we have $\sigma(E^{\prime})\subset\mathfrak{B}(E)$ and the following theorem, which is a consequence of e.g. [10, Theorem E.9] and which can be also found in [21, Proposition 1.1.1], shows that for separable $E$ both $\sigma$-algebras coincide.
Theorem 2.2.
Let $E$ be a separable Banach space. Then we have $\sigma(E^{\prime})=\mathfrak{B}(E)$.
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. Then $X$ is called weakly measurable if it is $({\cal A},\sigma(E^{\prime}))$-measurable. By a standard measurability argument, see e.g. [27, Corollary 1.82], we see that this is equivalent to the $({\cal A},\mathfrak{B})$-measurability of $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ for all $x^{\prime}\in E^{\prime}$. Moreover, $X$ is called strongly measurable if it is $({\cal A},\mathfrak{B}(E))$-measurable and $f(\Omega)$ is separable. If $E$ is separable, then one can quickly check with Theorem 2.2 that weak and strong measurability are equivalent. Finally, if $(\Omega,{\cal A},\mathrm{P})$ is a probability space, then a strongly measurable $X:\Omega\to E$ is called a random variable, and in this case its image measure $\mathrm{P}_{X}$ on $\mathfrak{B}(E)$ is called the distribution of $X$. Of course, if $E$ is separable, every weakly measurable $X:\Omega\to E$ is a random variable, and for this reason we sometimes simply speak of measurable $X$.
We also need to recall the notion of weak convergence of probability measures. To this end, let $E$ be a Banach space, $(\mathrm{P}_{n})$ be a sequence of probability measures on $\mathfrak{B}(E)$, and $\mathrm{P}$ be another probability measure on $\mathfrak{B}(E)$. Then we say that $(\mathrm{P}_{n})$ converges weakly to $\mathrm{P}$, if for all bounded and continuous $f:E\to\mathbb{R}$ we have
$\int_{E}f\,\mathrm{d}\mathrm{P}_{n}\to\int_{E}f\,\mathrm{d}\mathrm{P}\,.$ (6)
It turns out, that in this case the limit $\mathrm{P}$ is unique, see e.g. [46, Chapter I.3.7]. Of course, for $E=\mathbb{R}^{d}$, Lévy’s continuity theorem shows that weak convergence of probability measures can be characterized by the pointwise convergence of the corresponding characteristic functions, see also Theorem C.3. A generalization to the case of separable Banach spaces $E$ is recalled in Theorem C.6. Finally, like in the finite dimensional case, we say that a sequence $(X_{n})$ of $E$-valued random variables converge to an $E$-valued random variable $X$ in distribution, if the distributions of $(X_{n})$ weakly converge to the distribution of $X$.
Let us now fix a probability space $(\Omega,{\cal A},\mathrm{P})$, a Banach space $E$, and $p\in[1,\infty)$. We write
${{\cal L}_{p}(\mathrm{P},E)}:=\biggl{\{}X:\Omega\to E\,\Bigl{|}\,X\mbox{ is strongly measurable and }\int_{\Omega}\|X(\omega)\|_{E}^{p}\,\mathrm{d}\mathrm{P}(\omega)<\infty\biggr{\}}\,.$
In addition, each $X\in{{\cal L}_{1}(\mathrm{P},E)}$ is called Bochner integrable and for these $X$ one can define the expectation by the $E$-valued Bochner integral
$\mathbb{E}X:=\int_{\Omega}X\,\mathrm{d}\mathrm{P}\,.$
We refer to Appendix A, which collects some information on Bochner integrals. Moreover, for $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},E)}$, where $F$ is another Banach space, the cross covariance of $X$ and $Y$ is the bounded linear operator $\operatorname{cov}(X,Y):F^{\prime}\to E$ defined by the Bochner integrals
$\operatorname{cov}(X,Y)(y^{\prime})=\int_{\Omega}\langle y^{\prime},Y-\mathbb{E}Y\rangle_{F^{\prime},F}\,(X-\mathbb{E}X)\,\mathrm{d}\mathrm{P}\,,\qquad\qquad y^{\prime}\in F^{\prime}.$
Recall that $\operatorname{cov}(X,Y):F^{\prime}\to E$ is a bounded linear operator, which is nuclear and thus compact, see e.g. [46, Theorem 2.5 in Chapter III.2] and Lemma B.7 for an alternative argument. In particular, the covariance of $X$, which is defined by
$\operatorname{cov}(X):=\operatorname{cov}(X,X)\,,$
is a nuclear operator $\operatorname{cov}(X):E^{\prime}\to E$. For more information on (cross) covariances we refer to Appendix B. Finally, in the case $E=\mathbb{R}^{m}$ and $F=\mathbb{R}^{n}$ the cross covariance operator $\operatorname{cov}(X,Y):(\mathbb{R}^{n})^{\prime}\to\mathbb{R}^{m}$ can be described by the $m\times n$-covariance matrix
$K_{X,Y}:=\bigl{(}\operatorname{cov}(X_{i},Y_{j})\bigr{)}_{i,j}\,,$ (7)
where $X_{i}$ and $Y_{j}$ denote the $i$th, respectively $j$th, component function of $X$ and $Y$. Note that in this description, we formally write elements of $(\mathbb{R}^{n})^{\prime}$ as row vectors and identify $(\mathbb{R}^{n})^{\prime}$ by $\mathbb{R}^{n}$ via the canonical isomorphism given by the transposition, that is $\,\cdot\,^{\mathsf{T}}:(\mathbb{R}^{n})^{\prime}\to\mathbb{R}^{n}$. In other words $\operatorname{cov}(X,Y)(y^{\prime})=K_{X,Y}\cdot(y^{\prime})^{\mathsf{T}}$ for all $y^{\prime}\in(\mathbb{R}^{n})^{\prime}$.
Recall that a probability measure $\mathrm{P}$ on $(\mathbb{R},\mathfrak{B})$ is a Gaussian measure, if there are $\mu\in\mathbb{R}$ and $\sigma\in[0,\infty)$ such that the characteristic function of $\mathrm{P}$ is given by
$\varphi_{\mathrm{P}}(t)=\mathrm{e}^{\mathrm{i}t\mu}\cdot\exp\biggl{(}-\frac{\sigma^{2}t^{2}}{2}\biggr{)}\,,\qquad\qquad t\in\mathbb{R}.$
In the case $\sigma>0$, we thus have $\mathrm{P}={\cal N}(\mu,\sigma^{2})$, i.e. $\mathrm{P}$ is a normal distribution with expectation $\mu$ and variance $\sigma^{2}$. If $\sigma=0$, the probability measure $\mathrm{P}$ has all its mass on $\mu$ and we write ${\cal N}(\mu,0):=\mathrm{P}$. Finally, if $(\Omega,{\cal A},\mathrm{P})$ is a probability space and $X:\Omega\to\mathbb{R}$ is measurable, then $X$ is said to be an $\mathbb{R}$-valued Gaussian random variable if its distribution $\mathrm{P}_{X}$ is a Gaussian measure on $(\mathbb{R},\mathfrak{B})$.
With these preparations we can now define $E$-valued Gaussian random variables and Gaussian measures on $E$. Since we will need the separability of $E$ at various places, we restrict the following definition to separable Banach spaces.
Definition 2.3.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space and $E$ be a separable Banach space. Then $X:\Omega\to E$ is a Gaussian random variable, if it is weakly measurable and $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ is a Gaussian random variable for all $x^{\prime}\in E^{\prime}$.
Moreover, a probability measure $\mathrm{Q}$ on $\sigma(E^{\prime})=\mathfrak{B}(E)$ is a Gaussian measure, if the image measure $\mathrm{Q}_{x^{\prime}}$ on $(\mathbb{R},\mathfrak{B})$ is Gaussian measure for all $x^{\prime}\in E^{\prime}$.
It turns out, that it suffices to verify the Gaussianity for a $\mathfrak{D}\subset B_{E^{\prime}}$ that is $w^{*}$-dense in $B_{E^{\prime}}$, see Theorem 5.1. In addition, it directly follows from the definition that $\mathrm{Q}$ is a Gaussian measure on $\sigma(E^{\prime})$, if and only if the identity map $\operatorname{id}_{E}:E\to E$ defined on the probability space $(E,\sigma(E^{\prime}),\mathrm{Q})$ is a Gaussian random variable. Finally, every affine linear transformation of a Gaussian random variable is a Gaussian random variable. To be more precise, if $X:\Omega\to E$ is a Gaussian random variable, $T:E\to F$ is a bounded linear operator mapping into a separable Banach space $F$, and $y\in F$, then $y+T\circ X$ is Gaussian random variable.
Recall that Fernique’s theorem, see e.g. [30, Corollary 3.2], shows that Gaussian random variables $X:\Omega\to E$ on some probability space $(\Omega,{\cal A},\mathrm{P})$ satisfy
$\int_{\Omega}\|X(\omega)\|_{E}^{p}\,\mathrm{d}\mathrm{P}(\omega)<\infty$ (8)
for all $p\in[1,\infty)$. Consequently, for every Gaussian random variable $X$ its expectation $\mathbb{E}X$ and covariance $\operatorname{cov}(X)$ are defined. Moreover, if $\mathrm{Q}$ is a Gaussian measure on $\sigma(E^{\prime})$, then we define its expectation and covariance by those of the Gaussian random variable $\operatorname{id}_{E}$ mentioned above, that is
$\mu_{\mathrm{Q}}$ $:=\mathbb{E}_{\mathrm{Q}}\operatorname{id}_{E}=\int_{\mathbb{E}}x\,\mathrm{d}\mathrm{Q}(x)\,,$
$\operatorname{cov}(\mathrm{Q})(x^{\prime})$ $:=\operatorname{cov}_{\mathrm{Q}}(\operatorname{id}_{E})(x^{\prime})=\int_{\Omega}\langle x^{\prime},x-\mu_{\mathrm{Q}}\rangle_{E^{\prime},E}\cdot(x-\mu_{\mathrm{Q}})\,\mathrm{d}\mathrm{Q}(x)\,,\qquad\qquad x^{\prime}\in E^{\prime}.$
Like in the finite dimensional case, the expectation and covariance uniquely determine a Gaussian measure. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are Gaussian measures on $\sigma(E^{\prime})$ with $\mu_{\mathrm{P}}=\mu_{\mathrm{Q}}$ and $\operatorname{cov}(\mathrm{P})=\operatorname{cov}(\mathrm{Q})$, then we have $\mathrm{P}=\mathrm{Q}$. This can be quickly verified with the help of characteristic functions, e.g. by combining [46, Theorem 2.2 in Chapter IV.2.1] with [46, Proposition 2.8 in Chapter IV.2.4], where the latter is restated in a simplified form in Proposition C.5. For this reason, we write
${\cal N}\bigl{(}\mu_{\mathrm{Q}},\operatorname{cov}(\mathrm{Q})\bigr{)}:=\mathrm{Q}$ (9)
for Gaussian measures $\mathrm{Q}$. More information on characteristic functions of Gaussian measures and random variables are collected in Appendix C. Finally, we call
${\cal G}(X):=\overline{\bigl{\{}\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}:x^{\prime}\in E^{\prime}\bigr{\}}}^{{L_{2}(\mathrm{P})}}$
the Gaussian Hilbert space associated to $X$, where, as usual, we identify the functions $\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}$ with their $\mathrm{P}$-equivalence classes. It turns out that ${\cal G}(X)$ is indeed a Gaussian Hilbert space in the sense of [23]. For this and additional properties of ${\cal G}(X)$ we refer to Lemma 5.4.
In this work we often deal with $E\times F$-valued Gaussian random variables. To describe such random variables, recall that the map $E^{\prime}\times F^{\prime}\to(E\times F)^{\prime}$ defined by $(x^{\prime},y^{\prime})(x,y):=x^{\prime}(x)+y^{\prime}(y)$ for all $x^{\prime}\in E^{\prime}$, $y^{\prime}\in F^{\prime}$, $x\in E$, and $y\in F$ is an isomorphism. This leads to the following definition.
Definition 2.4.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be a separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be random variables. Then we say that $X$ and $Y$ are jointly Gaussian random variables, if $(X,Y):\Omega\to E\times F$ is a Gaussian random variable, that is,
$\langle x^{\prime},X\rangle_{E^{\prime},E}+\langle y^{\prime},Y\rangle_{F^{\prime},F}:\Omega\to\mathbb{R}$ (10)
is a Gaussian random variable for all $x^{\prime}\in E^{\prime}$ and $y^{\prime}\in F^{\prime}$.
Obviously, if $X$ and $Y$ are jointly Gaussian random variables, then both $X$ and $Y$ are Gaussian random variables. Moreover, if $X$ is an $E$-valued Gaussian random variable, then $(X,X)$ is an $E\times E$-valued Gaussian random variable.
Given jointly Gaussian random variables $X:\Omega\to E$ and $Y:\Omega\to F$, the major goal of this work is to investigate the conditional distribution of $X$ given an observation $Y=y$. To make this mathematically rigorous, the following definition recalls the concept of regular conditional probabilities.
Definition 2.5.
Let $(T,{\cal B})$ and $(U,{\cal A})$ be measurable spaces and $\mathrm{P}$ be a probability measure on $(T\times U,{\cal B}\otimes{\cal A})$. Then we say that a map $\mathrm{P}(\,\cdot\,|\,\cdot\,):{\cal B}\times U\to[0,1]$ is a regular conditional probability of $\mathrm{P}$ given $U$, if the following three conditions are satisfied:
  1. i)
    For all $u\in U$ the map $\mathrm{P}(\,\cdot\,|u):{\cal B}\to[0,1]$ is a probability measure.
  2. ii)
    For all $B\in{\cal B}$ the map $\mathrm{P}(B|\,\cdot\,):U\to[0,1]$ is ${\cal A}$-measurable.
  3. iii)
    For all $B\in{\cal B}$ and $A\in{\cal A}$ we have
    $\mathrm{P}(B\times A)=\int_{U}\boldsymbol{1}_{A}(u)\mathrm{P}(B|u)\,\mathrm{d}\mathrm{P}_{U}(u)\,.$ (11)
    Here, $\mathrm{P}_{U}$ denotes the marginal distribution of $\mathrm{P}$ on $U$, that is $\mathrm{P}_{U}:=\mathrm{P}_{\pi_{U}}$, where $\mathrm{P}_{\pi_{U}}$ is the image measure of $\mathrm{P}$ under the coordinate projection $\pi_{U}:T\times U\to U$.
Note that if $E$ and $F$ are separable Banach space and $\mathrm{Q}$ is a probability measure on $\mathfrak{B}(E)\otimes\mathfrak{B}(F)$, then there always exists a regular conditional probability $\mathrm{Q}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\otimes F\to[0,1]$ of $\mathrm{Q}$ given $F$. In addition, this regular conditional probability is always $\mathrm{P}_{F}$-almost surely unique. These statements directly follow from a more general and well-known result on the existence and uniqueness of regular conditional probabilities, which is recalled in Theorem D.1.
Let us now apply these notion and results to jointly Gaussian random variables $X:\Omega\to E$ and $Y:\Omega\to F$ that are defined on some probability space $(\Omega,{\cal A},\mathrm{P})$. To this end, we consider their joint distribution $\mathrm{Q}:=\mathrm{P}_{(X,Y)}$ on $\mathfrak{B}(E)\otimes\mathfrak{B}(F)$. A quick calculation yields $\mathrm{Q}_{F}=\mathrm{P}_{Y}$, and our discussion above thus shows that there exists a $\mathrm{P}_{Y}$-almost surely unique regular conditional probability
$\mathrm{Q}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times F\to[0,1]$
of $\mathrm{Q}$ given $F$. We write $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):=\mathrm{Q}(\,\cdot\,|\,\cdot\,)$ and speak of the regular conditional probability of $X$ given $Y$. Note that this is justified as $\mathrm{P}(B|y)$ describes the probability of $X\in B$ under the condition that we have observed $y=Y(\omega)$. Consequently, our goal is to describe $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ modulo $\mathrm{P}_{Y}$-sets. This task will be accomplished by
  1. i)
    showing that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure, and
  2. ii)
    determining the expectation and covariance of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$.
As discussed around (9) these two steps completely describe $\mathrm{P}_{X|Y}(\,\cdot\,|y)$.
In the finite dimensional case, the solution of this task is folklore, see e.g. [36, Appendix A.2], [5, Chapter 2.3], [19, Lemma 4.3], [41, Theorem 6.20], and also [35, Section 4]. To formulate the solution in the most general setting, let us recall that a $n\times m$-Matrix $B$ is said to be a Moore-Penrose pseudo-inverse of a given $m\times n$-matrix $A$, if the following four equations hold true:
$ABA$ $=A$ and $BAB$ $=B\,,$ (12)
${(AB)}^{\mathsf{T}}$ $=AB$ and ${(BA)}^{\mathsf{T}}$ $=BA\,.$ (13)
It can be shown that for every $m\times n$-matrix $A$ there exists exactly one Moore-Penrose pseudo-inverse, which we denote by $A^{\dagger}$, see e.g. [34, Theorem 1]. Similarly, if we have a linear map $A:\mathbb{R}^{n}\to\mathbb{R}^{m}$, then a linear map $B:\mathbb{R}^{m}\to\mathbb{R}^{n}$ is a Moore-Penrose pseudo-inverse of $A$, if the four equations above are satisfied. The uniqueness of such a pseudo-inverse can be shown analogously to the matrix case and the existence follows from the matrix case by using the matrices representing $A$ and $B$. We again write $A^{\dagger}$ for the pseudo-inverse.
With these preparations, the description of the regular conditional probability $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ in the finite dimensional case reads as follows.
Theorem 2.6.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space and $X:\Omega\to\mathbb{R}^{m}$ and $Y:\Omega\to\mathbb{R}^{n}$ be jointly Gaussian random variables. Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}^{m}\times\mathbb{R}^{n}\to[0,1]$ of the regular conditional probability of $X$ given $Y$ there exists an $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$ such that for all $y\in\mathbb{R}^{n}\setminus N$ we have
$\mathrm{P}_{X|Y}(\,\cdot\,|y)={\cal N}\bigl{(}\mu_{X|Y}(y),K_{X|Y}\bigr{)}\,,$
where
$\mu_{X|Y}(y)$ $:=\mathbb{E}X+K_{X,Y}K_{Y,Y}^{\dagger}(y-\mathbb{E}Y)\,,$
$K_{X|Y}$ $:=K_{X,X}-K_{X,Y}K_{Y,Y}^{\dagger}K_{Y,X}\,.$
Finally, we also need the concept of conditional expectations, which, at least for $\mathbb{R}$-valued random variables, is probably better known than regular conditional probabilities. Its generalization to $E$-valued random variables is straightforward. Namely, if $E$ is a Banach space, $(\Omega,{\cal A},\mathrm{P})$ is a probability space, ${\cal C}\subset{\cal A}$ is a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$, then a $\mathrm{P}$-integrable and strongly ${\cal C}$-measurable $Z:\Omega\to E$ is called a conditional expectation of $X$ given ${\cal C}$, if
$\int_{\Omega}\boldsymbol{1}_{C}Z\,\mathrm{d}\mathrm{P}=\int_{\Omega}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}\,,\qquad\qquad C\in{\cal C}.$
Like in the $\mathbb{R}$-valued case, the conditional expectation of $X$ given ${\cal C}$ is $\mathrm{P}$-almost surely unique, and hence we write $\mathbb{E}(X|{\cal C}):=Z$ and say that $Z$ is a version of $\mathbb{E}(X|{\cal C})$. Moreover, if ${\cal C}$ is generated by a random variable $Y$, that is ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|\sigma(Y))=\mathbb{E}(X|{\cal C})$. Banach space valued conditional expectations enjoy a lot of the properties of $\mathbb{R}$-valued conditional expectations, see Appendix D. In addition, Appendix D also contains a detailed description of the relationship between regular conditional probabilities and conditional expectations in the Banach space setting.
3 Main Results
In this section we present the main results of this work that investigate the regular conditional probabilities arising from conditioning one Banach space valued Gaussian random variable with respect to another one.
We begin by presenting a theorem that generalizes and extends the finite dimensional Theorem 2.6 to the case in which the random variable $X$ is Banach space valued. Of course, for such $X$ we can no longer work with matrices, but apart from this difference, the first part of the following theorem is an almost literal copy of Theorem 2.6.
Theorem 3.1.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ be a separable Banach space, and $X:\Omega\to E$ and $Y:\Omega\to\mathbb{R}^{n}$ be be jointly Gaussian random variables. Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ of the regular conditional probability of $X$ given $Y$ there exist an $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$ such that for all $y\in\mathbb{R}^{n}\setminus N$ we have
$\mathrm{P}_{X|Y}(\,\cdot\,|y)={\cal N}\bigl{(}\mu_{X|Y}(y),C_{X|Y}\bigr{)}\,,$
where
$\mu_{X|Y}(y)$ $=\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\,,$
$C_{X|Y}$ $=\operatorname{cov}(X)-\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)\,.$
In addition, the random variable $Z:\Omega\to E$ defined by
$Z:=\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)$
is both a version of $\mathbb{E}(X|Y)$ and a Gaussian random variable with $\mathbb{E}Z=\mathbb{E}X$ and
$\operatorname{cov}(Z)$ $=\operatorname{cov}(X)-C_{X|Y}\,,$
${\cal G}(Z)$ $\subset{\cal G}(Y)\,.$
As mentioned in the introduction, [6, Corollary 3.10.3] has already shown that there is a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$, such that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y$. Later, in Theorem 3.5 we will recover such a version, too. Moreover, for more specific cases, formulas for the mean and covariance have also been derived by other authors as laid out in the introduction. However, we could not find such formulas in the full generality presented in Theorem 3.1. Since these formulas play a crucial role for our approximation scheme in the case of infinite dimensional $Y$, we thus decided to present Theorem 3.1 as a main result.
Our next goal is to generalize Theorem 3.1 to infinite dimensional $Y$. Here we recall that we cannot expect a literal generalization. To illustrate this, let us assume for simplicity that $H:=F$ is an infinite dimensional Hilbert space. Using the canonical Riesz-Fréchet-identification of $H^{\prime}$ with $H$, we can then view $\operatorname{cov}(Y)$ as a compact linear operator $\operatorname{cov}(Y):H\to H$. Now, [18, Theorem 2.4] shows that a bounded linear operator $A:H\to H$ has a Moore-Penrose inverse if and only if $\operatorname{ran}A$ is closed. However, it is well-known that compact operators have a closed range, if and only if their range is actually finite dimensional, see e.g. [32, Proposition 3.4.6]. Consequently, the Moore-Penrose inverse $(\operatorname{cov}Y)^{\dagger}$ exists if and only if $\operatorname{ran}\operatorname{cov}(Y)$ is finite dimensional. The latter is only possible, if the range of $Y$ is almost surely contained in a finite-dimensional subspace of $F$, see e.g. [29, Theorem 2.1] or [42, Theorem 2.43].
In the following, we will resolve this issue by considering a sequence $(Y_{n})$ of suitably chosen finite dimensional projections of $Y$. For each $Y_{n}$ we can then apply Theorem 3.1, and it remains to show that the resulting sequence of regular conditional probabilities $\mathrm{P}_{X|Y_{n}}$ converges to the regular conditional probability $\mathrm{P}_{X|Y}$. The following definition introduces the type of projections we need for this approach.
Definition 3.2.
Let $F$ be a separable Banach space and $(A_{n})$ be a sequence of bounded linear operators $A_{n}:F\to\mathbb{R}^{n}$. Then we say that $(A_{n})$ is a filtering sequence for $F$, if we have $\sigma(A_{n})\subset\sigma(A_{n+1})$ for all $n\geq 1$ and
$\sigma(F^{\prime})=\sigma\bigl{(}A_{n}:n\geq 1\bigr{)}\,.$
The measure-theoretic description of filtering sequences is useful for establishing our general result on conditioning jointly Gaussian random variables, yet it is somewhat cumbersome to verify. The following proposition makes this verification substantially easier.
Proposition 3.3.
Let $F$ be a separable Banach space and $(y_{i}^{\prime})\subset F^{\prime}$ be a sequence with
$\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}=F^{\prime}\,,$ (14)
where on the left hand side, the $w^{*}$-sequential closure is considered. For $n\geq 1$ we define $A_{n}:F\to\mathbb{R}^{n}$ by
$A_{n}(y):=\bigl{(}\langle y_{1}^{\prime},y\rangle_{F^{\prime},F},\dots,\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}\bigr{)}\,,\qquad\qquad y\in F.$ (15)
Then $(A_{n})$ is a filtering sequence for $F$.
Using Proposition 3.3 it is easy to construct filtering sequences, both in an abstract setting as well as in concrete examples. This will be discussed in detail in Section 4.
With the help of filtering sequences we can now present our main result on conditioning one Gaussian random variable with respect to another one.
Theorem 3.4.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be jointly Gaussian random variables. Moreover, let $(A_{n})$ be a filtering sequence for $F$ and
$Y_{n}:=A_{n}\circ Y\,,\qquad\qquad n\geq 1.$
Then for all versions $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ and $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times F\to[0,1]$ of the regular conditional probabilities of $X$ given $Y_{n}$, respectively $Y$, there exists an $N\in\sigma(F^{\prime})=\mathfrak{B}(F)$ with $\mathrm{P}_{Y}(N)=0$ such that for all $y\in F\setminus N$ the following statements hold true:
  1. i)
    For all $n\geq 1$ we have $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))={\cal N}\bigl{(}\mu_{X|Y_{n}}(A_{n}(y)),C_{X|Y_{n}}\bigr{)}$ with
    $\mu_{X|Y_{n}}(A_{n}(y))$ $=\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(A_{n}(y)-\mathbb{E}Y_{n})\,,$
    $C_{X|Y_{n}}$ $=\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\,.$
  2. ii)
    The measure $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure on $E$.
  3. iii)
    The expectation $\mu_{X|Y}(y)$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is given by
    $\mu_{X|Y}(y)=\mathbb{E}X+\lim_{n\to\infty}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(A_{n}(y)-\mathbb{E}Y_{n})\,,$
    where the convergence is with respect to $\|\cdot\|_{E}$.
  4. iv)
    The covariance operator $C_{X|Y}:E^{\prime}\to E$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is given by
    $C_{X|Y}=\operatorname{cov}(X)-\lim_{n\to\infty}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\,,$
    where the convergence is with respect to the nuclear norm $\|\cdot\|_{\mathrm{nuc}}$ and thus also with respect to the operator norm.
  5. v)
    The weak convergence $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))\to\mathrm{P}_{X|Y}(\,\cdot\,|y)$ holds true.
  6. vi)
    If we define the random variable $Z:\Omega\to E$ by $Z(\omega):=0$ for all $\omega\in Y^{-1}(N)$ and
    $Z(\omega):=\mathbb{E}X+\lim_{n\to\infty}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\bigl{(}Y_{n}(\omega)-\mathbb{E}Y_{n}\bigr{)}\,,\qquad\qquad\omega\in\Omega\setminus Y^{-1}(N),$
    then $Z$ is both a version of $\mathbb{E}(X|Y)$ and a Gaussian random variable with $\mathbb{E}Z=\mathbb{E}X$ and
    $\operatorname{cov}(Z)$ $=\operatorname{cov}(X)-C_{X|Y}\,.$
    Finally, the limit in the definition of $Z$ also holds in the ${{\cal L}_{2}(\mathrm{P},E)}$-sense.
  7. vii)
    We have ${\cal G}(Z)\subset{\cal G}(Y)$ and the orthogonal projection $\Pi_{Y}:{\cal G}(X,Y)\to{\cal G}(X,Y)$ onto the subspace ${\cal G}(Y)$ is given by
    $\Pi_{Y}f=\mathbb{E}(f|Y)\,,\qquad\qquad f\in{\cal G}(X,Y).$
  8. viii)
    For all $x^{\prime}\in E^{\prime}$, for which $\langle x^{\prime},X\rangle_{E^{\prime},E}:\Omega\to\mathbb{R}$ is $\sigma(Y)$-measurable, and all $\omega\not\in Y^{-1}(N)$ we have
    $\bigl{\langle}x^{\prime},\mu_{X|Y}(Y(\omega))\bigr{\rangle}_{E^{\prime},E}$ $=\langle x^{\prime},X(\omega)\rangle_{E^{\prime},E}$
    $C_{X|Y}x^{\prime}$ $=0\,.$
Recall that [33] also obtain a nuclear norm convergence for their approximation scheme in the Hilbert space setting. Moreover, note that in the observational model (5), we have ${\cal G}(X,Y)={\cal G}(X)$ and ${\cal G}(Y)\subset{\cal G}(X)$, see Lemma 5.4, and therefore the conditional expectation considered in vi) becomes the orthogonal projection $\Pi_{Y}:{\cal G}(X)\to{\cal G}(X)$ onto the subspace ${\cal G}(Y)$.
Our next goal is to refine our analysis for both $\mathrm{P}_{X|Y}$ and its approximation. To this end, let us assume that we are in the situation of Theorem 3.4. Then each $\mu_{X|Y_{n}}\circ A_{n}:F\to E$ is strongly measurable, where we use the $\sigma$-algebra $\sigma(F^{\prime})=\mathfrak{B}(F)$ on $F$. Let us define $\mu_{X|Y}:F\to E$ by
$\mu_{X|Y}(y):=\lim_{n\to\infty}\boldsymbol{1}_{F\setminus N}(y)\cdot\mu_{X|Y_{n}}(A_{n}(y))\,,\qquad\qquad y\in Y\,,$ (16)
which is possible by and consistent with parts i) and iii). Then $\mu_{X|Y}$ is the pointwise limit of a sequence of strongly measurable functions, and therefore also (strongly) measurable. Moreover, the definition of $Z$ in part vi) gives
$Z=\mu_{X|Y}\circ Y\,,$
and since $Z$ is a version of $\mathbb{E}(X|Y)$, we see that $\mu_{X|Y}$ is a version of the “factorized” conditional expectation $\mathbb{E}(X|Y=\,\cdot\,):F\to E$, see also the discussion at the beginning of [27, Chapter 8.3] for the case of $\mathbb{R}$-valued $X$. Consequently, Theorem 3.4 shows that for each version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ we can construct a version $\mu_{X|Y}$ of $\mathbb{E}(X|Y=\,\cdot\,):F\to E$ by setting
$\mu_{X|Y}(y):=\begin{cases}\mathbb{E}\mathrm{P}_{X|Y}(\,\cdot\,|y)\,,&\mbox{ if }y\in F\setminus N\,,\\ 0\,,&\mbox{ else, }\end{cases}$
where $N\in\sigma(F^{\prime})$ is a suitable set with $\mathrm{P}_{Y}(N)=0$. Moreover, because of
$C_{X|Y}=\operatorname{cov}(X)-\operatorname{cov}(Z)=\operatorname{cov}(X)-\operatorname{cov}(\mu_{X|Y}\circ Y)$
and $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for all $y\in F\setminus N$, we see that $\mu_{X|Y}$ carries all information to reconstruct a version of $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ up to the $\mathrm{P}_{Y}$-zero set $N$. This raises the question, whether we also obtain a version of $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ by setting
$\mathrm{P}_{X|Y}(\,\cdot\,|y):={\cal N}\bigl{(}\mu(y),\operatorname{cov}(X)-\operatorname{cov}(\mu\circ Y)\bigr{)}\,,\qquad\qquad y\in Y$ (17)
for an arbitrary version $\mu$ of $\mathbb{E}(X|Y=\,\cdot\,):F\to E$, where we note that, for course, each such $\mu$ coincides $\mathrm{P}_{Y}$-almost surely with $\mu_{X|Y}$. The following theorem provides a positive answer to this question.
Theorem 3.5.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be jointly Gaussian random variables. Then (17) defines a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$ for all measurable $\mu:F\to E$, for which
$Z:=\mu\circ Y$
is a version of $\mathbb{E}(X|Y)$. In particular, in the case $F=\mathbb{R}^{n}$, these versions satisfy the assertions of Theorem 3.1 for $N=\emptyset$.
We have seen around (16) that a filtering sequence in combination with Theorem 3.4 can be used to construct a version of $\mathbb{E}(X|Y=\,\cdot\,):F\to E$. The next result shows that we can refine this construction to have a better control over the nuisance $\mathrm{P}_{Y}$-zero set $N$. Combining this with the construction (17) we can then also find a version $\mathrm{P}_{X|Y,\mathfrak{A}}(\,\cdot\,|\,\cdot\,)$ of which we have full control over its expectations.
Theorem 3.6.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be jointly Gaussian random variables. Moreover, let $\mathfrak{A}:=(A_{n})$ be a filtering sequence for $F$ and
$F_{\mathrm{conv},\mathfrak{A}}:=\bigl{\{}y\in F:\exists\lim_{n\to\infty}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(A_{n}(y)-\mathbb{E}Y_{n})\bigr{\}}\,,$
where $Y_{n}:=A_{n}\circ Y$ for all $n\geq 1$. Then the following statements hold true:
  1. i)
    We have $F_{\mathrm{conv},\mathfrak{A}}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(F_{\mathrm{conv},\mathfrak{A}})=1$.
  2. ii)
    The set $F_{X|Y,\mathfrak{A}}:=-\mathbb{E}Y+F_{\mathrm{conv},\mathfrak{A}}$ is a measurable subspace of $F$ with $\operatorname{ran}\operatorname{cov}(Y,X)\subset F_{X|Y,\mathfrak{A}}$.
  3. iii)
    The map $\mu_{X|Y,\mathfrak{A}}:F\to E$ defined by
    $\mu_{X|Y,\mathfrak{A}}(y):=\mathbb{E}X+\lim_{n\to\infty}\boldsymbol{1}_{F_{\mathrm{conv},\mathfrak{A}}}(y)\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(A_{n}(y)-\mathbb{E}Y_{n})\,,\qquad\qquad y\in Y\,,$
    is measurable and $\mu_{X|Y,\mathfrak{A}}\circ Y$ is a version of $\mathbb{E}(X|Y)$.
  4. iv)
    The map $\mathrm{P}_{X|Y,\mathfrak{A}}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ defined by
    $\mathrm{P}_{X|Y,\mathfrak{A}}(\,\cdot\,|y):={\cal N}\bigl{(}\mu_{X|Y,\mathfrak{A}}(y),\operatorname{cov}(X)-\operatorname{cov}(\mu_{X|Y,\mathfrak{A}}\circ Y)\bigr{)}\,,\qquad\qquad y\in Y$
    is a version of the regular conditional probability of $X$ given $Y$.
Let us assume for a moment, that $X$ and $Y$ are centered. Then $F_{\mathrm{conv},\mathfrak{A}}=F_{X|Y,\mathfrak{A}}$ holds, and therefore $F_{\mathrm{conv},\mathfrak{A}}$ is a measurable subspace of full measure $\mathrm{P}_{Y}$. Moreover, for all $y\in F_{\mathrm{conv},\mathfrak{A}}$ we have
$\mu_{X|Y,\mathfrak{A}}(y)=\lim_{n\to\infty}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}A_{n}(y)\,.$
Consequently, the restriction of $\mu_{X|Y,\mathfrak{A}}$ to $F_{\mathrm{conv},\mathfrak{A}}$ is linear and measurable. Note that such linear maps also occurred in [31] when considering conditional Gaussian distributions in the Hilbert space setup.
The theorems above also provide some concentration result. To illustrate this, let us assume that we have an $E$-valued Gaussian random variable $X$ and that we have fixed a filtering sequence $\mathfrak{A}:=(A_{n})$ for $E$. Moreover, assume that we cannot observe $X$ but only $X_{n}:=A_{n}\circ X$. This gives us versions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probabilities of $X$ given $X_{n}$ with
$\mathrm{P}_{X|X_{n}}(\,\cdot\,|A_{n}(x))={\cal N}\bigl{(}\mu_{X|X_{n}}(x),C_{X|X_{n}}\bigr{)}$
for all $x\in E$, where
$\mu_{X|X_{n}}(A_{n}(x))$ $=\mathbb{E}X+\operatorname{cov}(X,X_{n})(\operatorname{cov}X_{n})^{\dagger}(A_{n}(x)-\mathbb{E}X_{n})\,,$
$C_{X|X_{n}}$ $=\operatorname{cov}(X)-\operatorname{cov}(X,X_{n})(\operatorname{cov}X_{n})^{\dagger}\operatorname{cov}(X_{n},X)\,.$
To describe their behavior for $n\to\infty$, we set $Y:=X$. Then our results above give an $N\in{\cal A}$ with $\mathrm{P}(N)=0$ such that for all $\omega\in\Omega\setminus N$ we have
$\mu_{X|X_{n}}(X_{n}(\omega))\to\mu_{X|X}((X(\omega))=X(\omega)$ (18)
with convergence in $E$. Here, the last identity follows from part viii) of Theorem 3.4 applied to all $x^{\prime}\in\mathfrak{D}$ of a countable and $w^{*}$-dense subset $\mathfrak{D}$ of $B_{E^{\prime}}$, see Theorem 2.1, and a subsequent application of Hahn-Banach’s theorem. Moreover, part viii) of Theorem 3.4 also shows $C_{X|X}x^{\prime}=0$ for all $x^{\prime}\in E^{\prime}$, that is $C_{X|X}=0$, and hence we find
$C_{X|X_{n}}\to C_{X|X}=0$ (19)
with respect to the nuclear norm. Part v) of Theorem 3.4 thus yields the weak convergence
$\mathrm{P}_{X|X_{n}}(\,\cdot\,|X_{n}(\omega))\to{\cal N}\bigl{(}X(\omega),0\bigr{)}$ (20)
for all $\omega\in\Omega\setminus N$. In other words, with increasing $n$, the distributions $\mathrm{P}_{X|X_{n}}(\,\cdot\,|X_{n}(\omega))$ concentrate more and more around the observation $X(\omega)$. Note that these convergence results can also be interpreted from a frequentist point of view for the Bayes method given by the posterior distributions $\mathrm{P}_{X|X_{n}}$ and the prior distribution $\mathrm{P}_{X}$. Here, it is assumed that the “true” parameter is some $X(\omega_{0})$. The convergence results in (18) and (19) then provide a contraction result in the sense of [16, Section 7.3], while (20) is an additional, yet interesting convergence.
4 Examples with a Focus on the Gaussian Process Case
Theorem 3.4 crucially depends on the existence and choice of filtering sequences. In this section, we elaborate on this by first showing that for separable Banach spaces $F$ we can always find a filtering sequence. In addition, we quickly present some natural filtering sequences for Hilbert spaces and reproducing kernel Hilbert spaces. Finally, we provide a detailed discussion for the Gaussian process case presented in the introduction, by considering $E:=C(T)$ and $F:=C(S)$ for a closed $S\subset T$.
Our first result shows the existence of a filtering sequence for generic, separable $F$, where we recall from Theorem 2.1 that for such $F$ we always have a countable and $w^{*}$-dense $\mathfrak{D}_{F^{\prime}}\subset B_{F^{\prime}}$,
Proposition 4.1.
Let $F$ be a separable Banach space and $\mathfrak{D}_{F^{\prime}}\subset B_{F^{\prime}}$ be countable and $w^{*}$-dense with enumeration $\mathfrak{D}_{F^{\prime}}=\{y_{i}^{\prime}:i\geq 1\}$. For $n\geq 1$ we define $A_{n}:F\to\mathbb{R}^{n}$ by (15). Then $(A_{n})$ is a filtering sequence for $F$.
Combining Proposition 4.1 with Theorem 3.4 we see, for example, that the regular conditional probabilities $\mathrm{P}_{X|Y}$ of jointly Gaussian random variables are always almost surely Gaussian measures. Moreover, we can always approximate their means and covariances by the described finite dimensional projections.
We have already seen in Proposition 3.3 that sequences $(y_{i}^{\prime})\subset F^{\prime}$ whose span is $w^{*}$-sequentially dense in $F^{\prime}$ can be used to construct filtering sequences. The following result provides a characterization of such sequences that can be easier verified in some situations.
Proposition 4.2.
Let $F$ be a separable Banach space and $(y_{i}^{\prime})\subset F^{\prime}$ be a sequence. Then the following statements are equivalent:
  1. i)
    We have $\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}=F^{\prime}$.
  2. ii)
    The sequence $(y_{i}^{\prime})\subset F^{\prime}$ is separating, that is, for all $y\in F\setminus\{0\}$ there is an $i\geq 1$ with $\langle y^{\prime}_{i},y\rangle_{F^{\prime},F}\neq 0$.
Let us now turn to more specific sequences. The following lemma considers orthonormal bases (ONBs) in separable Hilbert spaces.
Lemma 4.3.
Let $H$ be a separable Hilbert space with $\dim H=\infty$ and $(e_{n})_{n\geq 1}$ be an ONB of $H$. For $n\geq 1$ we define $A_{n}:H\to\mathbb{R}^{n}$ by
$A_{n}(y):=\bigl{(}\langle e_{1},y\rangle_{H},\dots,\langle e_{n},y\rangle_{H}\bigr{)}\,,\qquad\qquad y\in H.$
Then $(A_{n})$ is a filtering sequence for $H$.
If $H$ is a separable reproducing kernel Hilbert space (RKHS) on a set $T$ with kernel $k:T\times T\to\mathbb{R}$, see e.g. [38, Chapter 4.2], then we can, of course, apply Lemma 4.3 to construct a filtering sequence. However, for RKHSs we have an additional generic way to construct a filtering sequence. To present this, let us recall, that $\operatorname{d}_{k}:T\times T\to[0,\infty)$ defined by
$\operatorname{d}_{k}(t_{1},t_{2}):=\|k(t_{1},\,\cdot\,)-k(t_{2},\,\cdot\,)\|_{H}=\sqrt{k(t_{1},t_{1})-2k(t_{1},t_{2})+k(t_{2},t_{2})}\,,\qquad\qquad t_{1},t_{2}\in T$
gives a pseudo-metric on $T$ called the kernel metric, see e.g. [38, Chapter 4.3]. With this preparation, the RKHS case reads as follows.
Proposition 4.4.
Let $H$ be a separable RKHS on $T$ with kernel $k$. Then there exists a countable $T_{0}\subset T$ that is dense in $T$ with respect to $\operatorname{d}_{k}$. Moreover, if we have a countable $T_{0}=\{t_{i}^{\prime}:i\geq 1\}$ that is $\operatorname{d}_{k}$-dense in $T$, then $A_{n}:H\to\mathbb{R}^{n}$ defined by
$A_{n}(y):=\bigl{(}\bigl{\langle}k(t_{1},\,\cdot\,),y\bigr{\rangle}_{H},\dots,\bigl{\langle}k(t_{n},\,\cdot\,),y\bigr{\rangle}_{H}\bigr{)}\,,\qquad\qquad y\in H$
gives filtering sequence $(A_{n})$ for $H$.
In most cases, working with the kernel metric may be somewhat cumbersome. For this reason we note that if $\operatorname{d}$ is a metric on $T$ such that $k$ is continuous with respect to $\operatorname{d}$, then $\operatorname{d}$-convergence implies $\operatorname{d}_{k}$-convergence, see e.g. [38, Lemma 4.29]. Consequently, if $(T,\operatorname{d})$ is separable, and $T_{0}$ is a countable and $\operatorname{d}$-dense subset of $T$, then it is also a $\operatorname{d}_{k}$-dense subset of $T$. In other words, we may replace $\operatorname{d}_{k}$ by $\operatorname{d}$ in Proposition 4.4.
Our next goal is to consider Gaussian random variables on the Banach space $C(T)$ of continuous functions, where $(T,\operatorname{d})$ is a compact metric space. To this end, we first note that in this case, $C(T)$ is a separable, see e.g. [11, Theorem V.6.6] or [14, Corollary 11.2.5]. To describe the dual of $C(T)$, we write ${\cal M}(T)$ for the space of all finite, signed Borel measures equipped with the total variation norm. Riesz’ representation theorem then shows that
^$\hat{\cdot}\,:{\cal M}(T)$ $\to C(T)^{\prime}$
$\nu$ $\mapsto\hat{\nu}:=\biggl{(}f\mapsto\int_{T}f\,\mathrm{d}\nu\biggr{)}$ (21)
is an isometric isomorphism, see e.g. [10, Theorem 7.3.6] and [4, Corollary 29.13], where for the first reference we note that compact metric spaces are Polish and Borel measures on Polish spaces are regular, and for the second reference we note that compact metric spaces have a countable base. In particular, the Dirac measures are mapped to the point the evaluation functionals $\delta_{t}:C(T)\to\mathbb{R}$ defined by $\delta_{t}f=f(t)$ for $t\in T$ and $f\in C(T)$.
To translate our main results to Gaussian processes, we also need a well-known, canonical one-to-one relationship between Gaussian processes with continuous paths and $C(T)$-valued Gaussian random variables. To present this relationship we fix a probability space $(\Omega,{\cal A},\mathrm{P})$. Now recall that a family $(X_{t})_{t\in T}$ of random variables $X_{t}:\Omega\to\mathbb{R}$ is called a stochastic process on $(\Omega,{\cal A},\mathrm{P})$. Moreover, it is called a Gaussian process if, for all $n\geq 1$ and all mutually distinct $t_{1},\dots,t_{n}\in T$, the random variable $(X_{t_{1}},\dots,X_{t_{n}}):\Omega\to\mathbb{R}^{n}$ is a Gaussian random variable. In addition, for a fixed $\omega\in\Omega$, the map
$X(\omega):T$ $\to\mathbb{R}$
$t$ $\mapsto X_{t}(\omega)$
is called a path of the stochastic process. We say that the process has continuous paths, if $X(\omega)\in C(T)$ for all $\omega\in\Omega$. Clearly, if only $\mathrm{P}$-almost all paths are continuous, then we can easily construct a version of the stochastic process that has continuous paths. In the following, we assume that we have already chosen such a version.
To describe the above mentioned relationship, let us now fix a map $X:\Omega\to C(T)$. Then we obtain a family $(X_{t})_{t\in T}$ of functions $X_{t}:\Omega\to\mathbb{R}$ by setting
$X_{t}(\omega):=\langle\delta_{t},X(\omega)\rangle_{C(T)^{\prime},C(T)}=(X(\omega))(t)\,,\qquad\qquad\omega\in\Omega.$ (22)
Obviously, this construction ensures that all maps $t\mapsto X_{t}(\omega)$ are continuous. Conversely, if we have a family $(X_{t})_{t\in T}$ of functions $X_{t}:\Omega\to\mathbb{R}$ such that for each $\omega\in\Omega$ the map $t\mapsto X_{t}(\omega)$ is continuous, then
$X:\Omega$ $\to C(T)$
$\omega$ $\mapsto\bigl{(}t\mapsto X_{t}(\omega)\bigr{)}$ (23)
defines a $C(T)$-valued map. Clearly, the operations (22) and (23) are inverse to each other. The following folklore result shows that they map $C(T)$-valued Gaussian random variables to Gaussian processes with continuous paths and vice versa.
Lemma 4.5.
Let $(T,\operatorname{d})$ be a compact metric space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $X:\Omega\to C(T)$ be a map, and $(X_{t})_{t\in T}$ be given by (22). Then the following statements hold true:
  1. i)
    $X$ is weakly measurable if and only if $(X_{t})_{t\in T}$ is a stochastic process with continuous paths.
  2. ii)
    $X$ is a Gaussian random variable if and only if $(X_{t})_{t\in T}$ is a Gaussian process with continuous paths.
It is well-known that a Gaussian process $(X_{t})_{t\in T}$ with continuous paths can be described by the mean and covariance functions given by (1) and (2). If $X$ is the Gaussian random variable associated to the process via (23), then the mean $\mathbb{E}X$ and the covariance operator $\operatorname{cov}(X)$ can be used to compute the mean and covariance function by
$m(t)$ $=\langle\delta_{t},\mathbb{E}X\rangle_{C(T)^{\prime},C(T)}\,,$ (24)
$k(s,t)$ $=\langle\delta_{s},\operatorname{cov}(X)\delta_{t}\rangle_{C(T)^{\prime},C(T)}$ (25)
for all $s,t\in T$, where we have used (A.11), respectively (B.1). Conversely, if we have the mean function $m$, then the identity above shows that we known the function $\mathbb{E}X\in C(T)$ pointwise, i.e. we know $\mathbb{E}X$. Moreover, the covariance function determines the covariance operator, namely, one can show
$\langle\hat{\mu},\operatorname{cov}(X)\hat{\nu}\rangle_{C(T)^{\prime},C(T)}=\int_{T}\int_{T}k(s,t)\,\mathrm{d}\mu(s)\,\mathrm{d}\nu(t)$
for all $\mu,\nu\in{\cal M}(T)$. These arguments justify the notation
${\cal N}(m,k):={\cal N}\bigl{(}\mathbb{E}X,\operatorname{cov}X\bigr{)}$
for the distribution of a given Gaussian random variable $X:\Omega\to C(T)$, where $m$ and $k$ are mean and covariance function of the associated stochastic process.
In summary, instead of working with a Gaussian random variable $X:\Omega\to C(T)$ we can alternatively consider Gaussian processes $(X_{t})_{t\in T}$ with continuous paths together with their mean and covariance function. In the following we will often adopt this point of view, which is common in the Gaussian process for machine learning literature, see e.g. [36, 24]. In particular, we speak of conditioned processes, when we condition the associated Gaussian random variables.
Let us now assume that we have a Gaussian processes $(X_{t})_{t\in T}$ with continuous paths and that we have an observation $(Y_{s})_{s\in S}$ of this process on a non-empty subset $S\subset T$. In other words, we have $Y_{s}(\omega)=X_{s}(\omega)$ for all $s\in S$ and $\omega\in\Omega$, where $(\Omega,{\cal A},\mathrm{P})$ is the underlying probability space and $(T,\operatorname{d})$ is a compact metric space. For technical simplicity, we assume that $S$ is closed, and thus also a compact metric space. To describe this setup by Gaussian random variables, we write $E:=C(T)$ and $F:=C(S)$ and denote the Gaussian random variables associated to the two processes above via (23) by $X:\Omega\to C(T)$ and $Y:\Omega\to C(S)$. This gives
$Y(\omega):=(X(\omega))_{|S}\,,\qquad\qquad\omega\in\Omega,$ (26)
that is $Y=\,\cdot\,_{|S}\circ X$, where the restriction operator $\,\cdot\,_{|S}:C(T)\to C(S)$ is bounded and continuous.
We first consider the case of a finite set $S=\{s_{1},\dots,s_{n}\}$. Here we write $I_{S}:C(T)\to\mathbb{R}^{n}$ for the bounded linear operator defined by
$I_{S}(f):=f(S):=\bigl{(}f(s_{1}),\dots,f(s_{n})\bigr{)}\,,\qquad\qquad f\in C(T)\,.$ (27)
Clearly, the general setup (26) reduces to $Y=I_{S}\circ X$ as $C(S)$ can be identified by $\mathbb{R}^{n}$. Finally, note that if we have another finite set $R=\{r_{1},\dots,r_{m}\}$, then the covariance operator $\operatorname{cov}(I_{R}\circ X,I_{S}\circ X):\mathbb{R}^{n}\to\mathbb{R}^{m}$ can be represented by the $m\times n$-matrix
$K_{R,S}:=\bigl{(}\operatorname{cov}(X_{r_{i}},X_{s_{j}})\bigr{)}_{i,j}=\bigl{(}k(r_{i},s_{j})\bigr{)}_{i,j}$
as discussed around (7). Furthermore, in the case $R=\{r\}$, we use the shorthand $K_{r,S}:=K_{R,S}$, and an analogous simplified notation is used in the case $S=\{s\}$.
With these preparations, the application of Theorem 3.1 to the observational model (26) reads as follows.
Theorem 4.6.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian processes on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (23) and $S:=\{s_{1},\dots,s_{n}\}\subset T$. Finally, let
$Y:=I_{S}\circ X\,,$ (28)
where $I_{S}:C(T)\to\mathbb{R}^{n}$ is defined by (27). Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(C(T))\times\mathbb{R}^{n}\to[0,1]$ of the regular conditional probability of $X$ given $Y$ there exist an $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$ such that for all $y\in\mathbb{R}^{n}\setminus N$ we have
$\mathrm{P}_{X|Y}(\,\cdot\,|y)={\cal N}\bigl{(}m_{X|Y=y},k_{X|Y}\bigr{)}\,,$
where the mean and covariance functions are given by
$m_{X|Y=y}(t)$ $=m(t)+K_{t,S}K_{S,S}^{\dagger}(y-m(S))\,,$
$k_{X|Y}(t_{1},t_{2})$ $=k(t_{1},t_{2})-K_{t_{1},S}K_{S,S}^{\dagger}K_{S,t_{2}}\,.$
Before we proceed with conditioning on observations on infinite $S$, we note that we can easily modify the observational model (28) to e.g.
$Y:=I_{S}\circ X+\varepsilon\,,$ (29)
where $\varepsilon\sim{\cal N}(0,\sigma^{2}I_{n})$ is independent of $X$, $\sigma\geq 0$, and $I_{n}$ denotes the $n$-dimensional identity matrix. Indeed, in view of Theorem 3.1, we only need to recompute the representing matrices of the operators $\operatorname{cov}(X,Y)$, $\operatorname{cov}(Y)$, and $\operatorname{cov}(Y,X)$. However, we have, for example,
$\operatorname{cov}(Y)$ $=\operatorname{cov}\bigl{(}I_{S}\circ X+\varepsilon,I_{S}\circ X+\varepsilon\bigr{)}$
$=\operatorname{cov}(I_{S}\circ X,I_{S}\circ X)+\operatorname{cov}(I_{S}\circ X,\varepsilon)+\operatorname{cov}(\varepsilon,I_{S}\circ X)+\operatorname{cov}(\varepsilon,\varepsilon)$
$=\operatorname{cov}(Y)+\operatorname{cov}(\varepsilon)\,,$
where in the last step we used the independence of $X$ and $\varepsilon$. Consequently, $\operatorname{cov}(Y)$ is represented by the matrix $K_{S,S}+\sigma^{2}I_{n}$, and the corresponding mean and covariance functions can be computed by
$m_{X|Y=y}(t)$ $=m(t)+K_{t,S}(K_{S,S}+\sigma^{2}I_{n})^{\dagger}(y-m(S))\,,$
$k_{X|Y}(t_{1},t_{2})$ $=k(t_{1},t_{2})-K_{t_{1},S}(K_{S,S}+\sigma^{2}I_{n})^{\dagger}K_{S,t_{2}}$
for all $t,t_{1},t_{2}\in T$. Not surprisingly, these formulas are the well-known textbook formulas from the Gaussian processes for machine learning literature, see e.g. [36]. However, there is is subtle but important difference in their meaning: The mentioned textbook formulas do not ensure that a Gaussian process with the above mean and covariance function has a version with continuous paths and even if this can be derived by other means, they do not ensure that the resulting measure of such a version on $C(T)$ equals the regular conditional measure we seek. In contrast, our approach establishes exactly these guarantees, so that the common functional interpretation of Gaussian processes for machine learning are indeed justified. This is similar in spirit of [45, Corollary 5], where more general functionals on $C(T)$ are considered in the noise free case $\varepsilon=0$.
Let us now turn to observations on infinite $S$. To apply Theorem 3.4 for this setting, we need a filtering sequence since the $C(S)$-valued random variable $Y:=X_{|S}$ lives in an infinite dimensional space. The following proposition provides a natural filtering sequence.
Proposition 4.7.
Let $(T,\operatorname{d})$ be a compact metric space and $T_{\infty}\subset T$ be a countable, dense subset with enumeration $T_{\infty}=\{t_{i}:i\geq 1\}$. For $n\geq 1$ we consider the evaluation operator $I_{n}:C(T)\to\mathbb{R}^{n}$ defined by
$I_{n}(f):=\bigl{(}f(t_{1}),\dots,f(t_{n})\bigr{)}\,,\qquad\qquad f\in C(T)$
Then $(I_{n})$ is a filtering sequence for $C(T)$.
Let us consider a closed subset $S\subset T$ of a compact metric space and a countable, dense $S_{\infty}\subset S$ with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. Then one can view the filtering sequence $(I_{n})$ of Proposition 4.7 as a mean to approximate the infinite dimensional observational random variable $Y:=X_{|S}$ with the help of the sequence $(Y_{n})$ of finite dimensional observations $Y_{n}:=X_{|S_{n}}$ taken at the points $S_{n}:=\{s_{1},\dots,s_{n}\}$. In other words, we approximate $Y$ on an arbitrarily fine, finite set of observational points. Since we have $Y_{n}=I_{n}\circ X=I_{S_{n}}\circ X$ the combination of Theorem 3.4 with Theorem 4.6 is straightforward and reads as follows.
Theorem 4.8.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian processes on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (23), $S\subset T$ be closed, and
$Y:=X_{|S}\,.$ (30)
Finally, let $S_{\infty}\subset S$ be countable and dense with enumeration $S_{\infty}=\{s_{i}:i\geq 1\}$. We write $S_{n}:=\{s_{1},\dots,s_{n}\}$ for all $n\geq 1$. Then for every version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(C(T))\times C(S)\to[0,1]$ of regular conditional probability of $X$ given $Y$ there exist an $N\in\mathfrak{B}(C(S))$ with $\mathrm{P}_{Y}(N)=0$ such that for all $g\in C(S)\setminus N$ we have
$\mathrm{P}_{X|Y}(\,\cdot\,|g)={\cal N}\bigl{(}m_{X|Y=y},k_{X|Y}\bigr{)}\,,$
where the mean and covariance functions are given by
$m_{X|Y=g}(t)$ $=m(t)+\lim_{n\to\infty}K_{t,S_{n}}K_{S_{n},S_{n}}^{\dagger}(g(S_{n})-m(S_{n}))\,,$ (31)
$k_{X|Y}(t_{1},t_{2})$ $=k(t_{1},t_{2})-\lim_{n\to\infty}K_{t_{1},S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t_{2}}\,$ (32)
where the convergence is uniform in $t\in T$, respectively in $t_{1},t_{2}\in T$. Moreover, for all $g\in C(S)\setminus N$, $s\in S$, and $t\in T$ we have
$m_{X|Y=g}(s)$ $=g(s)\,,$
$k_{X|Y}(s,t)$ $=0\,.$
Of course, the remaining assertions of Theorem 3.4 also hold in the special setup considered in Theorem 4.8. In particular note that the convergence of the covariance operators in Theorem 3.4 are with respect to the nuclear norm, while the uniform convergence of the covariance functions stated in Theorem 4.8 is a consequence of the weaker operator norm convergence. In addition, it is, of course, possible to translate Theorems 3.5 and 3.6 to the $C(T)$-valued case with the help of Lemma 6.2.
Similarly, the concentration phenomenon for $Y=X$ discussed around (20) holds true in the $C(T)$-valued case, too. However, a bit more can be said in the situation of Theorem 4.8. To this end, let us assume that we have a countable set $S_{\infty}=\{s_{i}:i\geq 1\}$ with $S=\overline{S_{\infty}}$ and we wish to approximate $m_{X|Y=g}$ and $k_{X|Y}$. Then Theorem 4.8 tells us that (31) and (32) give us a uniform approximation of $m_{X|Y=g}$, respectively $k_{X|Y}$, and that these approximations satisfy
$m(s)+K_{s,S_{n}}K_{S_{n},S_{n}}^{\dagger}(g(S_{n})-m(S_{n}))$ $\to g(s)\,,$
$k(s,t)-K_{s,S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t}$ $\to 0$
for $\mathrm{P}_{Y}$-almost all observed paths $g\in C(S)$ and all $s\in S$ and $t\in T$, where again the convergence is uniform in $s$, respectively $s$ and $t$.
5 Proofs of the Results of Section 3
5.1 Some Auxiliary Results for the Proof of Theorem 3.1
The following theorem shows that the Gaussianity of an $E$-valued random variable can be checked with the help of suitable, countable subsets of $E^{\prime}$.
Theorem 5.1.
Let $E$ be a separable Banach space and $\mathfrak{D}\subset B_{E^{\prime}}$ be $w^{*}$-dense in $B_{E^{\prime}}$. Then a probability measure $\mathrm{P}$ on $\mathfrak{B}(E)=\sigma(E^{\prime})$ is a Gaussian measure, if and only if the image measure $\mathrm{P}_{x^{\prime}}$ is a Gaussian measure on $\mathfrak{B}$ for all $x^{\prime}\in\mathfrak{D}$.
Moreover, an $E$-valued, weakly measurable random variable $X$ is a Gaussian random variable, if and only if $\langle x^{\prime},X\rangle_{E^{\prime},E}$ is an $\mathbb{R}$-valued Gaussian random variable for all $x^{\prime}\in\mathfrak{D}$.
Proof of Theorem 5.1: Let us first consider the case of an $E$-valued and weakly measurable random variable $X:\Omega\to E$. We begin by assuming that $\mathrm{P}_{x^{\prime}}$ is a Gaussian measure for all $x^{\prime}\in\mathfrak{D}$. To this end, we fix an $x^{\prime}\in E^{\prime}$. Since the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(x^{\prime}_{n})\subset\mathfrak{D}$ with $\langle x_{n}^{\prime},x\rangle\to\langle x^{\prime},x\rangle$ for all $x\in E$. Consequently, we have
$\langle x_{n}^{\prime},X(\omega)\rangle_{E^{\prime},E}\to\langle x^{\prime},X(\omega)\rangle_{E^{\prime},E}\,,\qquad\qquad\omega\in\Omega,$
and this in turn shows that $\langle x_{n}^{\prime},X\rangle_{E^{\prime},E}\to\langle x^{\prime},X\rangle_{E^{\prime},E}$ in distribution. Now, all $\langle x^{\prime}_{n},X\rangle$ are $\mathbb{R}$-valued Gaussian random variables by assumption, and consequently Theorem C.4 ensures that $\langle x^{\prime},X\rangle$ is also a Gaussian random variable. This shows that $X$ is a Gaussian random variable. The converse implication is trivial.
Finally, the assertion for Gaussian measures follows from considering the random variable $X:=\operatorname{id}_{E}:E\to E$ on the probability space $(E,\sigma(E^{\prime}),\mathrm{P})$, since in this case we have $\mathrm{P}_{X}=\mathrm{P}$. ∎
Given a bounded linear operator $T:E^{\prime}\to E$ and an $x^{\prime}\in E^{\prime}$ with dual $(x^{\prime})^{\prime}:\mathbb{R}^{\prime}\to E^{\prime}$, we can consider the bounded linear map $x^{\prime}\circ T\circ(x^{\prime})^{\prime}:\mathbb{R}^{\prime}\to\mathbb{R}$. Since both $\mathbb{R}^{\prime}$ and $\mathbb{R}$ are one-dimensional, this map can be described by a $1\times 1$-matrix with respect to the canonical bases of $\mathbb{R}^{\prime}$ and $\mathbb{R}$. The following lemma determines this matrix.
Lemma 5.2.
Let $E$ be a Banach space and $T:E^{\prime}\to E$ be a bounded, linear operator. Then we have
$\bigl{(}x^{\prime}\circ T\circ(x^{\prime})^{\prime}\bigr{)}(\langle t,\,\cdot\,\rangle_{\mathbb{R}})=\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}\cdot t\,,\qquad\qquad x^{\prime}\in E^{\prime},t\in\mathbb{R},$
where $(x^{\prime})^{\prime}:\mathbb{R}^{\prime}\to E^{\prime}$ denotes the dual of $x^{\prime}:E\to\mathbb{R}$ and $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathbb{R}}$ denotes the inner product of the Hilbert space $\mathbb{R}$.
Proof of Lemma 5.2: We fix an $x^{\prime}\in E^{\prime}$ and recall that the elements in $\mathbb{R}^{\prime}$ can be represented by $\mathbb{R}$ via the canonical isomorphism $t\mapsto\langle t,\,\cdot\,\rangle_{\mathbb{R}}$. Now observe that for $t\in\mathbb{R}$ and $x\in E$ we have
$\bigl{\langle}(x^{\prime})^{\prime}(\langle t,\,\cdot\,\rangle_{\mathbb{R}}),x\bigr{\rangle}_{E^{\prime},E}=\bigl{\langle}\langle t,\,\cdot\,\rangle_{\mathbb{R}},x^{\prime}(x)\bigr{\rangle}_{\mathbb{R}^{\prime},\mathbb{R}}=\langle t,x^{\prime}(x)\rangle_{\mathbb{R}}=t\cdot x^{\prime}(x)=\langle tx^{\prime},x\rangle_{E^{\prime},E}\,.$
In other words, we have $(x^{\prime})^{\prime}(\langle t,\,\cdot\,\rangle_{\mathbb{R}})=tx^{\prime}$ for all $t\in\mathbb{R}$. This in turn yields
$\bigl{(}x^{\prime}\circ T\circ(x^{\prime})^{\prime}\bigr{)}(\langle 1,\,\cdot\,\rangle_{\mathbb{R}})=(x^{\prime}\circ T)\bigl{(}(x^{\prime})^{\prime}(\langle 1,\,\cdot\,\rangle_{\mathbb{R}})\bigr{)}=(x^{\prime}\circ T)(x^{\prime})=x^{\prime}(Tx^{\prime})=\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}\,,$
and by multiplying both sides with $t\in\mathbb{R}$ we obtain the assertion. ∎
By definition, every one-dimensional projection $\langle x^{\prime},X\rangle_{E^{\prime},E}$ of a centered Gaussian random variable is a centered Gaussian random variable, and it is obvious that the set of these collections form a linear space. This motivates the following definition taken from [23].
Definition 5.3.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space and $H\subset{L_{2}(\mathrm{P})}$. Then $H$ is called a Gaussian Hilbert space, if $H$ is a closed subspace of ${L_{2}(\mathrm{P})}$ and every $h\in H$ is a centered Gaussian random variable.
The next result connects Gaussian Hilbert space to Gaussian random variables.
Lemma 5.4.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ be a separable Banach space, and $X:\Omega\to E$ be a Gaussian random variable. Then
${\cal G}(X):=\overline{\bigl{\{}\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}:x^{\prime}\in E^{\prime}\bigr{\}}}^{{L_{2}(\mathrm{P})}}$
is a Gaussian Hilbert space and each $\mathrm{P}$-equivalence class in ${\cal G}(X)$ has a $\sigma(X)$-measurable representative. In addition, if $F$ is another separable Banach space, then the following statements hold true:
  1. i)
    If $A:E\to F$ is a bounded and linear operator, then ${\cal G}(A\circ X)\subset{\cal G}(X)$.
  2. ii)
    If $Y:\Omega\to F$ is a random variable such that $X$ and $Y$ are jointly Gaussian random variables, then
    ${\cal G}(X,Y):={\cal G}((X,Y))=\overline{{\cal G}(X)+{\cal G}(Y)}^{{L_{2}(\mathrm{P})}}\,.$
Proof of Lemma 5.4: Without loss of generality we assume that $X$ and $Y$ are centered.
Now, [23, Example 1.8] shows that ${\cal G}(X)$ is a Gaussian Hilbert space. However, this also directly follows from the observation that ${\cal G}(X)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by construction and from the fact that ${L_{2}(\mathrm{P})}$-limits of centered Gaussian random variables are centered Gaussian random variables, see also Theorem C.4.
Moreover, each $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{\cal G}(X)$ is $\sigma(X)$-measurable, and therefore there is nothing to prove for the $\mathrm{P}$-equivalence classes of these functions. Let us now fix an arbitrary $f\in{\cal G}(X)$, so that our goal is to find a $\sigma(X)$-measurable $g:\Omega\to\mathbb{R}$ with $\mathrm{P}(\{f\neq g\})=0$. To this end, we first note that by the definition of ${\cal G}(X)$, there exists a sequence $(x^{\prime}_{n})\subset E^{\prime}$ with $f_{n}:=\langle x^{\prime}_{n},X\rangle_{E^{\prime},E}\to f$ in ${L_{2}(\mathrm{P})}$. Since there exists a $\mathrm{P}$-almost surely converging subsequence, see e.g. [27, Corollary 6.14], we may assume without loss of generality that we actually have the $\mathrm{P}$-almost sure convergence $f_{n}\to f$. Moreover, each $f_{n}$ is $\sigma(X)$-measurable, and therefore we have
$A:=\bigl{\{}\omega\in\Omega:\exists\lim_{n\to\infty}f_{n}(\omega)\bigr{\}}\in\sigma(X)\,,$
see e.g. [7, Lemma 2.1.7]. In addition, the $\mathrm{P}$-almost sure convergence $f_{n}\to f$ shows $\mathrm{P}(A)=1$. We define $g_{n}:=\boldsymbol{1}_{A}f_{n}$ and $g:\Omega\to\mathbb{R}$ by $g(\omega):=\lim f_{n}(\omega)$ for $\omega\in A$ and $g(\omega):=0$ for $\omega\in\Omega\setminus A$. Since each $g_{n}$ is $\sigma(X)$-measurable and $g_{n}(\omega)\to g(\omega)$ for all $\omega\in\Omega$ we then see that $g$ is $\sigma(X)$-measurable. Moreover, we have
$\mathrm{P}(\{f\neq g\})=\mathrm{P}(A\cap\{f\neq g\})+\mathrm{P}\bigl{(}(\Omega\setminus A)\cap\{f\neq g\})\bigr{)}=\mathrm{P}(A\cap\{f\neq\lim f_{n}\})=0\,,$
and hence we have found the desired $\sigma(X)$-measurable representative of the $\mathrm{P}$-equivalence class of $f$.
i). For $y^{\prime}\in F^{\prime}$ we have $x^{\prime}:=y^{\prime}\circ A\in E^{\prime}$, and this yields
$\langle y^{\prime},A\circ X\rangle_{F^{\prime},F}=y^{\prime}\circ A\circ X=\langle x^{\prime},X\rangle_{E^{\prime},E}\in{\cal G}(X)\,.$
In other words we have $\{\langle y^{\prime},A\circ X\rangle_{F^{\prime},F}:y^{\prime}\in F^{\prime}\}\subset{\cal G}(X)$ and taking the ${L_{2}(\mathrm{P})}$-closure then yields the assertion.
ii). The inclusion $\supset$” follows from i) by considering the projections $\pi_{E}:E\times F\to E$ and $\pi_{F}:E\times F\to F$ and the fact that ${\cal G}(X,Y)$ is a closed linear subspace of ${L_{2}(\mathrm{P})}$. To prove the converse implication, we fix an $z^{\prime}\in(E\times F)^{\prime}$. As discussed in front of Definition 2.4, there then exist $x^{\prime}\in E^{\prime}$ and $y^{\prime}\in F^{\prime}$ such that $z^{\prime}(x,y)=x^{\prime}(x)+y^{\prime}(y)$ for all $x\in E$ and $y\in F$, and this in turn implies
$\langle z^{\prime},(X,Y)\rangle_{(E\times F)^{\prime},E\times F}=\langle x^{\prime},X\rangle_{E^{\prime},E}+\langle y^{\prime},Y\rangle_{F^{\prime},F}\in{\cal G}(X)+{\cal G}(Y)\,.$
Taking the ${L_{2}(\mathrm{P})}$-closure thus gives the inclusion $\subset$”. ∎
5.2 Proof of Theorem 3.1
Proof of Theorem 3.1: We begin by some preparations. To this end, we fix a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$. Moreover, we choose a countable, $w^{*}$-dense $\mathfrak{D}\subset B_{E^{\prime}}$ with the help of Theorem 2.1. Finally, we fix an enumeration $(x^{\prime}_{i})$ of $\mathfrak{D}$, that is, $\mathfrak{D}=\{x_{i}^{\prime}:i\geq 1\}$.
For $i\geq 1$ we now define $X_{i}:=\langle x^{\prime}_{i},X\rangle_{E^{\prime},E}$. Then $X_{i}:\Omega\to\mathbb{R}$ and $Y:\Omega\to\mathbb{R}^{n}$ are jointly Gaussian random variables with joint distribution $\mathrm{P}_{(X_{i},Y)}$ on ×$\mathbb{R}\times\mathbb{R}^{n}$. We fix a version $\mathrm{P}_{X_{i}|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X_{i}$ given $Y$. By Theorem 2.6 there then exists an $N_{i}^{\prime}\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N_{i}^{\prime})=0$ such that for all $y\in\mathbb{R}^{n}\setminus N_{i}^{\prime}$ we have
$\mathrm{P}_{X_{i}|Y}(\,\cdot\,|y)={\cal N}\bigl{(}\mu_{X_{i}|Y}(y),K_{X_{i}|Y}\bigr{)}\,,$
where the expectation and covariance are given by
$\mu_{X_{i}|Y}(y)$ $:=\mathbb{E}X_{i}+K_{X_{i},Y}K_{Y,Y}^{\dagger}(y-\mathbb{E}Y)\,,$
$K_{X_{i}|Y}$ $:=K_{X_{i},X_{i}}-K_{X_{i},Y}K_{Y,Y}^{\dagger}K_{Y,X_{i}}\,.$
Using the definition of $X_{i}$, the expression for the expectation reads as
$\mu_{X_{i}|Y}(y)$ $=\mathbb{E}X_{i}+K_{X_{i},Y}K_{Y,Y}^{\dagger}(y-\mathbb{E}Y)$
$=\langle x^{\prime}_{i},\mathbb{E}X\rangle_{E^{\prime},E}+\operatorname{cov}(x^{\prime}_{i}\circ X,Y)\operatorname{cov}(Y)^{\dagger}(y-\mathbb{E}Y)$
$=\langle x^{\prime}_{i},\mathbb{E}X\rangle_{E^{\prime},E}+\bigl{\langle}x^{\prime}_{i},\operatorname{cov}(X,Y)\operatorname{cov}(Y)^{\dagger}(y-\mathbb{E}Y)\bigr{\rangle}_{E^{\prime},E}$
$=\bigl{\langle}x^{\prime}_{i},\mathbb{E}X+\operatorname{cov}(X,Y)\operatorname{cov}(Y)^{\dagger}(y-\mathbb{E}Y)\bigr{\rangle}_{E^{\prime},E}\,,$ (33)
where we have used the permutability of integration and operator application (A.11) and the cross covariance formula for linearly transformed random variables (B.3). To investigate $K_{X_{i}|Y}$ in a similar fashion, we consider the bounded linear operator $R_{X,Y}:E^{\prime}\to E$ defined by
$R_{X,Y}:=\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)\,.$
By (B.3) we then find
$\operatorname{cov}(X_{i},Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X_{i})=\operatorname{cov}(x^{\prime}_{i}\circ X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,x^{\prime}_{i}\circ X)=x^{\prime}_{i}\circ R_{X,Y}\circ(x^{\prime}_{i})^{\prime}\,.$
Moreover, Lemma 5.2 shows that the operator $x^{\prime}_{i}\circ R_{X,Y}\circ(x^{\prime}_{i})^{\prime}:\mathbb{R}^{\prime}\to\mathbb{R}$ is described by the $1\times 1$-matrix $\langle x_{i}^{\prime},R_{X,Y}x_{i}^{\prime}\rangle_{E^{\prime},E}$, while the definition of the covariance matrices $K_{X_{i},Y}$, $K_{Y,Y}$, and $K_{Y,X_{i}}$ shows that $\operatorname{cov}(X_{i},Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X_{i}):\mathbb{R}^{\prime}\to\mathbb{R}$ is described by the $1\times 1$-matrix $K_{X_{i},Y}K_{Y,Y}^{\dagger}K_{Y,X_{i}}$. Consequently, we have
$K_{X_{i},Y}K_{Y,Y}^{\dagger}K_{Y,X_{i}}=\langle x_{i}^{\prime},R_{X,Y}x_{i}^{\prime}\rangle_{E^{\prime},E}\,.$ (34)
In addition, (B.6) shows
$K_{X_{i},X_{i}}=\operatorname{var}(X_{i})=\operatorname{var}(\langle x_{i}^{\prime},X\rangle_{E^{\prime},E})=\langle x_{i}^{\prime},\operatorname{cov}(X)x_{i}^{\prime}\rangle_{E^{\prime},E}\,.$
Combining both considerations, we thus find
$K_{X_{i}|Y}=K_{X_{i},X_{i}}-K_{X_{i},Y}K_{Y,Y}^{\dagger}K_{Y,X_{i}}=\langle x_{i}^{\prime},\operatorname{cov}(X)x_{i}^{\prime}\rangle_{E^{\prime},E}-\langle x_{i}^{\prime},R_{X,Y}x_{i}^{\prime}\rangle_{E^{\prime},E}\,.$ (35)
Our next goal is to describe the relationship between $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ and the just considered $\mathrm{P}_{X_{i}|Y}(\,\cdot\,|\,\cdot\,)$ with the help of Theorem D.2, which describes regular conditional probabilities under transformations. To this end, we set $T_{0}:=E$ and $T_{1}:=\mathbb{R}$, as well as $U_{0}:=U_{1}:=\mathbb{R}^{n}$ and ${\cal A}_{0}:={\cal A}_{1}:=\mathfrak{B}^{n}$. Moreover, we consider the maps $\Phi:=x_{i}^{\prime}$ and $\Psi:=\operatorname{id}_{\mathbb{R}^{n}}$ as well as the measures $\mathrm{P}^{(0)}:=\mathrm{P}_{(X,Y)}$ and $\mathrm{P}^{(1)}:=\mathrm{P}_{(X_{i},Y)}=\mathrm{P}^{(1)}_{(\Phi,\Psi)}$. By Theorem D.2 there then exists an $N_{i}^{\prime\prime}\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N_{i}^{\prime\prime})=0$ such that for all $y\in\mathbb{R}^{n}\setminus N_{i}^{\prime\prime}$ we have
$\mathrm{P}_{X|Y}\bigl{(}(x_{i}^{\prime})^{-1}(B)|y\bigr{)}=\mathrm{P}_{X_{i}|Y}(B|y)\,,\qquad\qquad B\in\mathfrak{B}.$
Consequently, we find
$\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}=\mathrm{P}_{X_{i}|Y}(\,\cdot\,|y)={\cal N}\bigl{(}\mu_{X_{i}|Y}(y),K_{X_{i}|Y}\bigr{)}$ (36)
for all $y\in\mathbb{R}^{n}\setminus(N_{i}^{\prime}\cup N_{i}^{\prime\prime})$. We define
$N:=\bigcup_{i\geq 1}(N_{i}^{\prime}\cup N_{i}^{\prime\prime})\,.$
Obviously, this gives $N\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y}(N)=0$, and by construction (36) holds for all $y\in\mathbb{R}^{n}\setminus N$ and $x^{\prime}_{i}\in\mathfrak{D}$. Let us fix a $y\in\mathbb{R}^{n}\setminus N$. Then we have just seen in (36) that for all $x_{i}^{\prime}\in\mathfrak{D}$ the image measure $(\mathrm{P}_{X|Y}(\,\cdot\,|y))_{x^{\prime}_{i}}$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ under $x_{i}^{\prime}$ is Gaussian, and by Theorem 5.1 we can thus conclude that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian.
Next we verify the formulas for the expectation and covariance. Let us first consider the expectation: For all $y\in\mathbb{R}^{n}\setminus N$ and $x^{\prime}_{i}\in\mathfrak{D}$, the identity (36) shows
$\mu_{X_{i}|Y}(y)=\mathbb{E}\bigl{(}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}\bigr{)}=\int_{\mathbb{R}}\operatorname{id}_{\mathbb{R}}(t)\,\mathrm{d}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}(t)$ $=\int_{E}(\operatorname{id}_{\mathbb{R}}\circ x_{i}^{\prime})(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
$=\int_{E}\langle x_{i}^{\prime},x\rangle_{E^{\prime},E}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
$=\biggl{\langle}x_{i}^{\prime},\int_{E}x\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)\biggr{\rangle}_{E^{\prime},E}$
$=\bigl{\langle}x_{i}^{\prime},\mu_{X|Y}(y)\bigr{\rangle}_{E^{\prime},E}\,,$ (37)
where in the third step we used the transformation formula for image measures, and in the last step we used that $\operatorname{id}_{E}$ is Bochner-integrable with respect to $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ since $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure, and thus (8) holds true. Using (33) we thus find
$\bigl{\langle}x_{i}^{\prime},\mu_{X|Y}(y)\bigr{\rangle}_{E^{\prime},E}=\bigl{\langle}x^{\prime}_{i},\mathbb{E}X+\operatorname{cov}(X,Y)\operatorname{cov}(Y)^{\dagger}(y-\mathbb{E}Y)\bigr{\rangle}_{E^{\prime},E}$
for all $x_{i}^{\prime}\in\mathfrak{D}$. Since $\mathfrak{D}$ is $w^{*}$-dense in $B_{E^{\prime}}$, a simple limit argument then shows that the latter identity not only holds for all $x_{i}^{\prime}$ but actually for all $x^{\prime}\in B_{E^{\prime}}$. By Hahn-Banach’s theorem we thus find
$\mu_{X|Y}(y)=\mathbb{E}X+\operatorname{cov}(X,Y)\operatorname{cov}(Y)^{\dagger}(y-\mathbb{E}Y)\,,\qquad\qquad y\in\mathbb{R}^{n}\setminus N.$
To verify the formula for the covariance, we again fix an $y\in\mathbb{R}^{n}\setminus N$. For $x_{i}^{\prime}\in\mathfrak{D}$, the Identity (36) and the fact that $\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}$ is a distribution on $\mathbb{R}$ then yields
$K_{X_{i}|Y}=\operatorname{cov}\Bigl{(}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}\Bigr{)}=\operatorname{var}\Bigl{(}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}\Bigr{)}$ $=\int_{\mathbb{R}}\bigl{(}\operatorname{id}_{\mathbb{R}}(t)-\mu_{X_{i}|Y}(y)\bigr{)}^{2}\,\mathrm{d}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}(t)$
$=\int_{E}\bigl{(}\operatorname{id}_{\mathbb{R}}\circ x_{i}^{\prime}-\mu_{X_{i}|Y}(y)\bigr{)}^{2}(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
$=\int_{E}\bigl{(}\langle x_{i}^{\prime},x\rangle_{E^{\prime},E}-\mu_{X_{i}|Y}(y)\bigr{)}^{2}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
$=\int_{E}\bigl{\langle}x_{i}^{\prime},x-\mu_{X|Y}(y)\bigr{\rangle}_{E^{\prime},E}^{2}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
$=\bigl{\langle}x_{i}^{\prime},\operatorname{cov}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}x_{i}^{\prime}\bigr{\rangle}_{E^{\prime},E}\,,$
where we have also used a projection formula for covariances (B.6), the basic transformation formula for $\mathbb{R}$-valued integrals, and (37). Combining this with (35) we thus find
$\bigl{\langle}x_{i}^{\prime},\operatorname{cov}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}x_{i}^{\prime}\bigr{\rangle}_{E^{\prime},E}=\langle x_{i}^{\prime},\operatorname{cov}(X)x_{i}^{\prime}\rangle_{E^{\prime},E}-\bigl{\langle}x_{i}^{\prime},R_{X,Y}x_{i}^{\prime}\bigr{\rangle}_{E^{\prime},E}$
for all $x_{i}^{\prime}\in\mathfrak{D}$ and $y\in\mathbb{R}^{n}\setminus N$. Now the left hand side of this identity is $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3, and the right-hand side is $w^{*}$-continuous on $B_{E^{\prime}}$ by Theorem B.3 and Lemma B.4. Since $\mathfrak{D}$ is $w^{*}$-dense in $B_{E^{\prime}}$, we thus find
$\bigl{\langle}x^{\prime},\operatorname{cov}\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}x^{\prime}\bigr{\rangle}_{E^{\prime},E}$ $=\langle x^{\prime},\operatorname{cov}(X)x^{\prime}\rangle_{E^{\prime},E}-\bigl{\langle}x^{\prime},R_{X,Y}x^{\prime}\bigr{\rangle}_{E^{\prime},E}\,.$
for all $x^{\prime}\in B_{E^{\prime}}$. By scaling this implies
$\bigl{\langle}x^{\prime},\bigl{(}\operatorname{cov}(\mathrm{P}_{X|Y}(\,\cdot\,|y))+R_{X,Y}\bigr{)}x^{\prime}\bigr{\rangle}_{E^{\prime},E}=\langle x^{\prime},\operatorname{cov}(X)x^{\prime}\rangle_{E^{\prime},E}$ (38)
for all $x^{\prime}\in E^{\prime}$. Moreover, $(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$ is symmetric and non-negative, see the discussion around (B.8), and therefore Lemma B.4 shows that $R_{X,Y}$ is an abstract covariance operator. This in turn shows that
$\operatorname{cov}(\mathrm{P}_{X|Y}(\,\cdot\,|y))+R_{X,Y}:E^{\prime}\to E$
is an abstract covariance operator. In addition, $\operatorname{cov}(X):E^{\prime}\to E$ is also an abstract covariance operator, and (38) shows that both abstract covariance operators coincide on the diagonal. Lemma B.2 thus yields $\operatorname{cov}(\mathrm{P}_{X|Y}(\,\cdot\,|y))+R_{X,Y}=\operatorname{cov}(X)$, that is, we have found the formula for the covariance.
Let us now prove the claims made for $Z$. To this end, we write $\pi_{E}:E\times\mathbb{R}^{n}\to E$ and $\pi_{\mathbb{R}^{n}}:E\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ for the coordinate projections onto $E$ and $\mathbb{R}^{n}$. Moreover, we define $N_{0}:=Y^{-1}(N)$, which yields $N_{0}\in{\cal A}$ with $\mathrm{P}(N_{0})=0$. For $\omega\in\Omega\setminus N_{0}$ and $y:=Y(\omega)$ we then have $y\in\mathbb{R}^{n}\setminus N$ and therefore $\mu_{X|Y}(y)$ is the mean of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$. Moreover, since $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian, we have $\operatorname{id}_{E}\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),E)}$. For $g:=\operatorname{id}_{E}$ this gives $g\circ\pi_{E}(\,\cdot\,,y)=g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),E)}$ for all $y\in\mathbb{R}^{n}\setminus N$, and therefore, the random variable $Z_{g}:\mathbb{R}^{n}\to E$ defined by
$Z_{g}(y):=\int_{E}\boldsymbol{1}_{\mathbb{R}^{n}\setminus N}(y)\cdot\bigl{(}(g\circ\pi_{E})(x,y)\bigr{)}\,\mathrm{d}\mathrm{P}_{X|Y}(\,\mathrm{d}x|y)\,,\qquad\qquad y\in\mathbb{R}^{n},$
provides a version $Z_{g}\circ\pi_{\mathbb{R}^{n}}$ of $\mathbb{E}(g\circ\pi_{E}|\pi_{\mathbb{R}^{n}}^{-1}(\mathfrak{B}^{n}))=\mathbb{E}(g\circ\pi_{E}|E\times\mathfrak{B}^{n})$ by Theorem D.6 applied to the probability measure $\mathrm{P}_{(X,Y)}$ on $\mathfrak{B}(E)\otimes\mathfrak{B}^{n}=\sigma(E^{\prime})\otimes\mathfrak{B}^{n}$. In other words, we have
$\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot(Z_{g}\circ\pi_{\mathbb{R}^{n}})\,\mathrm{d}\mathrm{P}_{(X,Y)}=\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{P}_{(X,Y)}$ (39)
for all $A\in\mathfrak{B}^{n}$. Moreover, for $\omega\in\Omega\setminus N_{0}$ we find
$\mu_{X|Y}(Y(\omega))=\int_{E}x\,\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$ $=\int_{E}\boldsymbol{1}_{\mathbb{R}^{n}\setminus N}(Y(\omega))\cdot\bigl{(}(g\circ\pi_{E})(x,Y(\omega))\bigr{)}\,\mathrm{d}\mathrm{P}_{X|Y}(\,\mathrm{d}x|Y(\omega))$
$=Z_{g}(Y(\omega))\,,$
which shows $\mathrm{P}(\{Z\neq Z_{g}\circ Y\})=0$. Since $Z$ is $\sigma(Y)$-measurable by construction, it thus suffices to show that $Z_{g}\circ Y$ is a version of $\mathbb{E}(X|Y)$. Here the $\sigma(Y)$-measurability of $Z_{g}\circ Y$ is ensured by construction. Moreover, for $A_{0}\in\sigma(Y)$ there exists an $A\in\mathfrak{B}^{n}$ with $A_{0}=Y^{-1}(A)$, and with this we obtain
$\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot(Z_{g}\circ Y)\,\mathrm{d}\mathrm{P}=\int_{\Omega}(\boldsymbol{1}_{A}\circ Y)\cdot(Z_{g}\circ Y)\,\mathrm{d}\mathrm{P}$ $=\int_{\mathbb{R}^{n}}\boldsymbol{1}_{A}Z_{g}\,\mathrm{d}\mathrm{P}_{Y}$
$=\int_{E\times\mathbb{R}^{n}}(\boldsymbol{1}_{A}\circ\pi_{\mathbb{R}^{n}})\cdot(Z_{g}\circ\pi_{\mathbb{R}^{n}})\,\mathrm{d}\mathrm{P}_{(X,Y)}$
$=\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot(Z_{g}\circ\pi_{\mathbb{R}^{n}})\,\mathrm{d}\mathrm{P}_{(X,Y)}$
as well as
$\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot X\,\mathrm{d}\mathrm{P}=\int_{\Omega}\boldsymbol{1}_{A}\circ Y\cdot X\,\mathrm{d}\mathrm{P}$ $=\int_{\Omega}\boldsymbol{1}_{A}\circ\pi_{\mathbb{R}^{n}}(X,Y)\cdot g\circ\pi_{E}(X,Y)\,\mathrm{d}\mathrm{P}$
$=\int_{E\times\mathbb{R}^{n}}\boldsymbol{1}_{E\times A}\cdot g\circ\pi_{E}\,\mathrm{d}\mathrm{P}_{(X,Y)}\,.$
Combining the latter two calculations with (39) we obtain
$\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot(Z_{g}\circ Y)\,\mathrm{d}\mathrm{P}=\int_{\Omega}\boldsymbol{1}_{A_{0}}\cdot X\,\mathrm{d}\mathrm{P}$
for all $A_{0}\in\sigma(Y)$. This shows that $Z_{g}\circ Y$ is a version of $\mathbb{E}(X|Y)$ and hence so is $Z$. Moreover, the map $T:\mathbb{R}^{n}\to E$ defined by
$Ty:=\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\,,\qquad\qquad y\in\mathbb{R}^{n},$
is affine linear and continuous with $Z=T\circ Y$. Since $Y$ is Gaussian random variable, we then see that $Z$ is also a Gaussian random variable. In addition, we have $\mathbb{E}Z=\mathbb{E}(\mathbb{E}(X|Y))=\mathbb{E}X$ by Theorem D.4.
To verify the formula for $\operatorname{cov}(Z)$, we first consider the case $E=\mathbb{R}$, in which we only need to determine $\operatorname{var}Z$. Since we already know $\mathbb{E}Z=\mathbb{E}X$, we then obtain
$\operatorname{var}Z=\mathbb{E}Z^{2}-(\mathbb{E}Z)^{2}=\mathbb{E}Z^{2}-(\mathbb{E}X)^{2}\,.$ (40)
Moreover, since for $y\in\mathbb{R}^{n}\setminus N$ we have $\mu_{X|Y}(y)=Ty$ we find
$\mathbb{E}Z^{2}=\mathbb{E}\bigl{(}T\circ Y)^{2}=\int_{\mathbb{R}^{n}}(Ty)^{2}\,\mathrm{d}\mathrm{P}_{Y}(y)=\int_{\mathbb{R}^{n}}(\mu_{X|Y}(y))^{2}\,\mathrm{d}\mathrm{P}_{Y}(y)$ (41)
by the definition of $T$ and the already established formula for $\mu_{X|Y}(y)$. Since $E=\mathbb{R}$ we further have $K_{X|Y}=\operatorname{var}(\mathrm{P}_{X|Y}(\,\cdot\,|y))$ for all $y\in\mathbb{R}^{n}\setminus N$ and combining this with (41) and (40) leads to
$K_{X|Y}=\int_{\mathbb{R}^{n}}K_{X|Y}\,\mathrm{d}\mathrm{P}_{Y}(y)$ $=\int_{\mathbb{R}^{n}}\biggl{(}\int_{\mathbb{R}}x^{2}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)-\bigl{(}\mu_{X|Y}(y)\bigr{)}^{2}\biggr{)}\,\mathrm{d}\mathrm{P}_{Y}(y)$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}}x^{2}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)\,\mathrm{d}\mathrm{P}_{Y}(y)-\mathbb{E}Z^{2}$
=×$=\int_{\mathbb{R}\times\mathbb{R}^{n}}(\pi_{E})^{2}\,\mathrm{d}\mathrm{P}_{(X,Y)}-\operatorname{var}Z-(\mathbb{E}X)^{2}$
$=\int_{\mathbb{R}}x^{2}\,\mathrm{d}\mathrm{P}_{X}-\operatorname{var}Z-(\mathbb{E}X)^{2}$
$=\operatorname{var}X-\operatorname{var}Z\,,$
where in the third to last step we used the disintegration formula (D.9) in the case of non-negative functions. In other words, we have found the desired covariance formula in the case $E=\mathbb{R}$.
To verify the formula for $\operatorname{cov}(Z)$ in the case of general $E$, we pick an $x^{\prime}\in E^{\prime}$ and define $X_{0}:=\langle x^{\prime},X\rangle_{E^{\prime},E}$ as well as
$Z_{0}:=\mathbb{E}X_{0}+\operatorname{cov}(X_{0},Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)\,.$
Using (B.3) this definition can also be expressed as
$\langle x^{\prime},Z\rangle_{E^{\prime},E}$ $=\bigl{\langle}x^{\prime},\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)\bigr{\rangle}_{E^{\prime},E}$
$=\mathbb{E}X_{0}+\bigl{(}x^{\prime}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(Y-\mathbb{E}Y)$
$=\mathbb{E}X_{0}+\operatorname{cov}(X_{0},Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)$
$=Z_{0}\,.$
Since $X_{0}$ and $Y$ are jointly Gaussian random variables, the already established case $E=\mathbb{R}$ thus shows
$\langle x^{\prime},\operatorname{cov}(Z)x^{\prime}\rangle_{E^{\prime},E}=\operatorname{var}(\langle x^{\prime},Z\rangle_{E^{\prime},E})=\operatorname{var}Z_{0}=\operatorname{var}(X_{0})-K_{X_{0}|Y}$ $=K_{X_{0},Y}K_{Y,Y}^{\dagger}K_{Y,X_{0}}$
$=\langle x^{\prime},R_{X,Y}x^{\prime}\rangle_{E^{\prime},E}\,,$
where in first step we used (B.6) and in the last step we repeated the calculations of (34). Since $\operatorname{cov}(Z)$ is an abstract covariance operator and we have seen around (38) that $R_{X,Y}$ is also an abstract covariance operator, Lemma B.2 then yields $\operatorname{cov}(Z)=R_{X,Y}$.
Finally, note that our definition of $Z$ gives $Z-\mathbb{E}Z=Z-\mathbb{E}X=\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(Y-\mathbb{E}Y)$. Moreover, $Y-\mathbb{E}Y$ is a centered Gaussian random variable and $\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}:\mathbb{R}^{n}\to E$ is bounded and linear. For $x^{\prime}\in E^{\prime}$ we then have $y^{\prime}:=x^{\prime}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\in(\mathbb{R}^{n})^{\prime}$, and this gives
$\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}=\bigl{(}x^{\prime}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(Y-\mathbb{E}Y)=\langle y^{\prime},Y-\mathbb{E}Y\rangle_{(\mathbb{R}^{n})^{\prime},\mathbb{R}^{n}}\in{\cal G}(Y)\,.$
Taking the ${L_{2}(\mathrm{P})}$-closure then yields the assertion. ∎
5.3 Towards the Proof of Theorem 3.4: More on Filtering Sequences
In this subsection we elaborate on filtering sequences and prove Proposition 3.3. We begin with the following lemma that considers the behavior of generated $\sigma$-algebras under some natural operations in $F^{\prime}$.
Lemma 5.5.
Let $F$ be a Banach space and $B\subset F^{\prime}$ be non-empty. Then we have
$\sigma(B)=\sigma(\operatorname{span}B)=\sigma(\overline{B}^{w^{*}\mathrm{\!-seq}})=\sigma(\overline{\operatorname{span}B}^{w^{*}\mathrm{\!-seq}})\,.$
Proof of Lemma 5.5: We first prove the equation $\sigma(B)=\sigma(\operatorname{span}B)$. Here, the inclusion $\subset$” is obvious. To establish the converse, we fix an $y^{\prime}\in\operatorname{span}B$. Then there exist an $n\geq 1$ as well as $\alpha_{1},\dots,\alpha_{n}\in\mathbb{R}$, and $y_{1}^{\prime},\dots,y_{n}^{\prime}\in B$ with $y^{\prime}=\alpha_{1}y_{1}^{\prime}+\dots+\alpha_{n}y_{n}^{\prime}$. Clearly, each $y_{i}^{\prime}:F\to\mathbb{R}$ is $(\sigma(B),\mathfrak{B})$-measurable, and hence $y^{\prime}$ is also $(\sigma(B),\mathfrak{B})$-measurable. This gives $\sigma(y^{\prime})\subset\sigma(B)$, and since $y^{\prime}\in\operatorname{span}B$ was chosen arbitrarily, we find $\sigma(\operatorname{span}B)\subset\sigma(B)$.
For the proof of $\sigma(B)=\sigma(\overline{B}^{w^{*}\mathrm{\!-seq}})$, we only need to show $\supset$”. To this end, we fix a $y^{\prime}\in\overline{B}^{w^{*}\mathrm{\!-seq}}$. Then we find a sequence $(y_{i}^{\prime})\subset B$ with $y_{i}^{\prime}\to y^{\prime}$ in the $w^{*}$-topology. This gives
$\langle y^{\prime},y\rangle_{F^{\prime},F}=\lim_{i\to\infty}\langle y_{i}^{\prime},y\rangle_{F^{\prime},F}\,,\qquad\qquad y\in F.$
Since each $y_{i}^{\prime}:F\to\mathbb{R}$ is $(\sigma(B),\mathfrak{B})$-measurable, we then conclude that their pointwise limit $y^{\prime}$ is also $(\sigma(B),\mathfrak{B})$-measurable. This shows $\sigma(y^{\prime})\subset\sigma(B)$, and again, we can then conclude that $\sigma(\overline{B}^{w^{*}\mathrm{\!-seq}})\subset\sigma(B)$.
The third equation is a direct consequence of the first two equations. ∎
Our next goal is to investigate the $\sigma$-algebra generated by some bounded linear $A:F\to\mathbb{R}^{n}$. To this end, let $e_{1},\dots,e_{n}$ be the standard ONB of $\mathbb{R}^{n}$. We then write
$\mathfrak{C}_{A}:=\bigl{\{}a_{i}:F\to\mathbb{R}\,\bigl{|}\,a_{i}:=\langle e_{i},A\rangle,i=1,\dots,n\bigr{\}}$
for the set of its coordinate functions $a_{1},\dots,a_{n}$. Clearly, we have $\mathfrak{C}_{A}\subset F^{\prime}$. To relate $\sigma(A)$ with $\mathfrak{C}_{A}$ we further need to recall that the annihilator of a non-empty $B\subset F$ is defined by
$B^{\perp\perp}$ :=$:=\bigl{\{}y^{\prime}\in F^{\prime}:\langle y^{\prime},y\rangle_{F^{\prime},F}=0\mbox{ for all }y\in B\bigr{\}}\,.$
Analogously, the annihilator of a non-empty set $B\subset F^{\prime}$ is defined by
$B_{\perp\perp}:=\bigl{\{}y\in F:\langle y^{\prime},y\rangle_{F^{\prime},F}=0\mbox{ for all }y^{\prime}\in B\bigr{\}}\,.$
With these preparations, our next intermediate result reads as follows.
Lemma 5.6.
Let $F$ be a Banach space and $A:F\to\mathbb{R}^{n}$ be a bounded linear operator. Then the following statements hold true:
  1. i)
    We have $\sigma(A)=\sigma(\mathfrak{C}_{A})=\sigma(\operatorname{span}\mathfrak{C}_{A})$.
  2. ii)
    We have $\operatorname{span}\mathfrak{C}_{A}=\operatorname{ran}A^{\prime}=(\ker A)^{\perp\perp}$.
Proof of Lemma 5.6: i). We first note that in view of Lemma 5.5 we only need to show $\sigma(A)=\sigma(\mathfrak{C}_{A})$. To establish this identity, let $a_{1}^{\prime},\dots,a_{n}^{\prime}:F\to\mathbb{R}$ be the coordinate functions of $A$, that is $\mathfrak{C}_{A}=\{a_{1}^{\prime},\dots,a_{n}^{\prime}\}$. Then, [27, Theorem 1.90] shows that for a given $\sigma$-algebra ${\cal A}^{\prime}$ on $F$, the map $A:F\to\mathbb{R}^{n}$ is $({\cal A}^{\prime},\mathfrak{B}^{n})$-measurable if and only if all $a_{1}^{\prime},\dots,a_{n}^{\prime}$ are $({\cal A}^{\prime},\mathfrak{B})$-measurable.
If we now consider ${\cal A}^{\prime}:=\sigma(A)$, then $A$ is $({\cal A}^{\prime},\mathfrak{B}^{n})$-measurable, and hence we conclude that $\sigma(a_{i}^{\prime})\subset{\cal A}^{\prime}=\sigma(A)$ for all $i=1,\dots,n$. This gives the inclusion $\sigma(a_{1}^{\prime},\dots,a_{n}^{\prime})\subset\sigma(A)$. Conversely, for ${\cal A}^{\prime}:=\sigma(a_{1}^{\prime},\dots,a_{n}^{\prime})$, we know that all $a_{1}^{\prime},\dots,a_{n}^{\prime}$ are $({\cal A}^{\prime},\mathfrak{B})$-measurable, and hence $A$ is ${\cal A}^{\prime}$-measurable, that is $\sigma(A)\subset\sigma(a_{1}^{\prime},\dots,a_{n}^{\prime})$.
ii). Let $e_{1},\dots,e_{n}$ be the standard ONB of $\mathbb{R}^{n}$. Then $A$ is given by $A=\sum_{i=1}^{n}a_{i}^{\prime}\otimes e_{i}$, and therefore (B.14) shows that its adjoint $A^{\prime}$ is given by
$A^{\prime}t=\biggl{(}\sum_{i=1}^{n}e_{i}\otimes a_{i}^{\prime}\biggr{)}(t)=\sum_{i=1}^{n}\langle e_{i},t\rangle\cdot a_{i}^{\prime}=\sum_{i=1}^{n}t_{i}a_{i}^{\prime}\,,\qquad\qquad t=(t_{1},\dots,t_{n})\in\mathbb{R}^{n},$
where we use the canonical identification of $\mathbb{R}^{n}$ with both $(\mathbb{R}^{n})^{\prime}$ and $(\mathbb{R}^{n})^{\prime\prime}$. This shows $\operatorname{span}\mathfrak{C}_{A}=\operatorname{ran}A^{\prime}$.
For the proof of the second identity, we first note that the linearity and continuity of $A$ gives
$(\ker A)^{\perp\perp}=\bigl{(}(\operatorname{ran}A^{\prime})_{\perp\perp}\bigr{)}^{\perp\perp}=\overline{\operatorname{ran}A^{\prime}}^{w^{*}}\,,$ (42)
where the first identity can be found in e.g. [32, Lemma 3.1.16] or [11, Proposition VI.1.8], and the second identity can be found in e.g. [32, Proposition 2.6.6]. Moreover, since the subspace $\operatorname{ran}A^{\prime}$ of $F^{\prime}$ is finite dimensional, it is $w^{*}$-closed, see e.g. [32, Corollary 2.2.32] or [37, Theorem 1.21]. Combining the latter with (42) yields $(\ker A)^{\perp\perp}=\operatorname{ran}A^{\prime}$. ∎
The next proposition provides a general sufficient condition for filtering sequences. It will be at the core of the proof of Proposition 3.3.
Proposition 5.7.
Let $F$ be a separable Banach space and $(A_{n})$ be a sequence of bounded linear operators $A_{n}:F\to\mathbb{R}^{n}$. Then, $(A_{n})$ is a filtering sequence for $F$ if we have both $\operatorname{span}\mathfrak{C}_{A_{n}}\subset\operatorname{span}\mathfrak{C}_{A_{n+1}}$ for all $n\geq 1$ and
$\overline{\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}}^{w^{*}\mathrm{\!-seq}}=F^{\prime}\,.$ (43)
Proof of Proposition 5.7: By Lemma 5.6 we find $\sigma(A_{n})=\sigma(\operatorname{span}\mathfrak{C}_{A_{n}})\subset\sigma(\operatorname{span}\mathfrak{C}_{A_{n+1}})=\sigma(A_{n+1})$. To verify the condition $\sigma(F^{\prime})=\sigma\bigl{(}A_{n}:n\geq 1\bigr{)}$, we write $F_{n}^{\prime}:=\operatorname{span}\mathfrak{C}_{A_{n}}$ and
$F_{\infty}^{\prime}:=\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\,.$
Since $F_{n}^{\prime}\subset F_{n+1}^{\prime}$ for all $n\geq 1$, we quickly see that $F_{\infty}^{\prime}$ is a subspace of $F^{\prime}$. Using our assumption and Lemma 5.5 we further have
$\sigma(F^{\prime})=\sigma(\overline{F_{\infty}^{\prime}}^{w^{*}\mathrm{\!-seq}})=\sigma(F_{\infty}^{\prime})=\sigma\biggl{(}\bigcup_{y^{\prime}\in F_{\infty}^{\prime}}\sigma(y^{\prime})\biggr{)}=\sigma\biggl{(}\bigcup_{n\geq 1}\bigcup_{y^{\prime}\in F_{n}^{\prime}}\sigma(y^{\prime})\biggr{)}$ $\subset\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(F_{n}^{\prime})\biggr{)}$
$=\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(A_{n})\biggr{)}\,,$
where in the last step we used the already observed $\sigma(F_{n}^{\prime})=\sigma(\operatorname{span}\mathfrak{C}_{A_{n}})=\sigma(A_{n})$. Hence we have found $\sigma(F^{\prime})\subset\sigma\bigl{(}A_{n}:n\geq 1\bigr{)}$. The converse inclusion is trivial. ∎
Proof of Proposition 3.3: We first note that our construction gives $\mathfrak{C}_{A_{n}}=\{y_{1}^{\prime},\dots,y_{n}^{\prime}\}$. This identity implies $\operatorname{span}\mathfrak{C}_{A_{n}}\subset\operatorname{span}\mathfrak{C}_{A_{n+1}}$ for all $n\geq 1$ as well as
$\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}=\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\,.$
Applying Proposition 5.7 then yields the assertion. ∎
Lemma 5.8.
Let $F$ be a separable Banach space and $(A_{n})$ be a filtering sequence for $F$. Moreover, let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $Y:\Omega\to F$ be $\sigma(F^{\prime})$- measurable, and $Y_{n}:=A_{n}\circ Y$. Then $(\sigma(Y_{n}))$ is a filtration in ${\cal A}$ and we have
$\sigma(Y)=\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(Y_{n})\biggr{)}\,.$
Proof of Lemma 5.8: We first note that we have
$\sigma(Y_{n})=\sigma(A_{n}\circ Y)=(A_{n}\circ Y)^{-1}(\mathfrak{B}^{n})=Y^{-1}(A_{n}^{-1}(\mathfrak{B}^{n}))=Y^{-1}(\sigma(A_{n}))\,,$
where in the second and the last step we used an elementary formula for computing the $\sigma$-algebra generated by a single map, see e.g. [27, Theorem 1.78]. Since $(\sigma(A_{n}))$ is a filtration, so is $(\sigma(Y_{n}))$. To establish the formula for $\sigma(Y)$, we write ${\cal E}:=\bigcup_{n\geq 1}\sigma(A_{n})$. By assumption we know that $\sigma({\cal E})=\sigma(F^{\prime})$, and hence we obtain
$\sigma(Y)=Y^{-1}(\sigma(F^{\prime}))=Y^{-1}(\sigma({\cal E}))=\sigma(Y^{-1}({\cal E}))=\sigma\biggl{(}\bigcup_{n\geq 1}Y^{-1}(\sigma(A_{n}))\biggr{)}=\sigma\biggl{(}\bigcup_{n\geq 1}\sigma(Y_{n})\biggr{)}\,,$
where in the third step we used another well-known fact on generated $\sigma$-algebras, see e.g. [27, Theorem 1.81] and its proof. ∎
5.4 Towards the Proof of Theorem 3.4: A Banach Space Martingale Theorem for Regular Conditional Probabilities
Let $E$ be a separable Banach space. Then a set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is called uniformly tight, if for all $\varepsilon>0$ there exists a compact $K\subset E$ with $\mu(E\setminus K)\leq\varepsilon$ for all $\mu\in{\cal M}$. Recall that given a probability measure $\mu$ on $(E,\mathfrak{B}(E))$, Ulam’s theorem, see e.g. [4, Lemma 26.2] or [46, Theorem 3.1 in Chapter I.3.2], shows that $\mu$ a Radon measure in the sense of inner regularity, that is
$\mu(B)=\sup\{\mu(K):K\subset B\mbox{ and $K$ is compact}\}$ (44)
for all $B\in\mathfrak{B}(E)$. In particular, $\{\mu\}$ is (uniformly) tight. Here we emphasize that we use the term “Radon measure” in the sense of [46, p. 28] and [6, p. 97], as we will need this definition also in the proof of Theorem C.6 when applying results from [46] and [6]. Consequently, every finite set ${\cal M}$ of probability measures on $(E,\mathfrak{B}(E))$ is uniformly tight and more generally, the finite union of uniformly tight sets is also uniformly tight.
The following lemma, which provides a simple characterization of uniform tightness, is taken from [8, Example 8.6.5]. In view of its formulation we note that [8, Example 8.6.5] additionally lists the condition $\mu(\{V=\infty\})=0$ for all $\mu\in{\cal M}$. However, this condition is implied by the finiteness of the integrals in ii), and hence we omitted it.
Lemma 5.9.
Let $E$ be a separable Banach space and ${\cal M}$ be a set of probability measures on $(E,\mathfrak{B}(E))$. Then the following statements are equivalent:
  1. i)
    ${\cal M}$ is uniformly tight.
  2. ii)
    There exists a $\mathfrak{B}(E)$-measurable $V:E\to[0,\infty]$ such that $\{V\leq c\}$ is compact for all $c\in[0,\infty)$ and
    $\sup_{\mu\in{\cal M}}\int_{E}V\,\mathrm{d}\mu<\infty\,.$
Theorem 5.10.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be separable Banach spaces, and $X:\Omega\to E$ and $Y:\Omega\to F$ be random variables. Moreover, let $(A_{n})$ be a filtering sequence for $F$ and
$Y_{n}:=A_{n}\circ Y\,,\qquad\qquad n\geq 1.$
Then for all versions $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times\mathbb{R}^{n}\to[0,1]$ of the regular conditional probabilities of $X$ given $Y_{n}$ there exists an $N\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N)=0$ such that
$\bigl{\{}\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)):n\geq 1\bigr{\}}$
is uniformly tight for all $y\in F\setminus N$.
In addition, for all Banach spaces $G$ and all $\mathrm{P}_{X}$-Bochner integrable $g:E\to G$ there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have $g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),G)}$ and
$g\in{{\cal L}_{1}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)),G)}\,,\qquad\qquad n\geq 1,$ (45)
as well as
$\int_{E}g(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\to\int_{E}g(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)\,,$ (46)
where the convergence is with respect to $\|\cdot\|_{G}$.
Proof of Theorem 5.10: Let us write ${\cal A}_{n}:=\sigma(A_{n})\subset\sigma(F^{\prime})$ and ${\cal A}_{\infty}:=\sigma(F^{\prime})$. Moreover, we define $A_{\infty}:=\operatorname{id}_{F}$ and $Y_{\infty}:=Y$, which gives $Y_{\infty}=A_{\infty}\circ Y$. Since $(A_{n})$ is a filtering sequence we then have
$\sigma(A_{\infty})={\cal A}_{\infty}=\sigma\bigl{(}A_{n}:n\geq 1\bigr{)}\,.$
In addition, we denote the restriction of $\mathrm{P}_{(X,Y)}$ onto $\mathfrak{B}(E)\otimes{\cal A}_{n}$ by $\mathrm{Q}^{(n)}$, that is
$\mathrm{Q}^{(n)}:=\bigl{(}\mathrm{P}_{(X,Y)}\bigr{)}_{|\mathfrak{B}(E)\otimes{\cal A}_{n}}$
for all $n\in\mathbb{N}\cup\{\infty\}$. For $n=\infty$, this gives $\mathrm{Q}^{(\infty)}=\mathrm{P}_{(X,Y)}$ and for $n\in\mathbb{N}$, the identity $Y_{n}=A_{n}\circ Y$ implies
$\mathrm{P}_{(X,Y_{n})}=\bigl{(}\mathrm{P}_{(X,Y)}\bigr{)}_{(\operatorname{id}_{E},A_{n})}=\mathrm{Q}^{(n)}_{(\operatorname{id}_{E},A_{n})}\,.$
With the help of Theorem D.1 we now fix a regular conditional probability $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ and write $\mathrm{Q}^{(\infty)}(\,\cdot\,|\,\cdot\,):=\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$. Moreover, for $n\in\mathbb{N}$, we analogously fix a regular conditional probability
$\mathrm{Q}^{(n)}(\,\cdot\,|\,\cdot\,):\mathfrak{B}(E)\times F\to[0,1]$
of $\mathrm{Q}^{(n)}$. Theorem D.2 applied to $\Phi:=\operatorname{id}_{E}$ and $\Psi:=A_{n}$ then gives an $N^{(1)}_{n}\in{\cal A}_{n}$ with $\mathrm{Q}^{(n)}_{F}(N^{(1)}_{n}))=0$ such that for all $y\in F\setminus N^{(1)}_{n}$ and all $B\in\mathfrak{B}(E)$ we have
$\mathrm{Q}^{(n)}(B|y)=\mathrm{P}_{X|Y_{n}}\bigl{(}B|A_{n}(y)\bigr{)}\,.$ (47)
Let us define
$N_{\mathrm{trafo}}:=\bigcup_{n\geq 1}N^{(1)}_{n}\,.$
Since for all $n\geq 1$ we have $N^{(1)}_{n}\in{\cal A}_{n}\subset{\cal A}_{\infty}=\sigma(F^{\prime})$ with
$\mathrm{P}_{Y}(N_{n}^{(1)})=\mathrm{P}_{(X,Y)}(E\times N_{n}^{(1)})=\mathrm{Q}^{(n)}(E\times N_{n}^{(1)})=\mathrm{Q}_{F}^{(n)}(N_{n}^{(1)})=0\,,$ (48)
we then find $N_{\mathrm{trafo}}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{\mathrm{trafo}})=0$. In addition, (47) holds for all $y\in F\setminus N_{\mathrm{trafo}}$ and all $n\in\mathbb{N}\cup\{\infty\}$ by our construction. Finally, our construction also ensures $\mathrm{Q}^{(n)}_{E}=\mathrm{P}_{X}$ for all $n\in\mathbb{N}\cup\{\infty\}$.
Our first goal is to establish (45) and (46). To this end, we fix a Banach space $G$ and a $\mathrm{P}_{X}$-Bochner integrable $g:E\to G$. Our construction then ensures
$\int_{E\times F}\|g\circ\pi_{E}\|_{G}\,\mathrm{d}\mathrm{Q}^{(n)}=\int_{E}\|g\|_{G}\,\mathrm{d}\mathrm{Q}^{(n)}_{E}=\int_{E}\|g\|_{G}\,\mathrm{d}\mathrm{P}_{X}<\infty\,,\qquad\qquad n\in\mathbb{N}\cup\{\infty\}.$ (49)
In other words, we have $g\circ\pi_{E}\in{{\cal L}_{1}(\mathrm{Q}^{(n)},G)}$ for all $n\in\mathbb{N}\cup\{\infty\}$. In the next step we will apply the martingale convergence Theorem D.5. To this end, we define the $\sigma$-algebras
${\cal C}_{n}:=\pi_{F}^{-1}({\cal A}_{n})=\{E\times A:A\in{\cal A}_{n}\}\,,\qquad\qquad n\geq 1$
and we further write
${\cal C}_{\infty}:=\pi_{F}^{-1}(\sigma(F^{\prime}))=\{E\times A:A\in\sigma(F^{\prime})\}=\{E\times A:A\in{\cal A}_{\infty}\}\,.$
Since $({\cal A}_{n})$ is a filtration with ${\cal A}_{\infty}=\sigma({\cal A}_{n}:n\geq 1)$, we quickly check that $({\cal C}_{n})$ is a filtration with ${\cal C}_{\infty}=\sigma({\cal C}_{n}:n\geq 1)$. For every fixed choice of versions $g_{n}$ of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{n})$ and $g_{\infty}$ of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{\infty})$ Theorem D.5 applied to the probability space $(E\times F,\mathfrak{B}(E)\otimes\sigma(F^{\prime}),\mathrm{P}_{(X,Y)})$ then gives an $N^{\prime}\in{\cal C}_{\infty}$ with $\mathrm{P}_{(X,Y)}(N^{\prime})=0$ and
$\|g_{n}(x,y)-g_{\infty}(x,y)\|_{G}\to 0\,,\qquad\qquad(x,y)\in(E\times F)\setminus N^{\prime}\,.$ (50)
Moreover, note that $N^{\prime}$ is of the form $N^{\prime}=E\times N_{\mathrm{mart}}$ with $N_{\mathrm{mart}}\in\sigma(F^{\prime})$ since $N^{\prime}\in{\cal C}_{\infty}$. In addition, we have $\mathrm{P}_{Y}(N_{\mathrm{mart}})=\mathrm{P}_{(X,Y)}(N^{\prime})=0$.
Let us now translate this convergence to regular conditional probabilities. To this end, we apply Theorem D.6 to $g\circ\pi_{E}$ and the measures $\mathrm{Q}^{(n)}$ defined on $\mathfrak{B}(E)\otimes{\cal A}_{n}$: Because of (49) we then obtain an $N^{(2)}_{n}\in{\cal A}_{n}$ with $\mathrm{Q}^{(n)}_{F}(N^{(2)}_{n})=0$ and
$g=g\circ\pi_{E}(\,\cdot\,,y)\in{{\cal L}_{1}(\mathrm{Q}^{(n)}(\,\cdot\,|y),G)}\,,\qquad\qquad y\in F\setminus N^{(2)}_{n}$
such that $Z_{n}:F\to G$ defined by
$Z_{n}(y):=\int_{E}\boldsymbol{1}_{F\setminus N^{(2)}_{n}}(y)\cdot(g\circ\pi_{E})(x,y)\,\mathrm{Q}^{(n)}(\mathrm{d}x|y)=\int_{E}\boldsymbol{1}_{F\setminus N^{(2)}_{n}}(y)\cdot g(x)\,\mathrm{Q}^{(n)}(\mathrm{d}x|y)\,,\qquad\qquad y\in F$
is strongly measurable, the disintegration formula (D.9) holds true, and $Z_{n}\circ\pi_{F}:E\times F\to G$ is a version of $\mathbb{E}_{\mathrm{Q}^{(n)}}(g\circ\pi_{E}|{\cal C}_{n})$, that is, we have
$\int_{E\times F}\boldsymbol{1}_{C}\cdot(Z_{n}\circ\pi_{F})\,\mathrm{d}\mathrm{Q}^{(n)}=\int_{E\times F}\boldsymbol{1}_{C}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{Q}^{(n)}$ (51)
for all $C\in{\cal C}_{n}$. Our next intermediate goal is to verify that $Z_{n}\circ\pi_{F}$ is also a version of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{n})$. Here, the strong ${\cal C}_{n}$-measurability is obviously satisfied since $Z_{n}\circ\pi_{F}$ is a version of $\mathbb{E}_{\mathrm{Q}^{(n)}}(g\circ\pi_{E}|{\cal C}_{n})$. Moreover, for $C\in{\cal C}_{n}$ we have
$\int_{E\times F}\boldsymbol{1}_{C}\cdot(Z_{n}\circ\pi_{F})\,\mathrm{d}\mathrm{P}_{(X,Y)}=\int_{E\times F}\boldsymbol{1}_{C}\cdot(Z_{n}\circ\pi_{F})\,\mathrm{d}\mathrm{Q}^{(n)}\,,$ (52)
where we used that $\boldsymbol{1}_{C}\cdot(Z_{n}\circ\pi_{F})$ is strongly ${\cal C}_{n}$-measurable and thus also strong $\mathfrak{B}(E)\otimes{\cal A}_{n}$-measurable and that $\mathrm{Q}^{(n)}$ is the restriction of $\mathrm{P}_{(X,Y)}$ onto $\mathfrak{B}(E)\otimes{\cal A}_{n}$. In addition, $g\circ\pi_{E}$ is strongly $\{B\times F:B\in\mathfrak{B}(E)\}$-measurable and thus also strongly $\mathfrak{B}(E)\otimes{\cal A}_{n}$-measurable. Consequently, $\boldsymbol{1}_{C}\cdot(g\circ\pi_{E})$ is also strongly $\mathfrak{B}(E)\otimes{\cal A}_{n}$-measurable, and since $\mathrm{Q}^{(n)}$ is the restriction of $\mathrm{P}_{(X,Y)}$ onto $\mathfrak{B}(E)\otimes{\cal A}_{n}$ we find
$\int_{E\times F}\boldsymbol{1}_{C}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{Q}^{(n)}=\int_{E\times F}\boldsymbol{1}_{C}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{P}_{(X,Y)}\,.$ (53)
Combining (51) with (52) and (53), we now find
$\int_{E\times F}\boldsymbol{1}_{C}\cdot(Z_{n}\circ\pi_{F})\,\mathrm{d}\mathrm{P}_{(X,Y)}=\int_{E\times F}\boldsymbol{1}_{C}\cdot(g\circ\pi_{E})\,\mathrm{d}\mathrm{P}_{(X,Y)}$
for all $C\in{\cal C}_{n}$. In other words, $Z_{n}\circ\pi_{F}$ is indeed a version of $\mathbb{E}_{\mathrm{P}_{(X,Y)}}(g\circ\pi_{E}|{\cal C}_{n})$. Moreover, these versions are independent of $x\in E$, and for all $y\in F\setminus N_{\mathrm{mart}}$ the martingale convergence (50) thus shows
$Z_{n}(y)\to Z_{\infty}(y)\,.$
Let us define $N_{g}:=N_{\mathrm{trafo}}\cup N_{\mathrm{mart}}\cup N^{(2)}_{\infty}\cup\bigcup_{n\geq 1}N^{(2)}_{n}$. Clearly, this definition gives $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$. For $y\in F\setminus N_{g}$ our considerations further give
$g\in{{\cal L}_{1}(\mathrm{Q}^{(n)}(\,\cdot\,|y),G)}={{\cal L}_{1}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)),G)}$
for all $n\in\mathbb{N}\cup\{\infty\}$ as well as
$\int_{E}g(x)\,\mathrm{P}_{X|Y_{n}}\bigl{(}\mathrm{d}x|A_{n}(y)\bigr{)}=\int_{E}g(x)\,\mathrm{Q}^{(n)}(\mathrm{d}x|y)=Z_{n}(y)\to Z_{\infty}(y)$ $=\int_{E}g(x)\,\mathrm{Q}^{(\infty)}(\mathrm{d}x|y)$
$=\int_{E}g(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
for $n\to\infty$, where the convergence is with respect to the norm $\|\,\cdot\,\|_{G}$. In other words we have shown both (45) and (46).
Let us finally establish the uniform tightness. To this end, we first note that $\{\mathrm{P}_{X}\}$ is uniformly tight as discussed around (44), and hence Lemma 5.9 gives a $\mathfrak{B}(E)$-measurable $V:E\to[0,\infty]$ such that $\{V\leq c\}$ is compact for all $c\in[0,\infty)$ and $V\in{{\cal L}_{1}(\mathrm{P}_{X})}$. Consequently, $N_{\mathrm{tight}}:=\{V=\infty\}\in\mathfrak{B}(E)$ satisfies $\mathrm{P}_{X}(N_{\mathrm{tight}})=0$. We write $G:=\mathbb{R}$ and $g:=\boldsymbol{1}_{E\setminus N_{\mathrm{tight}}}V$, where we use the convention $0\cdot\infty:=0$. Then we have $g\in{{\cal L}_{1}(\mathrm{P}_{X},\mathbb{R})}$, and therefore we already know that there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have $g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),\mathbb{R})}$ and both (45) and (46) hold true. Consequently, there exists an $n_{0}\geq 1$ such that
$\int_{E}g(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\leq 1+\int_{E}g(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)$
for all $n\geq n_{0}$ and all $y\in F\setminus N_{g}$. By applying (45) we conclude that
$\sup_{n\geq 1}\int_{E}g(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))<\infty$ (54)
for all $y\in F\setminus N_{g}$. Now note that $\{g\leq c\}$ may not be compact, and therefore we cannot apply Lemma 5.9 directly. For this reason, we translate (54) back to the function $V$. To this end, we note that we have
$0=\mathrm{P}_{X}(N_{\mathrm{tight}})=\mathrm{P}_{(X,Y)}(N_{\mathrm{tight}}\times F)=\mathrm{Q}^{(n)}(N_{\mathrm{tight}}\times F)=\int_{F}\mathrm{Q}^{(n)}(N_{\mathrm{tight}}|y)\,\mathrm{d}\mathrm{Q}^{(n)}_{F}(y)\,.$
Consequently, there exists an $N_{n}^{(3)}\in{\cal A}_{n}$ with $\mathrm{Q}^{(n)}_{F}(N_{n}^{(3)})=0$ and
$\mathrm{Q}^{(n)}(N_{\mathrm{tight}}|y)=0\,,\qquad\qquad y\in F\setminus N_{n}^{(3)}.$
Repeating the calculation (48) then shows $\mathrm{P}_{Y}(N_{n}^{(3)})=0$. We define $N:=N_{g}\cup\bigcup_{n\geq 1}(N_{n}^{(1)}\cup N_{n}^{(3)})$. This gives $N\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N)=0$ and for $y\in F\setminus N$ we have both (54) and
$\mathrm{P}_{X|Y_{n}}(N_{\mathrm{tight}}|A_{n}(y))=\mathrm{Q}^{(n)}(N_{\mathrm{tight}}|y)=0\,.$
Moreover, we have $N_{\mathrm{tight}}=\{V=\infty\}=\{V\neq g\}$ by the definition of $g$. The previous identity thus implies
$\int_{E}V(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))=\int_{E}g(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))$
for all $y\in F\setminus N$. Inserting this into (54) then gives the uniform tightness by Lemma 5.9. ∎
5.5 Proof of Theorem 3.4
Proof of Theorem 3.4: In the following, we sequentially prove all parts of Theorem 3.4. To this end, we construct for each part x) an $N^{x)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{x)})=0$ such that for all $y\in F\setminus N^{x)}$ the claims of the parts i) to x) hold true. The desired $\mathrm{P}_{Y}$-zero set $N$ is then given by $N:=N^{viii)}$.
i). Since $(\operatorname{id}_{E},A_{n}):E\times F\to E\times\mathbb{R}^{n}$ is a bounded and linear operator we quickly see that $(X,Y_{n})$ is a Gaussian random variable. For each $n\geq 1$, Theorem 3.1 then gives an $N_{n}^{(1)}\in\mathfrak{B}^{n}$ with $\mathrm{P}_{Y_{n}}(N_{n}^{(1)})=0$ such for all $y_{n}\in\mathbb{R}^{n}\setminus N_{n}^{(1)}$ we have
$\mathrm{P}_{X|Y_{n}}(\,\cdot\,|y_{n})={\cal N}\bigl{(}\mu_{X|Y_{n}}(y_{n}),C_{X|Y_{n}}\bigr{)}\,,$ (55)
where $\mu_{X|Y_{n}}(y_{n})\in E$ and $C_{X|Y_{n}}:E^{\prime}\to E$ are described in Theorem 3.1. We define
$N^{i)}:=\bigcup_{n\geq 1}A_{n}^{-1}(N_{n}^{(1)})\,.$
Note that we have $A_{n}^{-1}(N_{n}^{(1)})\in\sigma(A_{n})\subset\sigma(F^{\prime})$, and $Y_{n}=A_{n}\circ Y$ shows $\mathrm{P}_{Y}(A_{n}^{-1}(N_{n}^{(1)}))=\mathrm{P}_{Y_{n}}(N_{n}^{(1)})=0$. Consequently, we have $N^{i)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{i)})=0$. Moreover, for $y\in F\setminus N^{i)}$ we have $y_{n}:=A_{n}(y)\in\mathbb{R}^{n}\setminus N_{n}^{(1)}$ and therefore, the claims follow from (55).
ii). We first consider the case $E=\mathbb{R}$. Here, we fix an enumeration $(t_{m})_{m\geq 3}$ of $\mathbb{Q}$ and consider the functions $g_{m}:\mathbb{R}\to\mathbb{C}$ defined by
$g_{m}(x):=\exp(\mathrm{i}t_{m}x)\,,\qquad\qquad x\in\mathbb{R}$
as well as the two functions $g_{1},g_{2}:\mathbb{R}\to\mathbb{R}$ defined by $g_{i}(x):=x^{i}$ for $i=1,2$. Let $G:=\mathbb{C}$ with $\mathbb{C}$ being interpreted as the $\mathbb{R}$-vector space $\mathbb{R}^{2}$. Since for all $m\geq 3$ the function $g_{m}$ is bounded and continuous, it is $\mathrm{P}_{X}$-Bochner integrable. Therefore, Theorem 5.10 shows that for each $m\geq 3$ there exists an $N_{m}^{(2)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{m}^{(2)})=0$ such that for all $y\in F\setminus N_{m}^{(2)}$ we have
$\varphi_{\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))}(t_{m})=\int_{E}g_{m}(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\to\int_{E}g_{m}(x)\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)=\varphi_{\mathrm{P}_{X|Y}(\,\cdot\,|y)}(t_{m})\,.$
In addition, for $i=1,2$ we have
$\int_{\mathbb{R}}|g_{i}|\,\mathrm{d}\mathrm{P}_{X}=\int_{\mathbb{R}}|x|^{i}\,\mathrm{d}\mathrm{P}_{X}(x)<\infty$
since $\mathrm{P}_{X}$ is a Gaussian measure on $\mathbb{R}$. Consequently, $g_{1}$ and $g_{2}$ are also $\mathrm{P}_{X}$-Bochner integrable, and therefore Theorem 5.10 shows that there exist $N_{1}^{(2)},N_{2}^{(2)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{i}^{(2)})=0$ such that for $i=1,2$ and all $y\in F\setminus N_{i}^{(2)}$ we have
$\mu_{n,y}^{(i)}:=\int_{E}x^{i}\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\to\int_{E}x^{i}\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)=:\mu_{\infty,y}^{(i)}\,.$
In other words, the first two raw moments $\mu_{n,y}^{(i)}$ of $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))$ converge to the first two raw moments $\mu_{\infty,y}^{(i)}$ of $\mathrm{P}_{X|Y}(\,\cdot\,|y)$. Now, the first raw moments are, of course, the expectations. Moreover, the variances can be computed by the first and second raw moments, and hence we obtain
$\sigma_{n,y}^{2}:=\operatorname{var}\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))$ $\to\operatorname{var}\mathrm{P}_{X|Y}(\,\cdot\,|y)=:\sigma_{\infty,y}^{2}\,.$
Let us define $N^{\prime}:=N^{i)}\cup\bigcup_{m\geq 1}N_{m}^{(2)}$. Obviously, this implies $N^{\prime}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{\prime})=0$ and
$\varphi_{\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))}(t)\to\varphi_{\mathrm{P}_{X|Y}(\,\cdot\,|y)}(t)$
for all $t\in\mathbb{Q}$ and $y\in F\setminus N^{\prime}$ as well as
$\varphi_{\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))}(t)=\mathrm{e}^{\mathrm{i}t\mu_{n,y}^{(1)}}\cdot\exp\biggl{(}-\frac{t^{2}\sigma_{n,y}^{2}}{2}\biggr{)}\to\mathrm{e}^{\mathrm{i}t\mu_{\infty,y}^{(1)}}\cdot\exp\biggl{(}-\frac{t^{2}\sigma_{\infty,y}^{2}}{2}\biggr{)}=\varphi_{{\cal N}(\mu_{\infty,y},\sigma_{\infty,y}^{2})}(t)$
for all $t\in\mathbb{R}$ and $y\in F\setminus N^{\prime}$. Consequently, for each fixed $y\in F\setminus N^{\prime}$, we have
$\varphi_{\mathrm{P}_{X|Y}(\,\cdot\,|y)}(t)=\varphi_{{\cal N}(\mu_{\infty,y},\sigma_{\infty,y}^{2})}(t)$ (56)
for all $t\in\mathbb{Q}$, and since both characteristic functions are continuous, we conclude that (56) actually holds for all $t\in\mathbb{R}$. This shows that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is a Gaussian measure.
Let us now consider the case of a general separable Banach space $E$. Here we fix a countable $\mathfrak{D}_{E^{\prime}}\subset B_{E^{\prime}}$ that is $w^{*}$-dense in $B_{E^{\prime}}$ and an enumeration $\mathfrak{D}_{E^{\prime}}=\{x_{i}^{\prime}:i\geq 1\}$. We write $X_{i}:=x_{i}^{\prime}\circ X$. Our considerations in the case $E=\mathbb{R}$ then show that for each $i\geq 1$ there exists an $N_{i}^{\prime}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{i}^{\prime})=0$ such that $\mathrm{P}_{X_{i}|Y}(\,\cdot\,|y)$ is a Gaussian measure for all $y\in F\setminus N_{i}^{\prime}$.
As in the proof of Theorem 3.1 we now translate this result to $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ with the help of Theorem D.2. To this end, we set $T_{0}:=E$ and $T_{1}:=\mathbb{R}$, as well as $U_{0}:=U_{1}:=F$ and ${\cal A}_{0}:={\cal A}_{1}:=\sigma(F^{\prime})$. Moreover, we consider the maps $\Phi:=x_{i}^{\prime}$ and $\Psi:=\operatorname{id}_{F}$ as well as the measures $\mathrm{P}^{(0)}:=\mathrm{P}_{(X,Y)}$ and $\mathrm{P}^{(1)}:=\mathrm{P}_{(X_{i},Y)}=\mathrm{P}^{(1)}_{(\Phi,\Psi)}$. By Theorem D.2 there then exists an $N_{i}^{\prime\prime}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{i}^{\prime\prime})=0$ such that for all $y\in F\setminus N_{i}^{\prime\prime}$ we have
$\mathrm{P}_{X|Y}\bigl{(}(x_{i}^{\prime})^{-1}(B)|y\bigr{)}=\mathrm{P}_{X_{i}|Y}(B|y)\,,\qquad\qquad B\in\mathfrak{B}.$
Consequently, we obtain
$\bigl{(}\mathrm{P}_{X|Y}(\,\cdot\,|y)\bigr{)}_{x^{\prime}_{i}}=\mathrm{P}_{X_{i}|Y}(\,\cdot\,|y)$
for all $y\in F\setminus N_{i}^{\prime\prime}$. Setting
$N^{ii)}:=\bigcup_{i\geq 1}(N_{i}^{\prime}\cup N_{i}^{\prime\prime})$
we then find $N^{ii)}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{ii)})=0$ and our construction ensures that $(\mathrm{P}_{X|Y}(\,\cdot\,|y))_{x^{\prime}_{i}}$ is a Gaussian measure for all $i\geq 1$ and all $y\in F\setminus N^{ii)}$. Theorem 5.1 thus shows that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is Gaussian for $y\in F\setminus N^{ii)}$.
iii). Let us consider the map $g:=\operatorname{id}_{E}$. Since $E$ is separable, $g$ is strongly measurable. Moreover, since $X$ is a Gaussian random variable, Fernique’s theorem, see (8), ensures
$\int_{E}\|g\|_{E}\,\mathrm{d}\mathrm{P}_{X}=\int_{E}\|x\|_{E}\,\mathrm{d}\mathrm{P}_{X}(x)<\infty\,,$
that is, $g:E\to E$ is $\mathrm{P}_{X}$-Bochner integrable. Therefore, Theorem 5.10 shows that there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus(N^{ii)}\cup N_{g})$ we have
$\mu_{X|Y_{n}}(A_{n}(y))=\int_{E}x\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\to\int_{E}x\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)=\mu_{X|Y}(y)\,.$
Setting $N^{iii)}:=N^{ii)}\cup N_{g}$ thus shows iii).
iv). Let us consider the maps $h:E\to[0,\infty)$ and $g:E\to\mathfrak{N}(E^{\prime},E)$ defined by
$h(x):=\|x\|_{E}^{2}\,,$
$g(x):=x\otimes x\,,$
where the tensor products as well as the space of nuclear operators $\mathfrak{N}(E^{\prime},E)$ are discussed in some detail in front of Lemma B.7 and Lemma B.5, respectively. In addition, we consider the probability space $(E,\mathfrak{B}(E),\mathrm{P}_{X})$ and the function $\tilde{X}:=\operatorname{id}_{E}$. Because $E$ is separable, $\tilde{X}$ is strongly measurable. Since $\mathrm{P}_{X}$ is a Gaussian measure, Fernique’s theorem, see (8), then shows
$\int_{E}\|\operatorname{id}_{E}\|^{2}_{E}\,\mathrm{d}\mathrm{P}_{X}=\int_{E}\|x\|_{E}^{2}\,\mathrm{d}\mathrm{P}_{X}(x)<\infty\,,$
and hence we find $\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X},E)}$. Consequently, Lemma B.7 shows that $g=\tilde{X}\otimes\tilde{X}$ is also strongly measurable with $g\in{{\cal L}_{1}(\mathrm{P}_{X},\mathfrak{N}(E^{\prime},E))}$. Therefore, Theorem 5.10 applied with $G:=\mathfrak{N}(E^{\prime},E)$ shows that there exists an $N_{g}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{g})=0$ such that for all $y\in F\setminus N_{g}$ we have
$g\in{{\cal L}_{1}(\mathrm{P}_{X|Y}(\,\cdot\,|y),G)}\qquad\qquad\mbox{ and }\qquad\qquad g\in{{\cal L}_{1}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)),G)}$
for all $n\geq 1$ as well as the $\mathfrak{N}(E^{\prime},E)$-convergence
$\int_{E}x\otimes x\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))$ $\to\int_{E}x\otimes x\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)\,.$ (57)
Moreover, Theorem 5.10 applied to $h=\|\tilde{X}\|_{E}^{2}$ with $G:=\mathbb{R}$ gives an $N_{h}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N_{h})=0$ such that for all $y\in F\setminus N_{h}$ we have
$\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X|Y}(\,\cdot\,|y))}\qquad\qquad\mbox{ and }\qquad\qquad\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)))}$
for all $n\geq 1$. We define $N^{iv)}:=N^{iii)}\cup N_{g}\cup N_{h}$. For $y\in F\setminus N^{iv)}$, Lemma B.7 applied to $\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)))}$ thus shows
$C_{X|Y_{n}}$
$=\operatorname{cov}\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))$
$=\int_{E}(\tilde{X}\otimes\tilde{X})(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))-\biggl{(}\int_{E}\tilde{X}(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\biggr{)}\otimes\biggl{(}\int_{E}\tilde{X}(x)\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))\biggr{)}$
$=\int_{E}x\otimes x\,\mathrm{P}_{X|Y_{n}}(\mathrm{d}x|A_{n}(y))-\mu_{X|Y_{n}}(A_{n}(y))\otimes\mu_{X|Y_{n}}(A_{n}(y))$
$\to\int_{E}x\otimes x\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)-\mu_{X|Y}(y)\otimes\mu_{X|Y}(y)$
with convergence in $\mathfrak{N}(E^{\prime},E)$, where in the last step we used (57) and part iii) in combination with Lemma B.5. Moreover, Lemma B.7 applied to $\tilde{X}\in{{\cal L}_{2}(\mathrm{P}_{X|Y}(\,\cdot\,|y))}$ also shows
$\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)=\int_{E}x\otimes x\,\mathrm{P}_{X|Y}(\mathrm{d}x|y)-\mu_{X|Y}(y)\otimes\mu_{X|Y}(y)$
for all $y\in F\setminus N^{iv)}$. Combining these considerations, we obtain the claimed $C_{X|Y_{n}}\to\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$ with convergence in $\mathfrak{N}(E^{\prime},E)$. Furthermore, all $C_{X|Y_{n}}$ are independent of $y$, and hence so is their limit $\operatorname{cov}\mathrm{P}_{X|Y}(\,\cdot\,|y)$, that is, the notation $C_{X|Y}$ is justified.
v). Let us fix a $y\in F\setminus N^{iv)}$. We then know from i) and Proposition C.5 that
$\varphi_{\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))}(x^{\prime})=\mathrm{e}^{\mathrm{i}\langle x^{\prime},\mu_{X|Y_{n}}(A_{n}(y))\rangle_{E^{\prime},E}}\cdot\exp\biggl{(}-\frac{\langle x^{\prime},C_{X|Y_{n}}x^{\prime}\rangle_{E^{\prime},E}}{2}\biggr{)}\,,\qquad\qquad x^{\prime}\in E^{\prime}\,.$ (58)
Moreover, note that iv) implies
$\bigl{|}\langle x^{\prime},C_{X|Y_{n}}x^{\prime}\rangle_{E^{\prime},E}-\langle x^{\prime},C_{X|Y}x^{\prime}\rangle_{E^{\prime},E}\ \bigr{|}=\bigl{|}\langle x^{\prime},(C_{X|Y_{n}}-C_{X|Y})x^{\prime}\rangle_{E^{\prime},E}\bigr{|}$ $\leq\|C_{X|Y_{n}}-C_{X|Y}\|\cdot\|x^{\prime}\|_{E^{\prime}}^{2}$
$\to 0$
since the nuclear norm dominates the operator norm. Combining this with iii) and (58) we thus obtain
$\varphi_{\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))}(x^{\prime})\to\mathrm{e}^{\mathrm{i}\langle x^{\prime},\mu_{X|Y}(y)\rangle_{E^{\prime},E}}\cdot\exp\biggl{(}-\frac{\langle x^{\prime},C_{X|Y}x^{\prime}\rangle_{E^{\prime},E}}{2}\biggr{)}\,,\qquad\qquad x^{\prime}\in E^{\prime}\,.$ (59)
Furthermore, by Theorem 5.10 we know that there exists an $N^{\star}\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N^{\star})=0$ such that
$\bigl{\{}\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y)):n\geq 1\bigr{\}}$ (60)
is uniformly tight for all $y\in F\setminus N^{\star}$. We define $N^{v)}:=N^{iv)}\cup N^{\star}$. For $y\in F\setminus N^{v)}$ we then have both (59) and (60) and therefore Theorem C.6 shows that there then exists a probability measure $\nu_{y}$ on $X$ such that $\mathrm{P}_{X|Y_{n}}(\,\cdot\,|A_{n}(y))\to\nu_{y}$ weakly and
$\varphi_{\nu_{y}}(x^{\prime})=\mathrm{e}^{\mathrm{i}\langle x^{\prime},\mu_{X|Y}(y)\rangle_{E^{\prime},E}}\cdot\exp\biggl{(}-\frac{\langle x^{\prime},C_{X|Y}x^{\prime}\rangle_{E^{\prime},E}}{2}\biggr{)}\,,\qquad\qquad x^{\prime}\in E^{\prime}\,.$
Proposition C.5 then shows that $\nu_{y}$ is a Gaussian measure with mean $\mu_{X|Y}(y)$ and covariance operator $C_{X|Y}$. By ii), iii), and iv) we further know that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is also a Gaussian measure with mean $\mu_{X|Y}(y)$ and covariance operator $C_{X|Y}$. As discussed in front of (9) we can then conclude that $\nu_{y}=\mathrm{P}_{X|Y}(\,\cdot\,|y)$.
vi). We write $N:=N^{vi)}:=N^{v)}$. Then $N_{\Omega}:=Y^{-1}(N)\in{\cal A}$ satisfies $\mathrm{P}(N_{\Omega})=0$ since $N$ is a $\mathrm{P}_{Y}$-zero set. For $n\geq 1$ we now define the random variable $Z_{n}:\Omega\to E$ by
$Z_{n}:=\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(Y_{n}-\mathbb{E}Y_{n})\,.$
Since for $\omega\in\Omega$ and $y:=Y(\omega)$ we have $\omega\in N_{\Omega}$ if and only if $y\in N$, part iii) together with $A_{n}(y)=A_{n}(Y(\omega))=Y_{n}(\omega)$ then shows that the definition of $Z$ can be written as
$Z(\omega)=\begin{cases}\lim_{n\to\infty}Z_{n}(\omega)\,,&\mbox{ if }\omega\in\Omega\setminus N_{\Omega}\\ 0&\mbox{ else.}\end{cases}$
Consequently, $Z$ as defined in vi) is indeed a random variable and we have the $\mathrm{P}$-almost sure convergence $Z_{n}\to Z$. Moreover, each $Z_{n}$ is $\sigma(Y_{n})=\sigma(A_{n}\circ Y)$-measurable and since $N_{\Omega}\in\sigma(Y)$, we conclude that $Z$ is actually $\sigma(Y)$-measurable. In addition, by Theorem 3.1 each $Z_{n}$ is a version of $\mathbb{E}(X|Y_{n})$ and a Gaussian random variable with $\mathbb{E}Z_{n}=\mathbb{E}X$ and
$\operatorname{cov}Z_{n}$ $=\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\,,$
${\cal G}(Z_{n})$ $\subset{\cal G}(Y_{n})\,.$ (61)
Now, given an $x^{\prime}\in E^{\prime}$, we have $\langle x^{\prime},Z_{n}(\omega)\rangle_{E^{\prime},E}\to\langle x^{\prime},Z(\omega)\rangle_{E^{\prime},E}$ for all $\omega\in\Omega\setminus N_{\Omega}$, and therefore Theorem C.4 shows that $\langle x^{\prime},Z\rangle_{E^{\prime},E}$ is an $\mathbb{R}$-valued Gaussian random variable with
$\mathbb{E}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}$ $\to\mathbb{E}\langle x^{\prime},Z\rangle_{E^{\prime},E}\,,$ (62)
$\operatorname{var}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}$ $\to\operatorname{var}\langle x^{\prime},Z\rangle_{E^{\prime},E}\,.$ (63)
Now, since all $\langle x^{\prime},Z\rangle_{E^{\prime},E}$ are Gaussian, $Z$ is a Gaussian random variable. Moreover, we have $\mathbb{E}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}$, and hence (62) gives
$\langle x^{\prime},\mathbb{E}X\rangle_{E^{\prime},E}=\mathbb{E}\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}Z\rangle_{E^{\prime},E}\,.$
Since this holds for all $x^{\prime}\in E^{\prime}$ we obtain $\mathbb{E}X=\mathbb{E}Z$. Similarly, (B.6) yields
$\operatorname{var}\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}=\langle x^{\prime},\operatorname{cov}(Z_{n})x^{\prime}\rangle_{E^{\prime},E}$ $=\langle x^{\prime},\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)x^{\prime}\rangle_{E^{\prime},E}$
$\to\langle x^{\prime},(\operatorname{cov}(X)-C_{X|Y})x^{\prime}\rangle_{E^{\prime},E}\,,$
where in the last step we used that iv) implies $\|\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)-(\operatorname{cov}(X)-C_{X|Y})\|\to 0$. With  (B.6) and (63) this gives
$\langle x^{\prime},\operatorname{cov}(Z)x^{\prime}\rangle_{E^{\prime},E}=\operatorname{var}\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},(\operatorname{cov}(X)-C_{X|Y})x^{\prime}\rangle_{E^{\prime},E}\,,$
and since this holds for all $x^{\prime}\in E^{\prime}$, Lemma B.2 shows $\operatorname{cov}(Z)=\operatorname{cov}(X)-C_{X|Y}$.
Our next step is to show that $Z$ is a version of $\mathbb{E}(X|Y)$. To this end, we write ${\cal F}_{n}:=\sigma(Y_{n})$ and ${\cal F}_{\infty}:=\sigma(Y)$. Then we already know that $Z_{n}$ is a version of $\mathbb{E}(X|{\cal F}_{n})$. Moreover, Lemma 5.8 shows that $({\cal F}_{n})$ is a filtration with
${\cal F}_{\infty}=\sigma\biggl{(}\bigcup_{n\geq 1}{\cal F}_{n}\biggr{)}\,.$
By the martingale convergence Theorem D.5 we thus find
$Z_{n}\to\mathbb{E}(X|Y)$
in ${{\cal L}_{2}(\mathrm{P},E)}$ and $\mathrm{P}$-almost surely. Since we already have the $\mathrm{P}$-almost sure convergence $Z_{n}\to Z$, we conclude that $\mathrm{P}(\{Z=\mathbb{E}(X|Y)\})=1$, and this in turn implies that $Z$ is indeed a version of $\mathbb{E}(X|Y)$ and $Z_{n}\to Z$ in ${{\cal L}_{2}(\mathrm{P},E)}$.
vii). We set $N^{vii)}:=N^{vi)}$ and fix an $x^{\prime}\in E^{\prime}$. Since $Y_{n}-\mathbb{E}Y_{n}=A_{n}\circ Y-\mathbb{E}(A_{n}\circ Y)=A_{n}\circ(Y-\mathbb{E}Y)$ and $A_{n}$ is bounded and linear, Lemma 5.4 shows ${\cal G}(Y_{n})\subset{\cal G}(Y)$. Combining this with (61) yields
$\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\in{\cal G}(Y_{n})\subset{\cal G}(Y)\,.$
Moreover, $Z_{n}\to Z$ in ${{\cal L}_{2}(\mathrm{P},E)}$ implies $\langle x^{\prime},Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. By $\mathbb{E}Z_{n}=\mathbb{E}X=\mathbb{E}Z$ we conclude $\langle x^{\prime},Z_{n}-\mathbb{E}Z_{n}\rangle_{E^{\prime},E}\to\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}$ in ${L_{2}(\mathrm{P})}$. Since ${\cal G}(Y)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ by definition, we thus find $\langle x^{\prime},Z-\mathbb{E}Z\rangle_{E^{\prime},E}\in{\cal G}(Y)$. In other words, we have shown ${\cal G}(Z)\subset{\cal G}(Y)$.
To verify the formula for the orthogonal projection, we recall that a bounded linear operator on a Hilbert space is an orthogonal projection if and only if it is idempotent and self-adjoint, see e.g. [11, Proposition II.3.3]. Moreover, $\mathbb{E}(\,\cdot\,|Y):{L_{2}(\mathrm{P})}\to{L_{2}(\mathrm{P})}$ is known to be an orthogonal projection, where we note that this result is rarely stated in the formulation above, so we quickly prove it here using well-known properties of conditional expectations, see e.g. [27, Theorem 8.14]. First, for $f\in{L_{2}(\mathrm{P})}$, we have
$\mathbb{E}(\mathbb{E}(f|Y)|Y)=\mathbb{E}(f|Y)$ (64)
since $\mathbb{E}(f|Y)$ is $\sigma(Y)$-measurable by definition. In other words, $\mathbb{E}(\,\cdot\,|Y)$ is idempotent. Moreover, for $f,g\in{L_{2}(\mathrm{P})}$ we have
$\langle f,\mathbb{E}(g|Y)\rangle_{{L_{2}(\mathrm{P})}}=\mathbb{E}\bigl{(}f\cdot{\mathbb{E}(g|Y)}\bigr{)}=\mathbb{E}\bigl{(}\mathbb{E}(f\cdot{\mathbb{E}(g|Y)}|Y)\bigr{)}$ $=\mathbb{E}\bigl{(}{\mathbb{E}(f|Y)}\cdot{\mathbb{E}(g|Y)}\bigr{)}$
$=\bigl{\langle}\mathbb{E}(f|Y),\mathbb{E}(g|Y)\bigr{\rangle}_{{L_{2}(\mathrm{P})}}\,.$
By symmetry we then find
$\langle f,\mathbb{E}(g|Y)\rangle_{{L_{2}(\mathrm{P})}}=\langle\mathbb{E}(f|Y),g\rangle_{{L_{2}(\mathrm{P})}}\,,$ (65)
that is, $\mathbb{E}(\,\cdot\,|Y)$ is also self-adjoint. Our next goal is to show
$\mathbb{E}(f|Y)\in{\cal G}(Y)\,,\qquad\qquad f\in{\cal G}(X,Y).$ (66)
To this end, we fix an $f\in{\cal G}(X,Y)$. By the definition of ${\cal G}(X,Y)$, there then exists a sequence $(f_{n})$ of the form $f_{n}=\langle x_{n}^{\prime},X\rangle_{E^{\prime},E}+\langle y_{n}^{\prime},Y\rangle_{E^{\prime},E}$ with $f_{n}\to f$ in ${L_{2}(\mathrm{P})}$, where $x_{n}^{\prime}\in E^{\prime}$ and $y_{n}^{\prime}\in F^{\prime}$ are suitable functionals. Now, ${\cal G}(Z)\subset{\cal G}(Y)$ implies
$\mathbb{E}\bigl{(}\langle x_{n}^{\prime},X\rangle_{E^{\prime},E}|Y\bigr{)}=\bigl{\langle}x_{n}^{\prime},\mathbb{E}(X|Y)\bigr{\rangle}_{E^{\prime},E}=\langle x_{n}^{\prime},Z\rangle_{E^{\prime},E}\in{\cal G}(Y)\,,$
where in the first step we used Theorem D.4. Moreover, since $\langle y_{n}^{\prime},Y\rangle_{E^{\prime},E}$ is $\sigma(Y)$-measurable, we have
$\mathbb{E}\bigl{(}\langle y_{n}^{\prime},Y\rangle_{E^{\prime},E}|Y\bigr{)}=\langle y_{n}^{\prime},Y\rangle_{E^{\prime},E}\in{\cal G}(Y)\,.$
Combining both considerations shows $\mathbb{E}(f_{n}|Y)\in{\cal G}(Y)$ for all $n\geq 1$. Since $\mathbb{E}(\,\cdot\,|Y):{L_{2}(\mathrm{P})}\to{L_{2}(\mathrm{P})}$ is continuous, we further have the ${L_{2}(\mathrm{P})}$-convergence
$\mathbb{E}(f_{n}|Y)\to\mathbb{E}(f|Y)\,,$
and since ${\cal G}(Y)$ is a closed subspace of ${L_{2}(\mathrm{P})}$ we conclude $\mathbb{E}(f|Y)\in{\cal G}(Y)$. This finishes the proof of (66). Since ${\cal G}(Y)\subset{\cal G}(X,Y)$ we can now consider the operator
$\Pi:{\cal G}(X,Y)$ $\to{\cal G}(X,Y)$
$f$ $\mapsto\mathbb{E}(f|Y)\,.$
Since $\Pi$ is the restriction of $\mathbb{E}(\,\cdot\,|Y)$ to ${\cal G}(X,Y)$, we see that $\Pi$ is bounded and linear. Moreover, (64) shows that $\Pi$ is idempotent and since (65) holds for all $f,g\in{L_{2}(\mathrm{P})}$ it also holds for all $f,g\in{\cal G}(X,Y)$, and hence $\Pi$ is self-adjoint. In other words, $\Pi$ is an orthogonal projection.
It remains to show $\operatorname{ran}\Pi={\cal G}(Y)$. Here, we first note that the inclusion $\subset$” has already been established in (66). For the proof of the converse inclusion, we pick an $f\in{\cal G}(Y)$. By Lemma 5.4 we know that the $\mathrm{P}$-equivalence class of $f$ contains a $\sigma(Y)$-measurable representative, so we may assume without loss of generality that $f$ is $\sigma(Y)$-measurable. This leads to
$\Pi f=\mathbb{E}(f|Y)=f\,,$
that is we have found $f\in\operatorname{ran}\Pi$.
viii). We fix an $x^{\prime}\in E^{\prime}$, for which $\langle x^{\prime},X\rangle_{E^{\prime},E}$ is $\sigma(Y)$-measurable. Using Theorem D.4 we then obtain
$\bigl{\langle}x^{\prime},\mu_{X|Y}\circ Y\bigr{\rangle}_{E^{\prime},E}=\langle x^{\prime},Z\rangle_{E^{\prime},E}=\bigl{\langle}x^{\prime},\mathbb{E}(X|Y)\bigr{\rangle}_{E^{\prime},E}=\mathbb{E}\bigl{(}\langle x^{\prime},X\rangle_{E^{\prime},E}\bigl{|}Y\bigr{)}=\langle x^{\prime},X\rangle_{E^{\prime},E}\,,$
where the first identity holds up to the $\mathrm{P}$-zero set $Y^{-1}(N^{vii)})\in{\cal A}$ and in the second identity we simply considered the version $Z$ of $\mathbb{E}(X|Y)$. Moreover, the third and forth identities are meant in the sense that both $\langle x^{\prime},Z\rangle_{E^{\prime},E}=\langle x^{\prime},\mathbb{E}(X|Y)\rangle_{E^{\prime},E}$ and $\langle x^{\prime},X\rangle_{E^{\prime},E}$ are versions of $\mathbb{E}(\langle x^{\prime},X\rangle_{E^{\prime},E}|Y)$. Consequently, the set
$N_{0}:=\bigl{\{}\langle x^{\prime},Z\rangle_{E^{\prime},E}\neq\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{\}}$
is a $\mathrm{P}$-zero set with $N_{0}\in\sigma(Y)$. Therefore, there exists an $N^{\prime}\in\sigma(F^{\prime})$ with $Y^{-1}(N^{\prime})=N_{0}$, and in particular we have $\mathrm{P}_{Y}(N^{\prime})=0$. Setting $N^{viii)}:=N^{vii)}\cup N^{\prime}$ then gives the first assertion.
Finally, by (B.6), part vi), Theorem D.4, and the assumed $\sigma$-measurability we obtain
$\langle x^{\prime},\operatorname{cov}(Z)x^{\prime}\rangle_{E^{\prime},E}=\operatorname{var}(\langle x^{\prime},Z\rangle_{E^{\prime},E})=\operatorname{var}\bigl{(}\langle x^{\prime},\mathbb{E}(X|Y)\rangle_{E^{\prime},E}\bigr{)}$ $=\operatorname{var}\bigl{(}\mathbb{E}(\langle x^{\prime},X\rangle_{E^{\prime},E}|Y)\bigr{)}$
$=\operatorname{var}\bigl{(}\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{)}$
$=\langle x^{\prime},\operatorname{cov}(X)x^{\prime}\rangle_{E^{\prime},E}\,.$
Using vi), we then find
$\langle x^{\prime},C_{X|Y}x^{\prime}\rangle_{E^{\prime},E}=\bigl{\langle}x^{\prime},(\operatorname{cov}(X)-\operatorname{cov}(Z))x^{\prime}\bigr{\rangle}_{E^{\prime},E}=\langle x^{\prime},\operatorname{cov}(X)x^{\prime}\rangle_{E^{\prime},E}-\bigl{\langle}x^{\prime},\operatorname{cov}(Z)x^{\prime}\bigr{\rangle}_{E^{\prime},E}=0\,.$
By (B.9) we then find $C_{X|Y}x^{\prime}=0$ as discussed there. ∎
5.6 Proofs of Theorems 3.5 and 3.6
Proof of Theorem 3.5: Throughout the proof we use the shorthand $C_{X|Y}:=\operatorname{cov}(X)-\operatorname{cov}(\mu_{X|Y}\circ Y)$. It suffices to verify the three conditions of Definition 2.5.
We begin by observing that $\mathrm{P}_{X|Y}(\,\cdot\,|y)$ is, of course, a probability measure on $\mathfrak{B}(E)$ for every $y\in F$ by our construction (17).
To verify the measurability condition ii), we define
${\cal E}$ $:=\bigcup_{x^{\prime}\in E^{\prime}}\sigma(x^{\prime})\,,$
${\cal B}$ :=$:=\bigl{\{}B\in\sigma(E^{\prime}):y\mapsto\mathrm{P}_{X|Y}(B|y)\mbox{ is measurable }\bigr{\}}\,.$
We first show that ${\cal E}\subset{\cal B}$. To this end, we choose a $B\in{\cal E}$. Then there exists an $x^{\prime}\in E^{\prime}$ and an $A\in\mathfrak{B}$ such that $B=(x^{\prime})^{-1}(A)$. For fixed $y\in F$ we then obtain
$\mathrm{P}_{X|Y}(B|y)={\cal N}\bigl{(}\mu(y),C_{X|Y}\bigr{)}\bigl{(}(x^{\prime})^{-1}(A)\bigr{)}=\Bigl{(}{\cal N}\bigl{(}\mu(y),C_{X|Y}\bigr{)}\Bigr{)}_{x^{\prime}}(A)={\cal N}(m(y),\sigma^{2})(A)\,,$
where we have set $m(y):=\langle x^{\prime},\mu(y)\rangle_{E^{\prime},E}$ and $\sigma^{2}$ denotes the $1\times 1$-matrix representing the linear operator $x^{\prime}\circ C_{X|Y}\circ(x^{\prime})^{\prime}:\mathbb{R}^{\prime}\to\mathbb{R}$. Note the map $y\mapsto m(y)$ is $(\sigma(F^{\prime}),\mathfrak{B})$-measurable since $\mu$ is $(\sigma(F^{\prime}),\sigma(E^{\prime}))$-measurable and $x^{\prime}$ is $(\sigma(E^{\prime}),\mathfrak{B})$-measurable.
It remains to verify the measurability of the map $y\mapsto{\cal N}(m(y),\sigma^{2})(A)$. In the case $\sigma^{2}=0$ this measurability immediately follows from
${\cal N}\bigl{(}m(y),\sigma^{2}\bigr{)}(A)=\delta_{\{m(y)\}}(A)=\boldsymbol{1}_{A}(m(y))$
and the measurability of $m$ and $A$. Moreover, in the case $\sigma^{2}>0$ we have
${\cal N}(m(y),\sigma^{2})(A)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\int_{\mathbb{R}}\boldsymbol{1}_{A}(t)\exp\Bigl{(}-\frac{(t-m(y))^{2}}{2\sigma^{2}}\Bigr{)}\,\mathrm{d}t\,.$
Now, the integrand $f(t,y):=\boldsymbol{1}_{A}(t)\exp\bigl{(}-\frac{(t-m(y))^{2}}{2\sigma^{2}}\bigr{)}$ is a non-negative $(\mathfrak{B}\otimes\sigma(F^{\prime}),\mathfrak{B})$-measurable function, and therefore Fubini-Tonelli’s theorem, see e.g. [27, Theorem 14.16], yields the measurability of $y\mapsto{\cal N}(m(y),\sigma^{2})(A)$. In both cases we have thus found $B\in{\cal B}$, and hence we have established ${\cal E}\subset{\cal B}$.
Let us now show that ${\cal B}$ is a $\sigma$-algebra. Here, $\emptyset\in{\cal B}$ is obvious since $y\mapsto\mathrm{P}_{X|Y}(\emptyset|y)=0$ for all $y\in F$. Moreover, if we have a $B\in{\cal B}$, then
$\mathrm{P}_{X|Y}(E\setminus B|y)=1-\mathrm{P}_{X|Y}(B|y)\,,\qquad\qquad y\in F$
quickly shows that $E\setminus B\in{\cal B}$. Similarly, if we have mutually disjoint $B_{1},B_{2},\dots\in{\cal B}$ and we set $B:=B_{1}\cup B_{2}\cup\dots$, then we have
$\mathrm{P}_{X|Y}(B|y)=\sum_{i=1}^{\infty}\mathrm{P}_{X|Y}(B_{i}|y)\,,\qquad\qquad y\in F$
and since each $y\mapsto\mathrm{P}_{X|Y}(B_{i}|y)$ is measurable, we conclude that $B\in{\cal B}$.
With these preparations we obtain $\sigma({\cal E})\subset\sigma({\cal B})={\cal B}\subset\sigma(E^{\prime})$, and since we also have $\sigma({\cal E})=\sigma(E^{\prime})$, we conclude that ${\cal B}=\sigma(E^{\prime})=\mathfrak{B}(E)$. This shows condition ii) of Definition 2.5.
Let us finally verify condition iii) of Definition 2.5. To this end, we fix an arbitrary version $\tilde{\mathrm{P}}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$ and an $N\in\sigma(F^{\prime})$ with $\mathrm{P}_{Y}(N)=0$ according to Theorem 3.4. We define $\mu_{X|Y}:F\to E$ by (16) and write
$\tilde{Z}:=\mu_{X|Y}\circ Y\,.$
Then part vi) of Theorem 3.4 shows that $\tilde{Z}$ is a version of $\mathbb{E}(X|Y)$, and hence we have
$0=\mathrm{P}(\{Z\neq\tilde{Z}\})=\mathrm{P}(\{\mu\circ Y\neq\mu_{X|Y}\circ Y\})=\mathrm{P}\bigl{(}Y^{-1}(\{\mu\neq\mu_{X|Y}\})\bigr{)}=\mathrm{P}_{Y}(\{\mu\neq\mu_{X|Y}\})\,,$
where in the first step we used the almost sure uniqueness of conditional expectations in the Banach space valued case, which is discussed after Definition D.3. In other words, the set $N_{1}:=\{\mu\neq\mu_{X|Y}\}$ satisfies $N_{1}\in\sigma(F^{\prime})=\mathfrak{B}(F)$ and $\mathrm{P}_{Y}(N_{1})=0$. For the $\mathrm{P}_{Y}$-zero set $N_{2}:=N\cup N_{1}$ and $y\in F\setminus N_{2}$ we then have
$\mathrm{P}_{X|Y}(\,\cdot\,|y)={\cal N}\bigl{(}\mu(y),C_{X|Y}\bigr{)}={\cal N}\bigl{(}\mu_{X|Y}(y),C_{X|Y}\bigr{)}=\tilde{\mathrm{P}}_{X|Y}(\,\cdot\,|y)\,.$
For $B\in\mathfrak{B}(E)$ and $A\in\mathfrak{B}(F)$ we thus find
$\mathrm{P}_{(X,Y)}(B\times A)=\int_{F}\boldsymbol{1}_{A}(y)\tilde{\mathrm{P}}(B|y)\,\mathrm{d}\mathrm{P}_{Y}(y)=\int_{F}\boldsymbol{1}_{A}(y)\mathrm{P}(B|y)\,\mathrm{d}\mathrm{P}_{Y}(y)\,,$
i.e. we have verified condition iii) of Definition 2.5. ∎
Proof of Theorem 3.6: Let $B_{n}:F\to E$ be the bounded linear operator given by
$B_{n}:=\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}A_{n}\,.$
Since $\mathbb{E}Y_{n}=A_{n}(\mathbb{E}Y)$, we then find
$F_{\mathrm{conv},\mathfrak{A}}=\bigl{\{}y\in F:\exists\lim_{n\to\infty}B_{n}(y-\mathbb{E}Y)\bigr{\}}\,.$
We now pick a version $\mathrm{P}_{X|Y}(\,\cdot\,|\,\cdot\,)$ of the regular conditional probability of $X$ given $Y$ and a $\mathrm{P}_{Y}$-zero set $N\in\sigma(F^{\prime})$ according to Theorem 3.4. In addition, we define $\mu_{X|Y}:F\to E$ with the help of part iii) of Theorem 3.4 and (16). Note that part iii) of Theorem 3.4 shows $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$.
i). Since $E$ is complete, we have
$F_{\mathrm{conv},\mathfrak{A}}$ =$=\bigl{\{}y\in F:(B_{n}(y-\mathbb{E}Y))\mbox{ is Cauchy sequence in $E$}\bigr{\}}$
$=\bigcap_{k\geq 1}\bigcup_{n_{0}\geq 1}\bigcap_{n,m\geq n_{0}}\bigl{\{}y\in F:\|B_{n}(y-\mathbb{E}Y)-B_{m}(y-\mathbb{E}Y)\|_{E}\leq 1/k\bigr{\}}\,.$
By the continuity of the map $y\mapsto\|B_{n}(y-\mathbb{E}Y)-B_{m}(y-\mathbb{E}Y)\|_{E}$, we further see that
$\bigl{\{}y\in F:\|B_{n}(y-\mathbb{E}Y)-B_{m}(y-\mathbb{E}Y)\|_{E}\leq 1/k\bigr{\}}\in\mathfrak{B}(F)$
for all $k,n,m\in\mathbb{N}$. Combining both observations yields $F_{\mathrm{conv},\mathfrak{A}}\in\mathfrak{B}(F)=\sigma(F^{\prime})$.
For the proof of $\mathrm{P}_{Y}(F_{\mathrm{conv},\mathfrak{A}})=1$ we simply combine the already observed inclusion $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$ with $\mathrm{P}_{Y}(F\setminus N)=1$.
ii). We first note that $B_{n}(0)=0$ shows $\mathbb{E}Y\in F_{\mathrm{conv},\mathfrak{A}}$, and hence $0\in F_{X|Y,\mathfrak{A}}$. Moreover, for $y_{1},y_{2}\in F_{X|Y,\mathfrak{A}}$ and $\alpha_{1},\alpha_{2}\in\mathbb{R}$ we have $\tilde{y}_{i}:=y_{i}+\mathbb{E}Y\in F_{\mathrm{conv},\mathfrak{A}}$ and
$B_{n}\bigl{(}(\alpha_{1}y_{1}+\alpha_{2}y_{2}+\mathbb{E}Y)-\mathbb{E}Y\bigr{)}$ $=\alpha_{1}B_{n}y_{1}+\alpha_{2}B_{n}y_{2}$
$=\alpha_{1}B_{n}(\tilde{y}_{1}-\mathbb{E}Y)+\alpha_{2}B_{n}(\tilde{y}_{2}-\mathbb{E}Y)\,.$
From this we easily conclude that $\alpha_{1}y_{1}+\alpha_{2}y_{2}+\mathbb{E}Y\in F_{\mathrm{conv},\mathfrak{A}}$, that is $\alpha_{1}y_{1}+\alpha_{2}y_{2}\in F_{X|Y,\mathfrak{A}}$. Finally, $F_{\mathrm{conv},\mathfrak{A}}\in\sigma(F^{\prime})=\mathfrak{B}(F)$ implies $F_{X|Y,\mathfrak{A}}\in\mathfrak{B}(F)=\sigma(F^{\prime})$ since translations $F\to F$ are continuous and thus $(\mathfrak{B}(F),\mathfrak{B}(F))$-measurable.
Let us now show $\operatorname{ran}\operatorname{cov}(Y,X)\subset F_{X|Y,\mathfrak{A}}$. To this end, we fix an $x^{\prime}\in E^{\prime}$ and set $y:=\operatorname{cov}(Y,X)(x^{\prime})$. This gives $A_{n}y=(A_{n}\circ\operatorname{cov}(Y,X))(x^{\prime})=\operatorname{cov}(Y_{n},X)(x^{\prime})$ and by part iv) of Theorem 3.4 we know that
$\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(x^{\prime})=\lim_{n\to\infty}\bigl{(}\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\bigr{)}(A_{n}y)$
exists. In other words, we have $y+\mathbb{E}Y\in F_{\mathrm{conv},\mathfrak{A}}$, and hence $y\in-\mathbb{E}Y+F_{\mathrm{conv},\mathfrak{A}}=F_{X|Y,\mathfrak{A}}$.
iii). The definition of $\mu_{X|Y,\mathfrak{A}}:F\to E$ can be rewritten as
$\mu_{X|Y,\mathfrak{A}}(y)=\mathbb{E}X+\lim_{n\to\infty}\boldsymbol{1}_{F_{\mathrm{conv},\mathfrak{A}}}(y)\cdot B_{n}(y-\mathbb{E}Y)\,.$
In particular, $\mu_{X|Y,\mathfrak{A}}$ is the pointwise limit of a sequence of strongly measurable functions $F\to E$, and hence also strongly measurable. Moreover, our construction ensures $\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}=F_{\mathrm{conv},\mathfrak{A}}\setminus(F\setminus N)$ since we already know $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$. This shows
$\mathrm{P}\bigl{(}\{\mu_{X|Y}\circ Y\neq\mu_{X|Y,\mathfrak{A}}\circ Y\}\bigr{)}=\mathrm{P}_{Y}\bigl{(}\{\mu_{X|Y}\neq\mu_{X|Y,\mathfrak{A}}\}\bigr{)}=\mathrm{P}_{Y}(F_{\mathrm{conv},\mathfrak{A}})-\mathrm{P}_{Y}(F\setminus N)=0\,,$
where we again used $F\setminus N\subset F_{\mathrm{conv},\mathfrak{A}}$. Since $\mu_{X|Y}\circ Y$ is a version of $\mathbb{E}(X|Y)$ by part iv) of Theorem 3.4, we conclude that $\mu_{X|Y,\mathfrak{A}}\circ Y$ is also a version of $\mathbb{E}(X|Y)$.
iv). This is a direct consequence of iii) and Theorem 3.5. ∎
6 Proofs for the Examples in Section 4
Proof of Proposition 4.1: We will apply Proposition 3.3. To verify (14) we pick a $y^{\prime}\in F^{\prime}$. Without loss of generality we may assume $y^{\prime}\neq 0$. We define $y_{\infty}^{\prime}:=\|y^{\prime}\|^{-1}y^{\prime}$. Since $y_{\infty}^{\prime}\in B_{F^{\prime}}$ and the relative $w^{*}$-topology of $B_{E^{\prime}}$ is metrizable, see Theorem 2.1, there then exists a sequence $(y_{i_{k}}^{\prime})\subset\mathfrak{D}_{F^{\prime}}$ with $y_{i_{k}}^{\prime}\to y_{\infty}^{\prime}$ in the $w^{*}$-topology. This shows
$y_{\infty}^{\prime}\in\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}\,,$
and rescaling gives the inclusion $\supset$” in (14). The converse inclusion $\subset$” is trivially satisfied. ∎
Proof of Proposition 4.2: For $n\geq 1$ we define $A_{n}:F\to\mathbb{R}^{n}$ by (15). This definition ensures $\operatorname{span}\mathfrak{C}_{A_{n}}\subset\operatorname{span}\mathfrak{C}_{A_{n+1}}$ for all $n\geq 1$ and hence we have
$\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}=\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\,.$
Moreover, by Lemma 5.6 we know $\operatorname{span}\mathfrak{C}_{A_{n}}=(\ker{A_{n}})^{\perp\perp}$, and hence we find
$\biggl{(}\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\biggr{)}_{\perp\perp}=\bigcap_{n\geq 1}(\operatorname{span}\mathfrak{C}_{A_{n}})_{\perp\perp}=\bigcap_{n\geq 1}\bigl{(}(\ker A_{n})^{\perp\perp}\bigr{)}_{\perp\perp}=\bigcap_{n\geq 1}\overline{\ker A_{n}}^{\|\cdot\|_{F}}=\bigcap_{n\geq 1}\ker A_{n}\,,$
where the first identity is a simple exercise and the third identity can be found in e.g. [32, Proposition 1.10.15]. We thus obtain
$\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}=\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}}=\overline{\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}}^{w^{*}}$ $=\biggl{(}\biggl{(}\bigcup_{n\geq 1}\operatorname{span}\mathfrak{C}_{A_{n}}\biggr{)}_{\perp\perp}\biggr{)}^{\perp\perp}$
$=\biggl{(}\bigcap_{n\geq 1}\ker A_{n}\biggr{)}^{\perp\perp}\,,$
where the first identity follows from e.g. [32, Corollary 2.7.13] and the third identity can be found in e.g. [32, Proposition 2.6.6]. Moreover, we have $A_{n}(y)=0$ if and only if $\langle y_{1}^{\prime},y\rangle_{F^{\prime},F}=\dots=\langle y_{n}^{\prime},y\rangle_{F^{\prime},F}=0$. In other words, the identity $\ker A_{n}=\ker y_{1}^{\prime}\cap\dots\cap\ker y_{n}^{\prime}$ holds. Combining this with our previous considerations yields
$\overline{\operatorname{span}\{y_{i}^{\prime}:i\geq 1\}}^{w^{*}\mathrm{\!-seq}}=\biggl{(}\bigcap_{i\geq 1}\ker y_{i}^{\prime}\biggr{)}^{\perp\perp}\,.$
Finally note that $\{y_{i}^{\prime}:i\geq 1\}\subset F^{\prime}$ is separating if and only if we have
$\bigcap_{i\geq 1}\ker y_{i}^{\prime}=\{0\}\,.$
Now the equivalence follows from the well-known fact that $\{0\}\subset F$ is the only non-empty subset $B\subset F$ with $B^{\perp\perp}=F^{\prime}$. ∎
Proof of Lemma 4.3: Let us identify $H$ with $H^{\prime}$ via the canonical and isometric Fréchet-Riesz isomorphism $h\mapsto\hat{h}:=\langle h,\,\cdot\,\rangle_{H}$. Then $(\hat{e}_{n})_{n\geq 1}$ is an ONB of $H^{\prime}$, and hence the assertion follows from Proposition 3.3. ∎
Proof of Proposition 4.4: We first show the existence of a countable $\operatorname{d}_{k}$-dense subset $T_{0}$ of $T$. To this end, we first recall that subsets of separable metric spaces are separable, and therefore, the set
$K_{T}:=\bigl{\{}k(t,\,\cdot\,):t\in T\bigr{\}}$
is $\|\cdot\|_{H}$-separable. Consequently, there exists a countable $T_{0}\subset T$ such that
$K_{0}:=\bigl{\{}k(t,\,\cdot\,):t\in T_{0}\bigr{\}}$
is $\|\cdot\|_{H}$-dense in $K_{T}$. To verify, that $T_{0}$ is the desired set, we fix a $t\in T$. Then there exists a sequence $(t_{n})\subset T_{0}$ such that $k(t_{n},\,\cdot\,)\to k(t,\,\cdot\,)$ in $H$. By the definition of $\operatorname{d}_{k}$, this gives $\operatorname{d}_{k}(t,t_{n})\to 0$.
For the proof of the second assertion, we first note that for $h\in H$ and $t,t^{\prime}\in T$ the reproducing property of $k$ and the Cauchy-Schwarz inequality yield
$|h(t)-h(t^{\prime})|=\bigl{|}\langle h,k(t,\,\cdot\,)-k(t^{\prime},\,\cdot\,)\rangle_{H}\bigr{|}\leq\|h\|_{H}\cdot\|k(t,\,\cdot\,)-k(t^{\prime},\,\cdot\,)\|_{H}=\|h\|_{H}\cdot\operatorname{d}_{k}(t,t^{\prime})\,.$
Consequently, every $h\in H$ is $\operatorname{d}_{k}$-continuous. Identifying $H$ with $H^{\prime}$ via the isometric Fréchet-Riesz isomorphism $h\mapsto\hat{h}:=\langle h,\,\cdot\,\rangle_{H}$ we then conclude that the sequence $(\hat{k}(t_{i},\,\cdot\,))$ is separating. Combining Propositions 3.3 and 4.2 then yields the assertion. ∎
Lemma 6.1.
Let $(T,\operatorname{d})$ be a compact metric space and $T_{0}\subset T$ be dense. Then we have
$\overline{\operatorname{aco}\{\delta_{t}:t\in T_{0}\}}^{w^{*}\mathrm{\!-seq}}=B_{C(T)^{\prime}}\,,$
where $\operatorname{aco}A$ denotes the absolute convex hull of a set $A$.
Proof of Lemma 6.1: We first consider the case $T_{0}=T$. Here, we note that the set of extreme points $\operatorname{ex}B_{{\cal M}(T)}$ of $B_{{\cal M}(T)}$ is given by
$\operatorname{ex}B_{{\cal M}(T)}=\bigl{\{}\alpha\delta_{\{t\}}:t\in T\mbox{ and }\alpha\in\{-1,1\}\bigr{\}}\,,$
see e.g. [11, Theorem V.8.4]. Using the identification (21) of $C(T)^{\prime}$ with ${\cal M}(T)$ and the obvious $\hat{\delta}_{\{t\}}=\delta_{t}$ we conclude that the set $\operatorname{ex}B_{C(T)^{\prime}}$ of extreme points of $B_{C(T)^{\prime}}$ is given by
$\operatorname{ex}B_{C(T)^{\prime}}=\bigl{\{}\alpha\delta_{t}:t\in T\mbox{ and }\alpha\in\{-1,1\}\bigr{\}}\,.$
Krein-Milman’s theorem, see e.g. [32, Corollary 2.10.9], thus gives
$\overline{\operatorname{aco}\{\delta_{t}:t\in T\}}^{w^{*}}=\overline{\operatorname{co}\bigl{\{}\alpha\delta_{t}:t\in T\mbox{ and }\alpha\in\{-1,1\}\bigr{\}}}^{w^{*}}=B_{C(T)^{\prime}}\,,$
where the first identity is trivial and $\operatorname{co}A$ denotes the convex hull of a set $A$. Moreover, $C(T)$ is separable, and hence the relative $w^{*}$-topology on $B_{C(T)^{\prime}}$ is metrizable, see Theorem 2.1. This gives
$\overline{\operatorname{aco}\{\delta_{t}:t\in T\}}^{w^{*}}=\overline{\operatorname{aco}\{\delta_{t}:t\in T\}}^{w^{*}\mathrm{\!-seq}}\,,$
where we note this identity can also be deduced from [32, Corollary 2.7.13].
Let us now prove the assertion in the case of a general, dense $T_{0}\subset T$. To this end, we fix a $\hat{\mu}\in B_{C(T)^{\prime}}$. By the already established case we then find a sequence
$(\hat{\mu}_{n})\subset\operatorname{aco}\{\delta_{t}:t\in T\}$
with $\langle\hat{\mu}_{n},f\rangle_{C(T)^{\prime},C(T)}\to\langle\hat{\mu},f\rangle_{C(T)^{\prime},C(T)}$ for all $f\in C(T)$. Clearly, each $\hat{\mu}_{n}$ is of the form
$\hat{\mu}_{n}=\sum_{i=1}^{m_{n}}\alpha_{i,n}\delta_{t_{i,n}}\,,$ (67)
where $t_{i,n}\in T$ and $\sum_{i=1}^{m_{n}}|\alpha_{i,n}|\leq 1$. Since $T_{0}\subset T$ is dense, we then find $s_{i,n}\in T_{0}$ with $|t_{i,n}-s_{i,n}|\leq 1/n$ for all $i=1,\dots,m_{n}$ and $n\geq 1$. We define
$\hat{\nu}_{n}:=\sum_{i=1}^{m_{n}}\alpha_{i,n}\delta_{s_{i,n}}\,.$
Let us fix an $f\in C(T)$ and denote its uniform modulus of continuity by $\omega(f,\,\cdot\,)$. Then we have
$\bigl{|}\langle\hat{\mu}_{n}-\hat{\nu}_{n},f\rangle_{C(T)^{\prime},C(T)}\bigr{|}=\biggl{|}\sum_{i=1}^{m_{n}}\alpha_{i,n}\langle\delta_{t_{i,n}}-\delta_{s_{i,n}},f\rangle_{C(T)^{\prime},C(T)}\biggr{|}$ $\leq\sum_{i=1}^{m_{n}}|\alpha_{i,n}|\cdot|f(t_{i,n})-f(s_{i,n})|$
$\leq\sum_{i=1}^{m_{n}}|\alpha_{i,n}|\cdot\omega(f,t_{i,n}-s_{i,n})$
$\leq\omega(f,1/n)\,,$
and $\lim_{n\to\infty}\omega(f,1/n)=0$ as $f$ is uniformly continuous by the compactness of $T$. This gives
$\bigl{|}\langle\hat{\mu}-\hat{\nu}_{n},f\rangle_{C(T)^{\prime},C(T)}\bigr{|}\leq\bigl{|}\langle\hat{\mu}-\hat{\mu}_{n},f\rangle_{C(T)^{\prime},C(T)}\bigr{|}+\bigl{|}\langle\hat{\mu}_{n}-\hat{\nu}_{n},f\rangle_{C(T)^{\prime},C(T)}\bigr{|}\to 0\,.$
This gives $\hat{\mu}\in\overline{\operatorname{aco}\{\delta_{t}:t\in T_{0}\}}^{w^{*}\mathrm{\!-seq}}$, and hence we have shown the inclusion $\supset$”. The converse inclusion $\subset$” follows by monotonicity from the already established case $T=T_{0}$. ∎
Proof of Lemma 4.5: i). If $X$ is weakly measurable, then (22) ensures that $(X_{t})_{t\in T}$ is a stochastic process with continuous paths. To show the converse, we fix a $\mu\in{\cal M}(T)$ with Hahn-Jordan decomposition $\mu=\mu^{+}-\mu^{-}$ and we consider the map
$\Omega\times T$ $\to\mathbb{R}$ (68)
$(\omega,t)$ $\mapsto X_{t}(\omega)\,.$
Then, $\omega\mapsto X_{t}(\omega)$ is measurable for all $t\in T$ since we have a stochastic process. In addition, $t\mapsto X_{t}(\omega)$ is continuous for all $\omega\in\Omega$ since the stochastic process has continuous paths. Moreover, $(T,\operatorname{d})$ is separable, and therefore (68) is ${\cal A}\otimes\mathfrak{B}(T)$-measurable, see e.g. [2, Lemma 4.51]. Fubini-Tonelli’s theorem, see e.g. [27, Theorem 14.16], then shows that the two functions
$\omega\mapsto\langle\mu^{\pm},X(\omega)\rangle_{{\cal M}(T),C(T)}:=\int_{\Omega}X_{t}(\omega)\,\mathrm{d}\mu^{\pm}(t)$
are measurable, and hence $\omega\mapsto\langle\mu,X(\omega)\rangle_{{\cal M}(T),C(T)}$ is also measurable. Using the identification of $C(T)^{\prime}$ with ${\cal M}(T)$ given in (21) we conclude that $X$ is weakly measurable.
ii). Let us first assume that $X$ is Gaussian. We fix some mutually distinct $t_{1},\dots,t_{n}\in T$ and write $Y:=(X_{t_{1}},\dots,X_{t_{n}})$. Clearly, our goal is to show that the $\mathbb{R}^{n}$-valued random variable $Y$ is Gaussian. To this end, we fix some $a\in\mathbb{R}^{n}$ and write $\hat{\mu}:=a_{1}\delta_{t_{1}}+\dots+a_{n}\delta_{t_{n}}$. Then a simple calculation using (22) shows
$\langle a,Y\rangle_{\mathbb{R}^{n}}=\sum_{i=1}^{n}a_{i}X_{t_{i}}=\sum_{i=1}^{n}a_{i}\langle\delta_{t_{i}},X\rangle_{C(T)^{\prime},C(T)}=\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}\,,$ (69)
and since $X$ is a Gaussian random variable, so is $\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}$, and thus also $\langle a,Y\rangle_{\mathbb{R}^{n}}$. Since $a\in\mathbb{R}^{n}$ was chosen arbitrarily, we conclude that $Y$ is Gaussian.
Let us now assume that $(X_{t})_{t\in T}$ is a Gaussian process with continuous paths. It suffices to show that $X:\Omega\to C(T)$ is Gaussian. To this end, we consider the set
$\mathfrak{D}:=\operatorname{aco}\{\delta_{t}:t\in T\}\,.$
By Lemma 6.1 we know that $\mathfrak{D}$ is $w^{*}$-dense in $B_{C(T)^{\prime}}$. By (21) and Theorem 5.1 it therefore suffices to show that $\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}$ is Gaussian for all $\hat{\mu}\in C(T)^{\prime}$ of the form $\hat{\mu}=a_{1}\delta_{t_{1}}+\dots+a_{n}\delta_{t_{n}}$, where the coefficients $a_{i}\in\mathbb{R}$ satisfy $|a_{1}|+\dots+|a_{n}|\leq 1$. Let us fix such a $\hat{\mu}$. Defining $Y:=(X_{t_{1}},\dots,X_{t_{n}})$, we then find
$\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}=\langle a,Y\rangle_{\mathbb{R}^{n}}$
as in (69). Moreover, since $Y$ is Gaussian, so is $\langle a,Y\rangle_{\mathbb{R}^{n}}$, and thus $\langle\hat{\mu},X\rangle_{C(T)^{\prime},C(T)}$ is also Gaussian. ∎
Lemma 6.2.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $(T,\operatorname{d})$ be a compact metric space, and $(X_{t})_{t\in T}$ be a Gaussian processes on $(\Omega,{\cal A},\mathrm{P})$ with continuous paths. Moreover, let $X:\Omega\to C(T)$ be the Gaussian random variable associated to the process via (23) and $S:=\{s_{1},\dots,s_{n}\}\subset T$. Finally, let
$Y:=I_{S}\circ X$
and $m(S):=(m(s_{1}),\dots,m(s_{n}))$, where $I_{S}:C(T)\to\mathbb{R}^{n}$ is defined by (27). Then for every $y\in\mathbb{R}^{n}$ and all $t,t_{1},t_{2}\in T$ the following two identities hold true.
$\bigl{\langle}\delta_{t},\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\bigr{\rangle}_{C(T)^{\prime},C(T)}$ $=m(t)+K_{t,S}K_{S,S}^{\dagger}(y-m(S))\,,$
$\bigl{\langle}\delta_{t_{1}},\bigr{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)\bigr{)}(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}$ $=k(t_{1},t_{2})-K_{t_{1},S}K_{S,S}^{\dagger}K_{S,t_{2}}\,.$
Proof of Lemma 6.2: Using (B.3) the first identity follows from
$\bigl{\langle}\delta_{t},\mathbb{E}X+\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}(y-\mathbb{E}Y)\bigr{\rangle}_{C(T)^{\prime},C(T)}$ $=\mathbb{E}X_{t}+\bigl{(}\delta_{t}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(y-\mathbb{E}Y)$
$=m(t)+\bigl{(}\operatorname{cov}(\delta_{t}\circ X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(y-m(S))$
$=m(t)+\operatorname{cov}(\delta_{t}\circ X,Y)\bigl{(}K_{S,S}^{\dagger}(y-m(S))\bigr{)}$
$=m(t)+K_{t,S}K_{S,S}^{\dagger}(y-m(S))\,.$
For the proof of the second identity, we first observe that
$\Bigl{\langle}\delta_{t_{1}},\bigl{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)\bigr{)}(\delta_{t_{2}})\Bigr{\rangle}_{C(T)^{\prime},C(T)}$
$=\bigl{\langle}\delta_{t_{1}},\operatorname{cov}(X)(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}-\bigl{\langle}\delta_{t_{1}},\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}\,.$
Moreover, by (B.1) we obtain
$\bigl{\langle}\delta_{t_{1}},\operatorname{cov}(X)(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}=\operatorname{cov}\bigl{(}\langle\delta_{t_{1}},X\rangle_{C(T)^{\prime},C(T)},\langle\delta_{t_{2}},X\rangle_{C(T)^{\prime},C(T)}\bigr{)}=\operatorname{cov}(X_{t_{1}},X_{t_{2}})=k(t_{1},t_{2})\,.$
In addition, the definition of cross covariance operators gives
$\operatorname{cov}(Y,X)(\delta_{t_{2}})=\int_{\Omega}(X_{t_{2}}-\mathbb{E}X_{t_{2}})(Y-\mathbb{E}Y)\,\mathrm{d}\mathrm{P}=\operatorname{cov}(Y,X_{t_{2}})(1)=K_{S,t_{2}}\cdot 1=K_{S,t_{2}}\,,$
and hence we find
$\bigl{\langle}\delta_{t_{1}},\operatorname{cov}(X,Y)(\operatorname{cov}Y)^{\dagger}\operatorname{cov}(Y,X)(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}$ $=\bigl{(}\delta_{t_{1}}\circ\operatorname{cov}(X,Y)\circ(\operatorname{cov}Y)^{\dagger}\circ\operatorname{cov}(Y,X)\bigr{)}(\delta_{t_{2}})$
$=\bigl{(}\operatorname{cov}(\delta_{t_{1}}\circ X,Y)\circ(\operatorname{cov}Y)^{\dagger}\circ\operatorname{cov}(Y,X)\bigr{)}(\delta_{t_{2}})$
$=\bigl{(}\operatorname{cov}(\delta_{t_{1}}\circ X,Y)\circ(\operatorname{cov}Y)^{\dagger}\bigr{)}(K_{S,t_{2}})$
$=\operatorname{cov}(\delta_{t_{1}}\circ X,Y)\bigl{(}K_{S,S}^{\dagger}K_{S,t_{2}}\bigr{)}$
$=K_{t_{1},S}K_{S,S}^{\dagger}K_{S,t_{2}}\,.$
Combining these considerations yields the second identity. ∎
Proof of Theorem 4.6: In view of Theorem 3.1 it suffices to translate the formulas for $\mu_{X|Y}$ and $C_{X|Y}$ to the claimed formulas for the mean and covariance function via (24), respectively (25). This, however, is straightforward by using Lemma 6.2. ∎
Proof of Proposition 4.7: Since $T_{\infty}$ is dense in $T$, every $f\in C(T)$ is uniquely determined by its behavior on $T_{\infty}$. Consequently, the sequence $(\delta_{t_{i}})$ is separating. Now the assertion follows from Propositions 3.3 and 4.2
Proof of Theorem 4.8: We already know by Proposition 4.7 that the operators $I_{n}:C(S)\to\mathbb{R}$ defined by
$I_{n}(g):=\bigl{(}g(s_{1}),\dots,g(s_{n})\bigr{)}=g(S_{n})\,,\qquad\qquad g\in C(S),$
give us a filtering sequence $(I_{n})$ for $F:=C(S)$. We define $Y_{n}:=I_{n}\circ Y$. By $S_{n}\subset S$ we then obtain
$Y_{n}=I_{n}\circ Y=I_{S_{n}}\circ X\,,$
where $I_{S_{n}}:C(T)\to\mathbb{R}^{n}$ is defined by (27). Applying Theorem 3.4 with the spaces $E:=C(T)$ and $F:=C(S)$ gives us a $\mathrm{P}_{Y}$-zero set $N\in\mathfrak{B}(C(S))$ such that identities parts i) to viii) hold true. In particular, we know that $\mathrm{P}_{X|Y}(\,\cdot\,|g)$ is a Gaussian measure on $C(T)$ for all $g\in C(S)\setminus N$.
To compute its mean and covariance function we pick a $g\in C(S)\setminus N$. Then, part i) of Theorem 3.4 together with Lemma 6.2 shows
$\bigl{\langle}\delta_{t},\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(I_{n}(g)-\mathbb{E}Y_{n})\bigr{\rangle}_{C(T)^{\prime},C(T)}=m(t)+K_{t,S_{n}}K_{S_{n},S_{n}}^{\dagger}(g(S_{n})-m(S_{n}))$
and
$\bigl{\langle}\delta_{t_{1}},\bigr{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(\delta_{t_{2}})\bigr{\rangle}_{C(T)^{\prime},C(T)}=K_{t_{1},S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t_{2}}$
for all $t,t_{1},t_{2}\in T$. Combining part iii) of Theorem 3.4 with (24) we thus obtain
$\sup_{t\in T}\Bigl{|}m(t)+K_{t,S_{n}}K_{S_{n},S_{n}}^{\dagger}\bigl{(}g(S_{n})-m(S_{n})\bigr{)}-m_{X|Y=g}(t)\Bigr{|}$
$=\sup_{t\in T}\Bigl{|}\bigl{\langle}\delta_{t},\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(I_{n}(g)-\mathbb{E}Y_{n})\bigr{\rangle}_{C(T)^{\prime},C(T)}-\langle\delta_{t},\mu_{X|Y}(g)\rangle_{C(T)^{\prime},C(T)}\Bigr{|}$
$=\sup_{t\in T}\Bigl{|}\bigl{\langle}\delta_{t},\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(I_{n}(g)-\mathbb{E}Y_{n})-\mu_{X|Y}(g)\bigr{\rangle}_{C(T)^{\prime},C(T)}\Bigr{|}$
$=\bigl{\|}\,\mathbb{E}X+\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}(I_{n}(g)-\mathbb{E}Y_{n})-\mu_{X|Y}(g)\,\bigr{\|}_{C(T)}$
$\to 0\,,$
where in the second to last step we used that the norm of $C(T)$ is the supremums norm.
The formula for the covariance function can be derived similarly: Indeed, combining part iv) of Theorem 3.4 with (25), we obtain
$\sup_{t_{1},t_{2}\in T}\Bigl{|}k(t_{1},t_{2})-K_{t_{1},S_{n}}K_{S_{n},S_{n}}^{\dagger}K_{S_{n},t_{2}}-k_{X|Y}(t_{1},t_{2})\Bigr{|}$
$=\sup_{t_{1},t_{2}\in T}\Bigl{|}\Bigl{\langle}\delta_{t_{1}},\bigl{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)\bigr{)}(\delta_{t_{2}})\Bigr{\rangle}_{C(T)^{\prime},C(T)}-\langle\delta_{t_{1}},C_{X|Y}\delta_{t_{2}}\rangle_{C(T)^{\prime},C(T)}\Bigr{|}$
$=\sup_{t_{1},t_{2}\in T}\Bigl{|}\Bigl{\langle}\delta_{t_{1}},\bigl{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)-C_{X|Y}\bigr{)}(\delta_{t_{2}})\Bigr{\rangle}_{C(T)^{\prime},C(T)}\Bigr{|}$
$=\sup_{t_{2}\in T}\bigl{\|}\,\bigl{(}\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)-C_{X|Y}\bigr{)}(\delta_{t_{2}})\,\bigr{\|}_{C(T)}$
$\leq\bigl{\|}\,\operatorname{cov}(X)-\operatorname{cov}(X,Y_{n})(\operatorname{cov}Y_{n})^{\dagger}\operatorname{cov}(Y_{n},X)-C_{X|Y}\,\bigr{\|}$
$\to 0\,,$
where we used that the norm of $C(T)$ is the supremums norm and that $\delta_{t_{2}}\in B_{C(T)^{\prime}}$ for all $t_{2}\in T$.
To prove the last assertion, we note that $\langle\delta_{s},X\rangle_{C(T)^{\prime},C(T)}=X_{s}=\langle\delta_{s},Y\rangle_{C(S)^{\prime},C(S)}$. Consequently, the random variable $\langle\delta_{s},X\rangle_{C(T)^{\prime},C(T)}$ is $\sigma(Y)$-measurable. If we now pick an $\omega\in Y^{-1}(\{g\})$, then we have $\omega\not\in Y^{-1}(N)$ and $g=Y(\omega)=X_{|S}(\omega)$, and therefore part viii) of Theorem 3.4 gives
$m_{X|Y=g}(s)=\langle\delta_{s},\mu_{X|Y}(g)\rangle_{C(T)^{\prime},C(T)}=\bigl{\langle}\delta_{s},\mu_{X|Y}(Y(\omega))\bigr{\rangle}_{C(T)^{\prime},C(T)}=\bigl{\langle}\delta_{s},X(\omega)\bigr{\rangle}_{C(T)^{\prime},C(T)}=g(s)\,.$
Analogously, part viii) of Theorem 3.4 shows $k_{X|Y}(s,t)=\langle\delta_{t},C_{X|Y}\delta_{s}\rangle_{C(T)^{\prime},C(T)}=0$ for all $t\in T$. ∎
A Appendix: Banach Space Valued Integration
Let $(\Omega,{\cal A})$ be a measurable space, $E$ be a Banach space, and $X:\Omega\to E$. It is well known, that $X$ is strongly measurable if and only if it is weakly measurable and $X(\Omega)$ is separable, see e.g. [10, Theorem E.9] or [21, Theorem 1.1.6]. In addition, if $X$ is strongly measurable, then there exists a sequence $(X_{n})$ of (strongly) measurable step functions $X_{n}:\Omega\to E$ with
$\lim_{n\to\infty}X_{n}(\omega)$ $=X(\omega)\,,$ (A.1)
$\|X_{n}(\omega)\|_{E}$ $\leq\|X(\omega)\|_{E}$ (A.2)
for all all $\omega\in\Omega$ and $n\geq 1$, see e.g. [10, Proposition E.2]. Here we note that step functions are also called simple functions. In addition, for step functions all three notions of measurability coincide, and for this reason, we usually speak of measurable step functions. Finally, every pointwise limit $X$ of a sequence $(X_{n})$ of strongly measurable functions is strongly measurable, see e.g. [10, Proposition E.1]. Consequently, we can characterize strong measurability by pointwise approximations with measurable step functions in the sense of (A.1). In particular, the finite sum of strongly measurable functions is strongly measurable and if $f:\Omega\to\mathbb{R}$ is measurable and $X:\Omega\to E$ is strongly measurable, then $f\cdot X:\Omega\to E$ is strongly measurable.
If we have a probability measure $\mathrm{P}$ on $(\Omega,{\cal A})$ and a measurable step function $X:\Omega\to E$ with some fixed representation $X=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}x_{i}$ with $A_{i}\in{\cal A}$ and $x_{i}\in E$, then
$\int_{\Omega}X\,\mathrm{d}\mathrm{P}:=\sum_{i=1}^{m}\mathrm{P}(A_{i})x_{i}$ (A.3)
is actually independent of the chosen representation, where the proof is analogous to the $\mathbb{R}$-valued case. To extend this integral we say that a strongly measurable $X:\Omega\to E$ is Bochner integrable with respect to $\mathrm{P}$, if
$\int_{\Omega}\|X(\omega)\|_{E}\,\mathrm{d}\mathrm{P}(\omega)<\infty\,.$
In this case, using approximating sequences described in (A.1) and (A.2), one can define the Bochner integral by
$\mathbb{E}X:=\int_{\Omega}X\,\mathrm{d}\mathrm{P}:=\lim_{n\to\infty}\int_{\Omega}X_{n}\,\mathrm{d}\mathrm{P}\,,$ (A.4)
where the integral in the middle is independent of the chosen approximation $(X_{n})$, see e.g. [10, page 399]. Moreover, the limit is, of course, with respect to the norm $\|\cdot\|_{E}$. The set ${{\cal L}_{1}(\mathrm{P},E)}$ of all Bochner integrable $X:\Omega\to E$ is a vector space, see e.g. [10, Proposition E.4], which obviously contains all measurable step functions. Moreover, it is easy to see that
$\|X\|_{{{\cal L}_{1}(\mathrm{P},E)}}:=\int_{\Omega}\|X(\omega)\|_{E}\,\mathrm{d}\mathrm{P}(\omega)$ (A.5)
defines a semi-norm on ${{\cal L}_{1}(\mathrm{P},E)}$. In addition, the Bochner integral defines a linear map ${{\cal L}_{1}(\mathrm{P},E)}\to E$ that is also bounded since we have
$\Bigl{\|}\,\int_{\Omega}X\,\mathrm{d}\mathrm{P}\,\Bigr{\|}_{E}\leq\int_{\Omega}\|X(\omega)\|_{E}\,\mathrm{d}\mathrm{P}(\omega)\,,\qquad\qquad X\in{{\cal L}_{1}(\mathrm{P},E)}\,,$ (A.6)
see e.g. [10, Proposition E.5]. As usual, we write ${L_{1}(\mathrm{P},E)}$ for the space of $\mathrm{P}$-equivalence classes, which becomes a Banach space if we modify the norm above in the usual way, see e.g. [10, page 401].
The Bochner integral also enjoys a form of the dominated convergence theorem. Namely, if we have a sequence $(X_{n})$ of strongly measurable $X_{n}:\Omega\to E$ for which there exist both an $X:\Omega\to E$ and a $\mathrm{P}$-integrable $g:\Omega\to[0,\infty]$ such that
$\lim_{n\to\infty}X_{n}(\omega)$ $=X(\omega)\,,$ (A.7)
$\|X_{n}(\omega)\|_{E}$ $\leq\|g(\omega)\|_{E}$ (A.8)
for all all $\omega\in\Omega$ and $n\geq 1$, then $X$ and all $X_{n}$ are Bochner $\mathrm{P}$-integrable, and we have
$\lim_{n\to\infty}\int_{\Omega}X_{n}\,\mathrm{d}\mathrm{P}=\int_{\Omega}X\,\mathrm{d}\mathrm{P}\,.$ (A.9)
For a proof, we refer to e.g. [10, Proposition E.6], where we note that the strong measurability of $X$ follows from e.g. [10, Proposition E.1].
Moreover, if $T:E\to F$ is a bounded linear operator mapping into another Banach space $F$ and $X\in{{\cal L}_{1}(\mathrm{P},E)}$, then the map $T\circ X:\Omega\to F$ satisfies both $T\circ X\in{{\cal L}_{1}(\mathrm{P},F)}$ and
$\int_{\Omega}T\circ X\,\mathrm{d}\mathrm{P}=T\biggl{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggr{)}\,,$ (A.10)
see e.g. [13, Theorem 6 in Chapter II.2] or [21, Theorem 1.2.4] for a stronger result in this direction that goes back to E. Hille. In particular, for all $X\in{{\cal L}_{1}(\mathrm{P},E)}$, we have both $\langle x^{\prime},X\rangle_{E^{\prime},E}\in{{\cal L}_{1}(\mathrm{P})}$ and the “Pettis property”
$\int_{\Omega}\langle x^{\prime},X\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}=\biggl{\langle}x^{\prime},\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggr{\rangle}_{E^{\prime},E}\,,\qquad\qquad x^{\prime}\in E^{\prime}\,.$ (A.11)
This shows that for $X\in{{\cal L}_{1}(\mathrm{P},E)}$ the Bochner integral coincides with the so-called Pettis integral, which is discussed in e.g. [21, Chapter 1.2.c].
Similarly to ${{\cal L}_{1}(\mathrm{P},E)}$ one can define, for $p\in(1,\infty)$, the vector space ${{\cal L}_{p}(\mathrm{P},E)}$ as the set of all strongly measurable $X:\Omega\to E$ with
$\|X\|_{{{\cal L}_{p}(\mathrm{P},E)}}^{p}:=\int_{\Omega}\|X(\omega)\|_{E}^{p}\,\mathrm{d}\mathrm{P}(\omega)<\infty.$
Again this defines a semi-norm on ${{\cal L}_{p}(\mathrm{P},E)}$ and ${L_{p}(\mathrm{P},E)}$ denotes the Banach space of corresponding $\mathrm{P}$-equivalence classes.
B Appendix: Covariance Operators
Let us fix a probability space $(\Omega,{\cal A},\mathrm{P})$ and Banach spaces $E$ and $F$. Moreover, let us fix some $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},F)}$, as well as a $y^{\prime}\in F^{\prime}$. Since $X$ and $Y$ are strongly measurable, the map
$\langle y^{\prime},Y-\mathbb{E}Y\rangle_{F^{\prime},F}\cdot(X-\mathbb{E}X):\Omega\to E$
is also strongly measurable, see the discussion following (A.2). In addition, the Cauchy-Schwarz inequality shows that it is also Bochner integrable. Moreover, the resulting linear operator $\operatorname{cov}(X,Y):F^{\prime}\to E$ defined by
$\operatorname{cov}(X,Y)y^{\prime}=\int_{\Omega}\langle y^{\prime},Y-\mathbb{E}Y\rangle_{F^{\prime},F}\,(X-\mathbb{E}X)\,\mathrm{d}\mathrm{P}\,,\qquad\qquad y^{\prime}\in F^{\prime}$
is bounded by virtue of the norm-inequality for Bochner integrals (A.6) and another application of the Cauchy-Schwarz inequality. Applying the “Pettis property” (A.11) then gives
$\langle x^{\prime},\operatorname{cov}(X,Y)y^{\prime}\rangle_{E^{\prime},E}$ $=\int_{\Omega}\langle y^{\prime},Y-\mathbb{E}Y\rangle_{F^{\prime},F}\,\langle x^{\prime},X-\mathbb{E}X\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}$ (B.1)
$=\operatorname{cov}\bigl{(}\langle y^{\prime},Y\rangle_{F^{\prime},F},\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{)}\,.$
In other words, the cross covariance operator provides the covariances between all pairs of one-dimensional projections of $X$ and $Y$. Another routine check shows that the dual is given by
$\operatorname{cov}(X,Y)^{\prime}=\iota_{F}\circ\operatorname{cov}(Y,X)\,,$ (B.2)
where $\iota_{F}:F\to F^{\prime\prime}$ denotes the canonical embedding of $F$ into its bidual $F^{\prime\prime}$. In addition, if $S:E\to\tilde{E}$ and $T:F\to\tilde{F}$ are bounded linear operators and we have $x\in E$ and $y\in F$, then the following identities hold true:
$\operatorname{cov}(S\circ X,T\circ Y)$ $=S\circ\operatorname{cov}(X,Y)\circ T^{\prime}\,,$ (B.3)
$\operatorname{cov}(x+X,y+Y)$ $=\operatorname{cov}(X,Y)\,.$ (B.4)
Indeed, (B.3) follows by applying (A.10) with the operator $S$ and using the definition of $T^{\prime}$ and (B.4) is a direct consequence of the definition of $\operatorname{cov}(X,Y)$.
In the case $X=Y$, the covariance operator $\operatorname{cov}(X):=\operatorname{cov}(X,X)$ is an operator $\operatorname{cov}(X):E^{\prime}\to E$ and with (B.1) we find
$\langle x^{\prime}_{1},\operatorname{cov}(X)x^{\prime}_{2}\rangle_{E^{\prime},E}$ $=\langle x^{\prime}_{2},\operatorname{cov}(X)x^{\prime}_{1}\rangle_{E^{\prime},E}\,,$ (B.5)
$\langle x^{\prime},\operatorname{cov}(X)x^{\prime}\rangle_{E^{\prime},E}$ $=\operatorname{var}\bigl{(}\langle x^{\prime},X\rangle_{E^{\prime},E}\bigr{)}\geq 0$ (B.6)
for all $x^{\prime},x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$, where $\operatorname{var}(Z)$ denotes the variance of an $\mathbb{R}$-valued random variable $Z$. This motivates the following definition.
Definition B.1.
Let $E$ be a Banach space. A bounded linear operator $T:E^{\prime}\to E$ is called an abstract covariance operator if it is symmetric and non-negative, that is, if for all $x^{\prime},x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$, we have
$\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}=\langle x_{2}^{\prime},Tx_{1}^{\prime}\rangle_{E^{\prime},E}\qquad\qquad\mbox{and }\qquad\qquad\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}\geq 0\,.$
Analogously, a bounded linear operator $S:E\to E^{\prime}$ is symmetric and non-negative, if for all $x,x_{1},x_{2}\in E$ we have
$\langle Sx_{1},x_{2}\rangle_{E^{\prime},E}=\langle Sx_{2},x_{1}\rangle_{E^{\prime},E}\qquad\qquad\mbox{ and }\qquad\qquad\langle Sx,x\rangle_{E^{\prime},E}\geq 0\,.$
As discussed above, every covariance operator $\operatorname{cov}(X)$ is an abstract covariance operator. In addition, an elementary calculation shows that the sum $T_{1}+T_{2}:E^{\prime}\to E$ of two abstract covariance operators $T_{1},T_{2}:E^{\prime}\to E$ is an abstract covariance operator. Moreover, every abstract covariance operator $T:E^{\prime}\to E$ satisfies
$\langle x_{1}^{\prime}+x_{2}^{\prime},T(x_{1}^{\prime}+x_{2}^{\prime})\rangle_{E^{\prime},E}=\langle x_{1}^{\prime},Tx_{1}^{\prime}\rangle_{E^{\prime},E}+2\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}+\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}\,,$
which in turn implies
$\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}=\frac{\langle x_{1}^{\prime}+x_{2}^{\prime},T(x_{1}^{\prime}+x_{2}^{\prime})\rangle_{E^{\prime},E}-\langle x_{1}^{\prime},Tx_{1}^{\prime}\rangle_{E^{\prime},E}-\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}}{2}$ (B.7)
for all $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$. The following lemma uses this observation to show that abstract covariance operators are determined by the behavior of $x^{\prime}\mapsto\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}$.
Lemma B.2.
Let $E$ be a Banach space and $S,T:E^{\prime}\to E$ be two abstract covariance operators with
$\langle x^{\prime},Tx^{\prime}\rangle_{E^{\prime},E}=\langle x^{\prime},Sx^{\prime}\rangle_{E^{\prime},E}\,,\qquad\qquad x^{\prime}\in E^{\prime}.$
Then we have $S=T$.
Proof of Lemma B.2: Let us fix some $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$. By (B.7) and the assumed identity on the diagonal we then know $\langle x_{1}^{\prime},Sx_{2}^{\prime}\rangle_{E^{\prime},E}=\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}$. This implies $Sx_{2}^{\prime}=Tx_{2}^{\prime}$ for all $x^{\prime}_{2}\in E^{\prime}$ by Hahn-Banach’s Theorem. Consequently we have $S=T$. ∎
In the finite dimensional case $E=\mathbb{R}^{n}$, one can quickly check that a linear operator $T:(\mathbb{R}^{n})^{\prime}\to\mathbb{R}^{n}$ is an abstract covariance operator, if and only if its representing $n\times n$-matrix $A$ with respect to the canonical bases of $(\mathbb{R}^{n})^{\prime}$ and $\mathbb{R}^{n}$ is symmetric and positive semi-definite. In this case, the spectral theorem for matrices shows $A=S\Lambda S^{\mathsf{T}}$, where $\Lambda$ is a diagonal $n\times n$-matrix given by the eigenvalues $\lambda_{1},\dots,\lambda_{n}$ of $A$ and $S$ is an orthogonal $n\times n$-matrix $S$. Moreover, we have
$A^{\dagger}=(S\Lambda S^{\mathsf{T}})^{\dagger}=S\Lambda^{\dagger}S^{\mathsf{T}}\,,$ (B.8)
where the second identity can be found in e.g. [48, Theorem 1.1.2]. Moreover, a straightforward calculation shows that $\Lambda^{\dagger}$ is the diagonal $n\times n$-matrix having the entries $\lambda_{1}^{-1},\dots,\lambda_{n}^{-1}$ with $00^{-1}:=$ 0. Consequently, $A^{\dagger}$ is also symmetric and positive semi-definite. Moreover, this matrix describes the operator $T^{\dagger}:\mathbb{R}^{n}\to(\mathbb{R}^{n})^{\prime}$, and therefore $T^{\dagger}$ is symmetric and non-negative.
It can be quickly verified that an abstract covariance operator $T:E^{\prime}\to E$ defines a (reproducing) kernel $k_{T}:E^{\prime}\times E^{\prime}\to\mathbb{R}$ by
$k_{T}(x_{1}^{\prime},x_{2}^{\prime}):=\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}\,,\qquad\qquad x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}\,.$
Lemma B.2 shows that this kernel is determined by its behaviour on the diagonal. Moreover, the Cauchy-Schwarz inequality for kernels, see e.g. [38, (4.14)] yields
$|\langle x_{1}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}|^{2}=|k_{T}(x_{1}^{\prime},x_{2}^{\prime})|^{2}\leq k_{T}(x_{1}^{\prime},x_{1}^{\prime})\cdot k_{T}(x_{2}^{\prime},x_{2}^{\prime})=\langle x_{1}^{\prime},Tx_{1}^{\prime}\rangle_{E^{\prime},E}\cdot\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}\,.$ (B.9)
In particular, $\langle x_{2}^{\prime},Tx_{2}^{\prime}\rangle_{E^{\prime},E}=0$ implies $Tx_{2}^{\prime}=0$ by a simple application of Hahn-Banach’s theorem.
The next lemma collects some additional information about this kernel $k_{T}$ in the case of covariance operators $T=\operatorname{cov}(X)$.
Theorem B.3.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ be a separable Banach space, $X\in{{\cal L}_{2}(\mathrm{P},E)}$, and $k_{X}:=k_{\operatorname{cov}(X)}:E^{\prime}\times E^{\prime}\to\mathbb{R}$ be the resulting kernel. Then the restriction $(k_{X})_{|B_{E^{\prime}}\times B_{E^{\prime}}}$ of $k_{X}$ onto $B_{E^{\prime}}\times B_{E^{\prime}}$ is $w^{*}$-continuous.
Proof of Theorem B.3: Since $(B_{E^{\prime}},w^{*})$ is metrizable, see Theorem 2.1, it suffices to show that the restricted kernel is $w^{*}$-sequentially continuous. Moreover, by (B.7) it suffices to show that the kernel restricted to $2B_{E^{\prime}}$ is $w^{*}$-sequentially continuous on the diagonal. Finally, by a simple scaling argument, it actually suffices to verify that $x^{\prime}\mapsto k_{X}(x^{\prime},x^{\prime})$ is $w^{*}$-sequentially continuous on $B_{E^{\prime}}$.
Let us therefore fix some $x^{\prime}_{0}\in B_{E^{\prime}}$ as well as a sequence $(x_{n}^{\prime})\subset B_{E^{\prime}}$ such that $x_{n}^{\prime}\to x^{\prime}_{0}$ with respect to $w^{*}$. Moreover, we write $Z:=X-\mathbb{E}X$. By (B.4) and (B.1) we then find
$\langle x_{n}^{\prime},\operatorname{cov}(X)x_{n}^{\prime}\rangle_{E^{\prime},E}=\langle x_{n}^{\prime},\operatorname{cov}(Z)x_{n}^{\prime}\rangle_{E^{\prime},E}$ $=\int_{\Omega}\langle x_{n}^{\prime},Z\rangle_{E^{\prime},E}\cdot\langle x_{n}^{\prime},Z\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}$ (B.10)
for all $n\geq 0$. For $n\geq 0$ we define $g_{n}:\Omega\to[0,\infty)$ by
$g_{n}(\omega):=\langle x_{n}^{\prime},Z(\omega)\rangle_{E^{\prime},E}^{2}\,,\qquad\qquad\omega\in\Omega.$
By the assumed $w^{*}$-convergence $x_{n}^{\prime}\to x^{\prime}_{0}$ we then see that $g_{n}(\omega)\to g_{0}(\omega)$ for all $\omega\in\Omega$. Moreover, a simple estimate yields
$|g_{n}(\omega)|\leq\|x_{n}^{\prime}\|_{E^{\prime}}^{2}\cdot\|Z(\omega)\|_{E}^{2}\leq\|Z(\omega)\|^{2}_{E}\,,\qquad\qquad\omega\in\Omega.$
Using $Z\in{{\cal L}_{2}(\mathrm{P},E)}$, Lebesgue’s dominated convergence theorem, and (B.10), we then find
$\langle x_{n}^{\prime},\operatorname{cov}(X)x_{n}^{\prime}\rangle_{E^{\prime},E}\to\langle x_{0}^{\prime},\operatorname{cov}(X)x_{0}^{\prime}\rangle_{E^{\prime},E}\,,$
that is, we have shown the assertion. ∎
Lemma B.4.
Let $E$ and $F$ be separable Banach spaces, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $X:\Omega\to E$ and $Y:\Omega\to F$ be centered Gaussian random variables, and $S:F\to F^{\prime}$ be a bounded, linear, symmetric, and non-negative operator. Then the operator $K:E^{\prime}\to E$ defined by
$K:=\operatorname{cov}(X,Y)S\operatorname{cov}(Y,X)\,.$
is an abstract covariance operator, and for all $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$ we have
$\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}=\int_{\Omega\times\Omega}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\cdot\langle SY(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\,\mathrm{d}\mathrm{P}^{2}(\omega_{1},\omega_{2})\,.$
Moreover, the map $B_{E^{\prime}}\to\mathbb{R}$ defined by $x^{\prime}\mapsto\langle x^{\prime},Kx^{\prime}\rangle_{E^{\prime},E}$ is $w^{*}$-continuous.
Proof of Lemma B.4: Obviously, $K$ is bounded and linear. To verify that $K$ is an abstract covariance operator, we fix some $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$ and write $y_{i}:=\operatorname{cov}(Y,X)x_{i}^{\prime}\in F$. With the help of (B.2) we then obtain
$\langle x_{1}^{\prime},Kx_{2}^{\prime}\rangle_{E^{\prime},E}=\langle x_{1}^{\prime},\operatorname{cov}(X,Y)Sy_{2}\rangle_{E^{\prime},E}=\langle\operatorname{cov}(X,Y)^{\prime}x_{1}^{\prime},Sy_{2}\rangle_{F^{\prime\prime},F^{\prime}}$ $=\langle\iota_{F}\operatorname{cov}(Y,X)x_{1}^{\prime},Sy_{2}\rangle_{F^{\prime\prime},F^{\prime}}$
$=\langle Sy_{2},y_{1}\rangle_{F^{\prime},F}\,.$
Since we assumed that $S$ is symmetric and non-negative, we can now quickly verify that this is also true for $K$.
To establish the integral formula, we again fix some $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$. Then the definition of $\operatorname{cov}(Y,X)$ gives
$y^{\prime}_{1}:=S\operatorname{cov}(Y,X)x^{\prime}_{1}=S\biggl{(}\int_{\Omega}\langle x_{1}^{\prime},X\rangle_{E^{\prime},E}\cdot Y\,\mathrm{d}\mathrm{P}\biggr{)}=\int_{\Omega}\langle x_{1}^{\prime},X\rangle_{E^{\prime},E}\cdot(S\circ Y)\,\mathrm{d}\mathrm{P}\,,$
where in the last step we used (A.10). Another routine calculation for $y\in F$ thus yields
$\langle y^{\prime}_{1},y\rangle_{F^{\prime},F}=\langle\iota_{F}y,y^{\prime}_{1}\rangle_{F^{\prime\prime},F^{\prime}}$ $=\biggl{\langle}\iota_{F}y,\int_{\Omega}\langle x_{1}^{\prime},X\rangle_{E^{\prime},E}\cdot(S\circ Y)\,\mathrm{d}\mathrm{P}\biggr{\rangle}_{F^{\prime\prime},F^{\prime}}$
$=\int_{\Omega}\langle x_{1}^{\prime},X\rangle_{E^{\prime},E}\cdot\langle\iota_{F}y,S\circ Y\rangle_{F^{\prime\prime},F^{\prime}}\,\mathrm{d}\mathrm{P}$
$=\int_{\Omega}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),y\rangle_{F^{\prime},F}\,\mathrm{d}\mathrm{P}(\omega_{1})\,,$
where we used the “Pettis property” (A.11) in the third step. Combining the last identity with the definition of $\operatorname{cov}(X,Y)$ and (A.11), we then find
$\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$ $=\langle x_{2}^{\prime},\operatorname{cov}(X,Y)y_{1}^{\prime}\rangle_{E^{\prime},E}$
$=\biggl{\langle}x_{2}^{\prime},\int_{\Omega}\langle y^{\prime}_{1},Y(\omega_{2})\rangle_{F^{\prime},F}\cdot X(\omega_{2})\,\mathrm{d}\mathrm{P}(\omega_{2})\biggr{\rangle}_{E^{\prime},E}$
$=\int_{\Omega}\langle y^{\prime}_{1},Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}(\omega_{2})$
$=\int_{\Omega}\int_{\Omega}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\,\mathrm{d}\mathrm{P}(\omega_{1})\,\mathrm{d}\mathrm{P}(\omega_{2})\,.$
Moreover, we have
$g(\omega_{1},\omega_{2})$ $:=\Bigl{|}\langle x_{1}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{2}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}\Bigr{|}$
$\leq\|x_{1}^{\prime}\|\cdot\|x_{2}^{\prime}\|\cdot\|S\|\cdot\|X(\omega_{1})\|_{E}\cdot\|Y(\omega_{1})\|_{F}\cdot\|Y(\omega_{2})\|_{F}\cdot\|X(\omega_{2})\|_{E}$ (B.11)
for all $\omega_{1},\omega_{2}\in\Omega$. Now Fernique’s theorem ensures $\|X\|_{E},\|Y\|_{F}\in{{\cal L}_{2}(\mathrm{P})}$, and therefore we find $\|X\|_{E}\cdot\|Y\|_{F}\in{{\cal L}_{1}(\mathrm{P})}$ by the Cauchy-Schwarz inequality. This gives
$\int_{\Omega}\int_{\Omega}|g(\omega_{1},\omega_{2})|\,\mathrm{d}\mathrm{P}(\omega_{1})\,\mathrm{d}\mathrm{P}(\omega_{2})$
$\leq\|x_{1}^{\prime}\|\cdot\|x_{2}^{\prime}\|\cdot\|S\|\int_{\Omega}\|X(\omega_{1})\|_{E}\cdot\|Y(\omega_{1})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{1})\cdot\int_{\Omega}\|X(\omega_{2})\|_{E}\cdot\|Y(\omega_{2})\|_{F}\,\mathrm{d}\mathrm{P}(\omega_{2})$
[…]
and therefore Fubini’s theorem yields the formula for $\langle x_{2}^{\prime},Kx_{1}^{\prime}\rangle_{E^{\prime},E}$.
Let us now recall from Theorem 2.1 that $(B_{E^{\prime}},w^{*})$ is a compact metrizable space since we assumed that $E$ is separable. Consequently, it suffices to show that $x^{\prime}\mapsto\langle x^{\prime},Kx^{\prime}\rangle_{E^{\prime},E}$ is $w^{*}$-sequentially continuous. To this end, we fix an $x^{\prime}_{0}\in B_{E^{\prime}}$ and a sequence $(x_{n}^{\prime})\subset B_{E^{\prime}}$ that converges to $x^{\prime}_{0}$ with respect to $w^{*}$, that is, $\langle x_{n}^{\prime},x\rangle_{E^{\prime},E}\to\langle x_{0}^{\prime},x\rangle_{E^{\prime},E}$ for all $x\in E$. We write
$f_{n}(\omega_{1},\omega_{2}):=\langle x_{n}^{\prime},X(\omega_{1})\rangle_{E^{\prime},E}\cdot\langle S\circ Y(\omega_{1}),Y(\omega_{2})\rangle_{F^{\prime},F}\cdot\langle x_{n}^{\prime},X(\omega_{2})\rangle_{E^{\prime},E}$
for all $n\geq 0$ and $\omega_{1},\omega_{2}\in\Omega$. Then we have $f_{n}\to f_{0}$ pointwise and (B.11) shows
$|f_{n}(\omega_{1},\omega_{2})|\leq\|S\|\cdot\|X(\omega_{1})\|\cdot\|Y(\omega_{1})\|\cdot\|Y(\omega_{2})\|\cdot\|X(\omega_{2})\|$
for all $n\geq 0$ and $\omega_{1},\omega_{2}\in\Omega$. Since this upper bound of $f_{n}$ is $\mathrm{P}\otimes\mathrm{P}$-integrable by (B.12) and Fubini’s theorem, we then find
$\langle x_{n}^{\prime},Kx_{n}^{\prime}\rangle_{E^{\prime},E}=\int_{\Omega\times\Omega}f_{n}\,\mathrm{d}\mathrm{P}^{2}\to\int_{\Omega\times\Omega}f_{0}\,\mathrm{d}\mathrm{P}^{2}=\langle x_{0}^{\prime},Kx_{0}^{\prime}\rangle_{E^{\prime},E}$
by the already established formula for $\langle x_{n}^{\prime},Kx_{n}^{\prime}\rangle$ and Lebesgue’s dominated convergence theorem. ∎
So far, we have defined the cross covariance operator $\operatorname{cov}(Y,X):E^{\prime}\to F$ with the help of an $F$-valued Bochner integral that is defined pointwise, i.e. for all $x^{\prime}\in E^{\prime}$. Our next goal is to show that the cross covariance operator can also be expressed by an operator valued Bochner integral. Here we note that in the Hilbert space case such a representation can sometimes be found in the literature, while in the general Banach space case we could not find such representation. For this reason, we explain its construction in some detail.
Let us begin by quickly recalling some basic facts about tensor products: Let us fix some Banach spaces $E$ and $F$. For $x^{\prime}\in E^{\prime}$ and $y\in F$, we define the bounded linear operator $x^{\prime}\otimes y:E\to F$ by
$(x^{\prime}\otimes y)(x):=\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\,,\qquad\qquad x\in E\,.$
A simple calculation then shows that
$\|x^{\prime}\otimes y\|=\sup_{x\in B_{E}}\|\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\|_{F}=\sup_{x\in B_{E}}|\langle x^{\prime},x\rangle_{E^{\prime},E}|\cdot\|y\|_{F}=\|x^{\prime}\|_{E^{\prime}}\cdot\|y\|_{F}\,.$ (B.13)
As usual, $x^{\prime}\otimes y$ is called an elementary tensor. Elementary tensors can be added and a quick calculation shows bilinearity, or more precisely
$x^{\prime}_{1}\otimes y+x^{\prime}_{2}\otimes y$ $=(x^{\prime}_{1}+x^{\prime}_{2})\otimes y\,,$ $\alpha\cdot(x^{\prime}\otimes y)$ $=(\alpha x^{\prime})\otimes y\,,$
$x^{\prime}\otimes y_{1}+x^{\prime}\otimes y_{2}$ $=x^{\prime}\otimes(y_{1}+y_{2})\,,$ $\beta\cdot(x^{\prime}\otimes y)$ $=x^{\prime}\otimes(\beta y)$
for all $x^{\prime},x^{\prime}_{1},x^{\prime}_{2}\in E$, $y,y_{1},y_{2}\in F$, and $\alpha,\beta\in\mathbb{R}$, where, as usual, we assumed that $\otimes$ has a higher priority than $+$. We define the algebraic tensor product space by
$E^{\prime}\otimes F:=\operatorname{span}\{x^{\prime}\otimes y:x^{\prime}\in E^{\prime},y\in F\}\,.$
It is easy to see that this space equals the space of all bounded linear operators $A:E\to F$ with finite rank. Moreover, for $x^{\prime}\otimes y:E\to F$, its adjoint $(x^{\prime}\otimes y)^{\prime}:F^{\prime}\to E^{\prime}$ is given by
$(x^{\prime}\otimes y)^{\prime}=(\iota_{F}y)\otimes x^{\prime}\,.$ (B.14)
Indeed, for $x\in E$ and $y^{\prime}\in F^{\prime}$ we have $((\iota_{F}y)\otimes x^{\prime})(y^{\prime})=\langle\iota_{F}y,y^{\prime}\rangle_{F^{\prime\prime},F^{\prime}}\cdot x^{\prime}=\langle y^{\prime},y\rangle_{F^{\prime},F}\cdot x^{\prime}$, and hence we find
$\bigl{\langle}((\iota_{F}y)\otimes x^{\prime})(y^{\prime}),x\bigr{\rangle}_{E^{\prime},E}=\bigl{\langle}\langle y^{\prime},y\rangle_{F^{\prime},F}\cdot x^{\prime},x\bigr{\rangle}_{E^{\prime},E}=\langle y^{\prime},y\rangle_{F^{\prime},F}\cdot\langle x^{\prime},x\rangle_{E^{\prime},E}$ $=\bigl{\langle}y^{\prime},\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\bigr{\rangle}_{F^{\prime},F}$
$=\langle y^{\prime},(x^{\prime}\otimes y)(x)\rangle_{F^{\prime},F}\,.$
A bounded linear operator $A:E\to F$ is called nuclear, if there exist sequences $(x_{n}^{\prime})\subset E^{\prime}$ and $(y_{n})\subset F$ with
$\sum_{n=1}^{\infty}\|x_{n}^{\prime}\|_{E^{\prime}}\cdot\|y_{n}\|_{F}<\infty\qquad\qquad\mbox{ and }\qquad\qquad Tx=\sum_{n=1}^{\infty}\langle x_{n}^{\prime},x\rangle_{E^{\prime},E}\cdot y_{n}$ (B.15)
for all $x\in E$, where the convergence in the second series is in $F$. Note that by (B.13) the first condition is equivalent to $(\|x^{\prime}_{n}\otimes y_{n}\|)\in\ell_{1}$, and therefore we have $T=\sum_{n=1}^{\infty}x_{n}^{\prime}\otimes y_{n}$, where the series is absolutely convergent with respect to the operator norm. Finally, the nuclear norm of $T$ is defined by
[…]
It is well-known that the set $\mathfrak{N}(E,F)$ of nuclear operators $T:E\to F$ equipped with $\|\,\cdot\,\|_{\mathrm{nuc}}$ is a Banach space and we have $\|T\|\leq\|T\|_{\mathrm{nuc}}$ for all $T\in\mathfrak{N}(E,F)$, see e.g. [50, Satz VI.5.3] or [12, Theorem 5.25]. Moreover, every finite rank operator is nuclear, and in particular all elementary tensors are nuclear. In addition, for $x^{\prime}\in E^{\prime}$ and $y\in F$ we have
$\|x^{\prime}\|_{E^{\prime}}\cdot\|y\|_{F}=\|x^{\prime}\otimes y\|\leq\|x^{\prime}\otimes y\|_{\mathrm{nuc}}\leq\|x^{\prime}\|_{E^{\prime}}\cdot\|y\|_{F}$
where in the last step we used that $x^{\prime}\otimes y$ is a possible representation (B.15) of $T:=x^{\prime}\otimes y$. Consequently, we have
$\|x^{\prime}\otimes y\|=\|x^{\prime}\otimes y\|_{\mathrm{nuc}}=\|x^{\prime}\|_{E^{\prime}}\cdot\|y\|_{F}$ (B.16)
for all $x^{\prime}\in E^{\prime}$ and $y\in F$. In particular, for $x_{1}^{\prime},x_{2}^{\prime}\in E^{\prime}$ and $y_{1},y_{2}\in F$ with $\|x_{1}^{\prime}\|_{E^{\prime}}\leq\|x_{2}^{\prime}\|_{E^{\prime}}$ and $\|y_{1}\|_{F}\leq\|y_{2}\|_{F}$ we have
$\|x_{1}^{\prime}\otimes y_{1}\|_{\mathrm{nuc}}\leq\|x_{2}^{\prime}\otimes y_{2}\|_{\mathrm{nuc}}\,.$ (B.17)
The next lemma establishes a simple yet useful convergence result for elementary tensors.
Lemma B.5.
Let $E$ and $F$ be Banach spaces, $x^{\prime}\in E^{\prime}$, $y\in F$, and $(x_{n}^{\prime})\subset E^{\prime}$ and $(y_{n})\subset F$ be sequences with $x_{n}^{\prime}\to x^{\prime}$ and $y_{n}\to y$. Then we have
$\|x^{\prime}_{n}\otimes y_{n}-x^{\prime}\otimes y\|_{\mathrm{nuc}}\to 0\,.$
Proof of Lemma B.5: For the nuclear operator $T_{n}:=x^{\prime}_{n}\otimes y_{n}-x^{\prime}\otimes y$ we have the alternative representation $T_{n}=(x^{\prime}_{n}-x^{\prime})\otimes y_{n}-x^{\prime}\otimes(y-y_{n})$ in the sense of (B.15). By the definition of the nuclear norm we hence obtain
$\|x^{\prime}_{n}\otimes y_{n}-x^{\prime}\otimes y\|_{\mathrm{nuc}}\leq\|x^{\prime}_{n}-x^{\prime}\|_{E^{\prime}}\cdot\|y_{n}\|_{F}+\|x^{\prime}\|_{E^{\prime}}\cdot\|y-y_{n}\|_{F}\to 0\,,$
since the convergence of $(y_{n})$ ensures that the sequence $(\|y_{n}\|_{F})$ is bounded. ∎
Given two functions $X^{\prime}:\Omega\to E^{\prime}$ and $Y:\Omega\to F$ we define the $E^{\prime}\otimes F$-valued function $X^{\prime}\otimes Y:\Omega\to E^{\prime}\otimes F$ pointwise by elementary tensors, that is
$(X^{\prime}\otimes Y)(\omega):=(X^{\prime}(\omega))\otimes(Y(\omega))\,,\qquad\qquad\omega\in\Omega.$
Under suitable assumptions, it turns out that the Bochner integral of tensor space valued functions can be used to compute cross covariances. To establish these results we need the following elementary lemma, which investigates such functions in the case of $X$ or $Y$ being constant.
Lemma B.6.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be Banach spaces, and $X^{\prime}\in{{\cal L}_{1}(\mathrm{P},E^{\prime})}$ and $Y\in{{\cal L}_{1}(\mathrm{P},F)}$. Then for all $x^{\prime}\in E^{\prime}$ and $y\in F$ we have $X^{\prime}\otimes y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E,F))}$ and $x^{\prime}\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E,F))}$ with
$\int_{\Omega}X^{\prime}\otimes y\,\mathrm{d}\mathrm{P}=\biggr{(}\int_{\Omega}X^{\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y\qquad\qquad\mbox{ and }\qquad\qquad\int_{\Omega}x^{\prime}\otimes Y\,\mathrm{d}\mathrm{P}=x^{\prime}\otimes\biggr{(}\int_{\Omega}Y\,\mathrm{d}\mathrm{P}\biggl{)}\,.$
Here, the integrals of $X^{\prime}\otimes y$ and $x^{\prime}\otimes Y$ are $\mathfrak{N}(E,F)$-valued Bochner integrals and the integrals of $X^{\prime}$ and $Y$ are $E^{\prime}$-valued, respectively $F$-valued Bochner integrals.
Proof of Lemma B.6: We only prove the assertion for $X^{\prime}\otimes y$, since the proof for $x^{\prime}\otimes Y$ is analogous. Let us first assume that $X^{\prime}$ is a measurable step function. We fix a representation $X=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}x_{i}^{\prime}$, where $A_{i}\in{\cal A}$ and $x_{i}^{\prime}\in E^{\prime}$. For $\omega\in\Omega$ the bilinearity of elementary tensors then gives
$(X^{\prime}\otimes y)(\omega)=(X^{\prime}(\omega))\otimes y=\biggl{(}\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}(\omega)x_{i}^{\prime}\biggr{)}\otimes y=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}(\omega)(x_{i}^{\prime}\otimes y)\,.$
Consequently, $X^{\prime}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is a measurable step function, and therefore we have $X^{\prime}\otimes y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E,F))}$, see also the discussion in front of (A.5). Moreover, the construction of the Bochner integral for step functions (A.3) shows
$\int_{\Omega}X^{\prime}\otimes y\,\mathrm{d}\mathrm{P}=\sum_{i=1}^{m}\mathrm{P}({A_{i}})\cdot(x_{i}^{\prime}\otimes y)=\biggl{(}\sum_{i=1}^{m}\mathrm{P}({A_{i}})\cdot x_{i}^{\prime}\biggr{)}\otimes y=\biggr{(}\int_{\Omega}X^{\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y\,.$
Let us now consider the general case $X^{\prime}\in{{\cal L}_{1}(\mathrm{P},E^{\prime})}$. Since $X^{\prime}$ is strongly measurable, there exists a sequence $(X^{\prime}_{n})$ of measurable step functions approximating $X^{\prime}$ in the sense of (A.1) and (A.2). For each fixed $\omega\in\Omega$, Lemma B.5 in combination with (A.1) then shows
$\|(X_{n}^{\prime}\otimes y)(\omega)-(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}\to 0\,.$
Furthermore, (A.2) and (B.17) gives $\|(X^{\prime}_{n}\otimes y)(\omega)\|_{\mathrm{nuc}}\leq\|(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}$, and we already know that each $X^{\prime}_{n}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is a measurable step function. Consequently, $X^{\prime}\otimes y:\Omega\to\mathfrak{N}(E,F)$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.16) we obtain
$\int_{\Omega}\|(X^{\prime}\otimes y)(\omega)\|_{\mathrm{nuc}}\,\mathrm{d}\mathrm{P}(\omega)=\int_{\Omega}\|X^{\prime}(\omega)\|_{E}\cdot\|y\|_{F}\,\mathrm{d}\mathrm{P}(\omega)<\infty\,.$
In other words we have found $X^{\prime}\otimes y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E,F))}$. Finally, with convergence in $\mathfrak{N}(E,F)$ we have
$\int_{\Omega}X^{\prime}\otimes y\,\mathrm{d}\mathrm{P}=\lim_{n\to\infty}\int_{\Omega}X^{\prime}_{n}\otimes y\,\mathrm{d}\mathrm{P}=\lim_{n\to\infty}\biggr{(}\biggr{(}\int_{\Omega}X^{\prime}_{n}\,\mathrm{d}\mathrm{P}\biggr{)}\otimes y\biggl{)}$ $=\biggr{(}\lim_{n\to\infty}\int_{\Omega}X^{\prime}_{n}\,\mathrm{d}\mathrm{P}\biggr{)}\otimes y$
$=\biggr{(}\int_{\Omega}X^{\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y\,,$
where the first and last step used the construction of Bochner integrals (A.4), the second step used the already verified assertion in the case of step functions, and the third step used Lemma B.5. ∎
Let us now turn to the announced integral representation of cross covariance operators. To this end, we fix two Banach spaces $E$ and $F$, as well as $x\in E$ and $y\in F$. For $x^{\prime\prime}:=\iota_{E}x\in E^{\prime\prime}$, where $\iota_{E}:E\to E^{\prime\prime}$ denotes the canonical, isometric embedding of $E$ into its bidual $E$, we then write
$x\otimes y:=x^{\prime\prime}\otimes y=(\iota_{E}x)\otimes y\,.$
Obviously, $x\otimes y\in E^{\prime\prime}\otimes F$ is the bounded linear map $E^{\prime}\to F$ given by
$(x\otimes y)(x^{\prime})=\langle x^{\prime},x\rangle_{E^{\prime},E}\cdot y\,,\qquad\qquad x^{\prime}\in E^{\prime}.$
By construction the bilinearity is preserved, and since $\|x\|_{E}=\|\iota_{E}x\|_{E^{\prime\prime}}=\|x^{\prime\prime}\|_{E^{\prime\prime}}$ holds true, we quickly verify that (B.16) and (B.17) read as
$\|x\otimes y\|$ $=\|x\otimes y\|_{\mathrm{nuc}}=\|x\|_{E}\cdot\|y\|_{F}\,,$
$\|x_{1}\otimes y_{1}\|_{\mathrm{nuc}}$ $\leq\|x_{2}\otimes y_{2}\|_{\mathrm{nuc}}$
for all $x,x_{1},x_{2}\in E$ and $y,y_{1},y_{2}\in F$ with $\|x_{1}\|_{E}\leq\|x_{2}\|_{E}$ and $\|y_{1}\|_{F}\leq\|y_{2}\|_{F}$. Moreover, Lemma B.5 also holds, since $x_{n}\to x$ in $E$ implies $\iota_{E}x_{n}\to\iota_{E}x$ in $E^{\prime\prime}$.
Finally, Lemma B.6 remains true: Indeed, if we have for example an $X\in{{\cal L}_{1}(\mathrm{P},E)}$, then $X^{\prime\prime}:=\iota_{E}\circ X\in{{\cal L}_{1}(\mathrm{P},E^{\prime\prime})}$ as discussed around (A.10), and therefore Lemma B.6 gives $X\otimes y=X^{\prime\prime}\otimes y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$ with
$\int_{\Omega}X\otimes y\,\mathrm{d}\mathrm{P}=\int_{\Omega}X^{\prime\prime}\otimes y\,\mathrm{d}\mathrm{P}=\biggr{(}\int_{\Omega}X^{\prime\prime}\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y=\biggr{(}\int_{\Omega}\iota_{E}\circ X\,\mathrm{d}\mathrm{P}\biggl{)}\otimes y$ $=\biggl{(}\iota_{E}\biggr{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggl{)}\biggr{)}\otimes y$
$=\biggl{(}\int_{\Omega}X\,\mathrm{d}\mathrm{P}\biggr{)}\otimes y$
by an application of (A.10) and our nomenclature for tensors of the form $x\otimes y$. In particular, the Bochner integral of $X$ is an $E$-valued Bochner integral and not just an $E^{\prime\prime}$-valued integral as a naïve application of Lemma B.6 would suggest.
With these preparation we can now present an alternative way to compute the cross covariance operator.
Lemma B.7.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $E$ and $F$ be Banach spaces, and $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},F)}$. Then the map $X\otimes Y:\Omega\to\mathfrak{N}(E^{\prime},F)$ is strongly measurable with $X\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$. Moreover, $\operatorname{cov}(Y,X)$ can be computed by the $\mathfrak{N}(E^{\prime},F)$-valued Bochner integral
$\operatorname{cov}(Y,X)=\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}-(\mathbb{E}X)\otimes(\mathbb{E}Y)\,.$ (B.18)
Proof of Lemma B.7: Our first goal is to show that it suffices to consider centered $X$ and $Y$. To this end, we pick some $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},F)}$ and define $X_{0}:=X-\mathbb{E}X$ and $Y_{0}:=Y-\mathbb{E}Y$. Furthermore, we assume that the assertion is true for $X_{0}$ and $Y_{0}$. Our goal is therefore to show that the assertion is also true for $X$ and $Y$. To this end, we first observe that the bilinearity of tensors gives
$X_{0}\otimes Y_{0}=(X-\mathbb{E}X)\otimes(Y-\mathbb{E}Y)=X\otimes Y-X\otimes(\mathbb{E}Y)-(\mathbb{E}X)\otimes Y+(\mathbb{E}X)\otimes(\mathbb{E}Y)\,.$
Now, our assumption guarantees $X_{0}\otimes Y_{0}\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$, and Lemma B.6 ensures both $X\otimes(\mathbb{E}Y)\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$ and $(\mathbb{E}X)\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$. Finally, $(\mathbb{E}X)\otimes(\mathbb{E}Y)\in\mathfrak{N}(E^{\prime},F)$ does not depend on $\omega$. In summary, we thus have $X\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$. Moreover, our calculation above yields
$\operatorname{cov}(Y,X)=\operatorname{cov}(Y_{0},X_{0})$ $=\int_{\Omega}X_{0}\otimes Y_{0}\,\mathrm{d}\mathrm{P}$
$=\int_{\Omega}X\otimes Y-X\otimes(\mathbb{E}Y)-(\mathbb{E}X)\otimes Y+(\mathbb{E}X)\otimes(\mathbb{E}Y)\,\mathrm{d}\mathrm{P}$
$=\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}-\int_{\Omega}X\otimes(\mathbb{E}Y)\,\mathrm{d}\mathrm{P}-\int_{\Omega}(\mathbb{E}X)\otimes Y\,\mathrm{d}\mathrm{P}+(\mathbb{E}X)\otimes(\mathbb{E}Y)$
$=\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}-(\mathbb{E}X)\otimes(\mathbb{E}Y)-(\mathbb{E}X)\otimes(\mathbb{E}Y)+(\mathbb{E}X)\otimes(\mathbb{E}Y)$
$=\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}-(\mathbb{E}X)\otimes(\mathbb{E}Y)\,,$
where in the second to last step we used Lemma B.6 twice.
Let us now turn to the actual proof. Here, we begin by considering the case, in which $X$ and $Y$ are measurable step functions. As discussed above, we may assume without loss of generality that $\mathbb{E}X=0$ and $\mathbb{E}Y=0$. We now fix some representations $X=\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}x_{i}$ and $Y=\sum_{j=1}^{n}\boldsymbol{1}_{B_{j}}y_{j}$, where $A_{i},B_{j}\in{\cal A}$, $x_{i}\in E$, and $y_{j}\in F$. For $\omega\in\Omega$, the bilinearity of elementary tensors then gives
$(X\otimes Y)(\omega)=\biggl{(}\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}(\omega)x_{i}\biggr{)}\otimes\biggl{(}\sum_{j=1}^{n}\boldsymbol{1}_{B_{j}}(\omega)y_{j}\biggr{)}=\sum_{i=1}^{m}\sum_{j=1}^{n}\boldsymbol{1}_{A_{i}\cap B_{j}}(\omega)(x_{i}\otimes y_{j})\,.$
Consequently, $X\otimes Y:\Omega\to\mathfrak{N}(E^{\prime},F)$ is a measurable step function, and therefore we have $X\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$, see also the discussion in front of (A.5). Finally, for $x^{\prime}\in E^{\prime}$ the construction of the Bochner integral for step functions (A.3) shows
$\biggl{(}\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}\biggr{)}(x^{\prime})=\biggl{(}\sum_{i=1}^{m}\sum_{j=1}^{n}\mathrm{P}({A_{i}\cap B_{j}})\cdot(x_{i}\otimes y_{j})\biggr{)}(x^{\prime})$ $=\sum_{i=1}^{m}\sum_{j=1}^{n}\mathrm{P}({A_{i}\cap B_{j}})\cdot(x_{i}\otimes y_{j})(x^{\prime})$
$=\sum_{i=1}^{m}\sum_{j=1}^{n}\mathrm{P}({A_{i}\cap B_{j}})\cdot\langle x^{\prime},x_{i}\rangle_{E^{\prime},E}\cdot y_{j}\,,$
as well as
$\operatorname{cov}(Y,X)(x^{\prime})=\int_{\Omega}\langle x^{\prime},X\rangle_{E^{\prime},E}\,Y\,\mathrm{d}\mathrm{P}$ $=\int_{\Omega}\biggl{(}\sum_{i=1}^{m}\boldsymbol{1}_{A_{i}}\langle x^{\prime},x_{i}\rangle_{E^{\prime},E}\biggr{)}\,\biggr{(}\sum_{j=1}^{n}\boldsymbol{1}_{B_{j}}y_{j}\biggl{)}\,\mathrm{d}\mathrm{P}$
$=\sum_{i=1}^{m}\sum_{j=1}^{n}\int_{\Omega}\boldsymbol{1}_{A_{i}\cap B_{j}}\langle x^{\prime},x_{i}\rangle_{E^{\prime},E}\cdot y_{j}\,\mathrm{d}\mathrm{P}$
$=\sum_{i=1}^{m}\sum_{j=1}^{n}\mathrm{P}({A_{i}\cap B_{j}})\cdot\langle x^{\prime},x_{i}\rangle_{E^{\prime},E}\cdot y_{j}\,.$
Combining both calculations yields (B.18).
Let us finally consider the case of centered $X\in{{\cal L}_{2}(\mathrm{P},E)}$ and $Y\in{{\cal L}_{2}(\mathrm{P},F)}$. Since $X$ and $Y$ are strongly measurable, there then exist sequences $(X_{n})$ and $(Y_{n})$ of measurable step functions approximating $X$, respectively $Y$ in the sense of (A.1) and (A.2). For each fixed $\omega\in\Omega$, Lemma B.5 in combination with (A.1) for $X_{n}$ and $Y_{n}$ then shows
$\|(X_{n}\otimes Y_{n})(\omega)-(X\otimes Y)(\omega)\|_{\mathrm{nuc}}\to 0\,.$
Furthermore, (A.2) and (B.17) gives $\|(X_{n}\otimes Y_{n})(\omega)\|_{\mathrm{nuc}}\leq\|(X\otimes Y)(\omega)\|_{\mathrm{nuc}}$. In addition, we already know that each $X_{n}\otimes Y_{n}:\Omega\to\mathfrak{N}(E^{\prime},F)$ is a measurable step function. Consequently, $X\otimes Y$ is strongly measurable as discussed around (A.1) and (A.2). Moreover, with the help of (B.16) we obtain
$\int_{\Omega}\|(X\otimes Y)(\omega)\|_{\mathrm{nuc}}\,\mathrm{d}\mathrm{P}(\omega)=\int_{\Omega}\|X(\omega)\|_{E}\cdot\|Y(\omega)\|_{F}\,\mathrm{d}\mathrm{P}(\omega)<\infty\,,$
where in the last step we used the Cauchy-Schwarz inequality. In other words we have found $X\otimes Y\in{{\cal L}_{1}(\mathrm{P},\mathfrak{N}(E^{\prime},F))}$.
To establish (B.18), we first note that (A.1) and (A.2) for $X$ and $(X_{n})$ give $\mathbb{E}X_{n}\to\mathbb{E}X=0$ by the definition of Bochner integrals and our assumption that $X$ is centered. Analogously, we find $\mathbb{E}Y_{n}\to 0$. Combining both gives $\mathbb{E}X_{n}\otimes\mathbb{E}Y_{n}\to 0$ by Lemma B.5. For $x^{\prime}\in E^{\prime}$ we thus find
$\biggl{(}\int_{\Omega}X\otimes Y\,\mathrm{d}\mathrm{P}\biggr{)}(x^{\prime})$ $=\biggl{(}\lim_{n\to\infty}\int_{\Omega}X_{n}\otimes Y_{n}\,\mathrm{d}\mathrm{P}\biggr{)}(x^{\prime})$
$=\lim_{n\to\infty}\Biggl{(}\biggl{(}\int_{\Omega}X_{n}\otimes Y_{n}\,\mathrm{d}\mathrm{P}\biggr{)}(x^{\prime})\Biggr{)}$
$=\lim_{n\to\infty}\Biggl{(}\biggl{(}\int_{\Omega}X_{n}\otimes Y_{n}\,\mathrm{d}\mathrm{P}\biggr{)}(x^{\prime})-(\mathbb{E}X_{n}\otimes\mathbb{E}Y_{n})(x^{\prime})\Biggr{)}$
$=\lim_{n\to\infty}\Bigl{(}\operatorname{cov}(Y_{n},X_{n})(x^{\prime})\Bigr{)}\,,$ (B.19)
where in the first step we used the definition of Bochner integrals, the second step used that the limit in the second expression is with respect to the nuclear norm and thus also with respect to the operator norm, see also the explanation following (A.4), and in the last step we used the already established identity (B.18) for step functions. Moreover, we also have
$\lim_{n\to\infty}\Bigl{(}\operatorname{cov}(Y_{n},X_{n})(x^{\prime})\Bigr{)}=\lim_{n\to\infty}\int_{\Omega}\langle x^{\prime},X_{n}\rangle_{E^{\prime},E}\cdot Y_{n}\,\mathrm{d}\mathrm{P}$ $=\int_{\Omega}\langle x^{\prime},X\rangle_{E^{\prime},E}\,Y\,\mathrm{d}\mathrm{P}$
$=\operatorname{cov}(Y,X)(x^{\prime})\,,$ (B.20)
where the second step is based upon the dominated convergence property discussed around (A.7) to (A.9) with
$\langle x^{\prime},X_{n}(\omega)\rangle_{E^{\prime},E}\cdot Y_{n}(\omega)\to\langle x^{\prime},X(\omega)\rangle_{E^{\prime},E}\cdot Y(\omega)$
for all $\omega\in\Omega$ and the $\mathrm{P}$-integrable $g:\Omega\to[0,\infty]$ defined by $g(\omega):=\|x^{\prime}\|_{E^{\prime}}\cdot\|X(\omega)\|_{E}\cdot\|Y(\omega)\|_{F}$. Combining (B.19) and (B.20) yields (B.18). ∎
C Appendix: Characteristic Functions
Let $\mathrm{P}$ be a probability measure on $\mathbb{R}^{d}$. Then the characteristic function $\varphi_{\mathrm{P}}:\mathbb{R}^{d}\to\mathbb{C}$ of $\mathrm{P}$ is defined by
$\varphi_{\mathrm{P}}(t):=\int_{\mathbb{R}^{d}}\exp(\mathrm{i}\langle t,x\rangle)\,\mathrm{d}\mathrm{P}(x)\,,\qquad\qquad t\in\mathbb{R}^{d},$
[…]
It is well known, see e.g. [27, Theorem 15.21], that every characteristic function $\varphi_{X}$ is uniformly continuous and we have $|\varphi_{X}(t)|\leq 1$ for all $t\in\mathbb{R}$ with $\varphi_{X}(0)=1$. Moreover, the characteristic function can be used to identify measures. The following well-known theorem, which can be found in e.g. [10, Proposition 10.3.11] or [27, Theorem 15.8] formalizes this.
Theorem C.1.
Let $\mathrm{P}$ and $\mathrm{Q}$ be probability measures on $\mathbb{R}^{d}$. Then we have $\mathrm{P}=\mathrm{Q}$ if and only if $\varphi_{\mathrm{P}}=\varphi_{\mathrm{Q}}$.
In addition, characteristic functions of affine linear transformations of random variables can be easily computed, as the following simple and well-known lemma shows.
Lemma C.2.
Let $(\Omega,{\cal A},\mathrm{P})$ be a probability space, $X:\Omega\to\mathbb{R}^{d}$ be measurable, $A$ be an $m\times d$-matrix, $a\in\mathbb{R}^{d}$, and $b\in\mathbb{R}^{m}$. Then, for all $s\in\mathbb{R}$ and $t\in\mathbb{R}^{m}$, the following two identities hold true
$\varphi_{\langle a,X\rangle}(s)$ $=\varphi_{X}(sa)\,,$
$\varphi_{AX+b}(t)$ $=\exp(\mathrm{i}\langle t,b\rangle)\cdot\varphi_{X}(A^{\mathsf{T}}t)\,.$
Proof of Lemma C.2: The second identity can be found in e.g. [22, Theorem 13.3], and the first identity is a special case of the second since $\langle a,X\rangle=a^{\mathsf{T}}X$ and $as=sa$, where $as$ denotes the matrix product and $a$ are $s$ are viewed as $d\times 1$ and $1\times 1$-matrices, respectively. ∎
The following result, whose implication iii) $\Rightarrow$ i) is known as Lévy’s continuity theorem relates convergence in distribution to pointwise convergence of the characteristic functions.
Theorem C.3.
Let $(X_{n})$ be a sequence of $\mathbb{R}^{d}$-valued random variables and $X$ be another $\mathbb{R}^{d}$-valued random variable. Then the following statements are equivalent:
  1. i)
    We have $X_{n}\to X$ in distribution.
  2. ii)
    For all $a\in\mathbb{R}^{d}$ with $\|a\|_{2}=1$ we have $\langle a,X_{n}\rangle\to\langle a,X\rangle$ in distribution.
  3. iii)
    We have $\varphi_{X_{n}}(t)\to\varphi_{X}(t)$ for all $t\in\mathbb{R}^{d}$.
Proof of Theorem C.3: i) $\Rightarrow$ iii). Follows immediately from the definition of convergence in distribution and from the form of characteristic functions.
iii) $\Rightarrow$ i). See, for example, [14, Theorem 9.8.2] or [27, Theorem 15.23].
ii) $\Rightarrow$ iii). Let us fix a $t\in\mathbb{R}^{d}$ with $t\neq 0$. We write $a:=\|t\|_{2}^{-1}t$ and $s:=\|t\|_{2}$ as well as $Y_{n}:=\langle a,X_{n}\rangle$ and $Y:=\langle a,X\rangle$. By Lemma C.2 we then obtain
$\varphi_{X_{n}}(t)=\varphi_{X_{n}}(sa)=\varphi_{\langle a,X_{n}\rangle}(s)=\varphi_{Y_{n}}(s)$
and analogously we find $\varphi_{X}(t)=\varphi_{Y}(s)$. Moreover, we have $Y_{n}\to Y$ in distribution and hence the already established implication i) $\Rightarrow$ iii) gives $\varphi_{Y_{n}}(s)\to\varphi_{Y}(s)$. The previous two identities then yield $\varphi_{X_{n}}(t)\to\varphi_{X}(t)$ for all $t\neq 0$. Moreover, we always have $\varphi_{X_{n}}(0)=1=\varphi_{X}(0)$, and hence iii) has been shown.
i) $\Rightarrow$ ii). Let us fix a bounded and continuous $f:\mathbb{R}\to\mathbb{R}$. Then $f(\langle a,\,\cdot\,\rangle):\mathbb{R}^{d}\to\mathbb{R}$ is also bounded and continuous, and hence the transformation formula for image measures together with i) yields
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{\langle a,X_{n}\rangle}=\int_{\mathbb{R}^{d}}f(\langle a,\,\cdot\,\rangle)\,\mathrm{d}\mathrm{P}_{X_{n}}\to\int_{\mathbb{R}^{d}}f(\langle a,\,\cdot\,\rangle)\,\mathrm{d}\mathrm{P}_{X}=\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{\langle a,X\rangle}\,,$
that is, we have found ii). ∎
The next result applies Theorem C.3 to Gaussian random variables.
Theorem C.4.
Let $(X_{n})$ be a sequence of $\mathbb{R}$-valued Gaussian random variables and $X$ be an $\mathbb{R}$-valued random variable with $X_{n}\to X$ in distribution. Then $X$ is a Gaussian random variable and we have:
$\mathbb{E}X_{n}$ $\to\mathbb{E}X\,,$
$\operatorname{var}X_{n}$ $\to\operatorname{var}X\,.$
Proof of Theorem C.4: Let us write $\mu_{n}:=\mathbb{E}X_{n}$ and $\sigma_{n}:=\sqrt{\operatorname{var}X_{n}}$. By Theorem C.3 we then know that
$\mathrm{e}^{\mathrm{i}t\mu_{n}}\cdot\exp\Bigl{(}-\frac{\sigma_{n}^{2}t^{2}}{2}\Bigr{)}=\varphi_{X_{n}}(t)\to\varphi_{X}(t)\,,\qquad\qquad t\in\mathbb{R}.$ (C.1)
Taking the absolute value in $\mathbb{C}$ yields
$\exp\Bigl{(}-\frac{\sigma_{n}^{2}t^{2}}{2}\Bigr{)}=|\varphi_{X_{n}}(t)|\to|\varphi_{X}(t)|=:a_{t}\,,\qquad\qquad t\in\mathbb{R}.$ (C.2)
Since $\varphi_{X}$ is uniformly continuous with $\varphi_{X}(0)=1$, there exists a $t_{0}>0$ such that $a_{t}>0$ for all $t\in[-t_{0},t_{0}]$. Consequently, we have
$\sigma^{2}_{n}\to\sigma^{2}:=-\frac{2}{t_{0}^{2}}\,\ln(a_{t_{0}})\,,$ (C.3)
and in particular, we find $b:=1+\sup_{n}\sigma_{n}<\infty$.
Our next step is to show that $(\mu_{n})$ is a bounded sequence. To this end, let us first assume that we had a subsequence $\mu_{n_{k}}\to\infty$. Since $\mathrm{P}_{X}$ is a probability measure there exists an $a>0$ with $\mathrm{P}_{X}([-a,a])\geq 1/2$. Clearly, we may assume without loss of generality that $\mu_{n_{k}}>a+1$ for all $k\geq 1$. Let us fix a continuous function $f:\mathbb{R}\to[0,1]$ with $f(t)=1$ for all $t\in[-a,a]$ and $f(t)=0$ for all $t\not\in[-a-1,a+1]$. By the assumed convergence $X_{n}\to X$ in distribution, we then obtain
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{X_{n_{k}}}\to\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{X}\geq\int_{[-a,a]}f\,\mathrm{d}\mathrm{P}_{X}\geq 1/2\,.$ (C.4)
To find a contradiction, let us first assume that there is an $n_{k}$ with $\sigma_{n_{k}}=0$. This gives $\mathrm{P}_{X_{n_{k}}}=\delta_{\{\mu_{n_{k}}\}}$, and since $\mu_{n_{k}}>a+1$ our construction of $f$ would then yield
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{X_{n_{k}}}=f(\mu_{n_{k}})=0\,.$
In view of (C.4) we thus conclude that there exists a $k_{0}\geq 1$ with $\sigma_{n_{k}}>0$ for all $k\geq k_{0}$. For such $k$, however, we find
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{X_{n_{k}}}$ $=\frac{1}{\sqrt{2\pi\sigma_{n_{k}}^{2}}}\int_{\mathbb{R}}f(t)\exp\Bigl{(}-\frac{(t-\mu_{n_{k}})^{2}}{2\sigma_{n_{k}}^{2}}\Bigr{)}\,\mathrm{d}t$
$=\frac{1}{\sqrt{2\pi\sigma_{n_{k}}^{2}}}\int_{\mathbb{R}}f(t+\mu_{n_{k}})\exp\Bigl{(}-\frac{t^{2}}{2\sigma_{n_{k}}^{2}}\Bigr{)}\,\mathrm{d}t$
$=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}f(\sigma_{n_{k}}t+\mu_{n_{k}})\exp\Bigl{(}-\frac{t^{2}}{2}\Bigr{)}\,\mathrm{d}t$
$=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\sigma_{n_{k}}^{-1}(a+1-\mu_{n_{k}})}f(\sigma_{n_{k}}t+\mu_{n_{k}})\exp\Bigl{(}-\frac{t^{2}}{2}\Bigr{)}\,\mathrm{d}t$
$\leq{\cal N}(0,1)\bigl{(}(-\infty,\sigma_{n_{k}}^{-1}(a+1-\mu_{n_{k}})]\bigr{)}\,,$
where in the second to last step we used that $\sigma_{n_{k}}t+\mu_{n_{k}}\leq a+1$ is equivalent to $t\leq\sigma^{-1}_{n_{k}}(a+1-\mu_{n_{k}})$. Now observe that the definition of $b$ implies $\sigma_{n_{k}}\leq b$ and thus $\sigma_{n_{k}}^{-1}\geq b^{-1}>0$ for all $k\geq k_{0}$. Since we also have $a+1-\mu_{n_{k}}<0$ for all $k\geq 1$, we thus conclude
$\sigma_{n_{k}}^{-1}(a+1-\mu_{n_{k}})\leq b^{-1}(a+1-\mu_{n_{k}})$
for all $k\geq k_{0}$. Consequently, the estimation above can be continued to
$\int_{\mathbb{R}}f\,\mathrm{d}\mathrm{P}_{X_{n_{k}}}\leq{\cal N}(0,1)\bigl{(}(-\infty,\sigma_{n_{k}}^{-1}(a+1-\mu_{n_{k}})]\bigr{)}\leq{\cal N}(0,1)\bigl{(}(-\infty,b^{-1}(a+1-\mu_{n_{k}})]\bigr{)}\to 0$
for $k\to\infty$, where in the last step we used $\mu_{n_{k}}\to\infty$. Clearly, this contradicts (C.4) and hence there is no subsequence $\mu_{n_{k}}\to\infty$. Analogously, we find that there is no subsequence $\mu_{n_{k}}\to-\infty$. In other words, $(\mu_{n})$ is a bounded sequence.
Now, since $(\mu_{n})$ is a bounded sequence, there exist a $\mu\in\mathbb{R}$ and a subsequence $(\mu_{n_{k}})$ with $\mu_{n_{k}}\to\mu$, and this in turn yields
$\varphi_{X_{n_{k}}}(t)=\mathrm{e}^{\mathrm{i}t\mu_{n_{k}}}\cdot\exp\Bigl{(}-\frac{\sigma_{n_{k}}^{2}t^{2}}{2}\Bigr{)}\to\mathrm{e}^{\mathrm{i}t\mu}\cdot\exp\Bigl{(}-\frac{\sigma^{2}t^{2}}{2}\Bigr{)}$ (C.5)
for all $t\in\mathbb{R}$. Since the limit function is the characteristic function of ${\cal N}(\mu,\sigma^{2})$ we see by (C.1) and the uniqueness of the characteristic function, see Theorem C.1, that $\mathrm{P}_{X}={\cal N}(\mu,\sigma^{2})$, that is, $X$ is indeed a Gaussian random variable.
In addition, $\operatorname{var}X_{n}\to\operatorname{var}X$ immediately follows from (C.3). Finally, we already know that $\mathbb{E}X_{n_{k}}=\mu_{n_{k}}\to\mu=\mathbb{E}X$. If, however, the full sequence $(\mathbb{E}X_{n})$ did not converge to $\mathbb{E}X$, then there would be another subsequence $(\mu_{m_{k}})$ and an $\varepsilon>0$ with $|\mu_{m_{k}}-\mu|\geq\varepsilon$ for all $k\geq 1$. Since $(\mu_{n})$ is bounded, so is $(\mu_{m_{k}})$, and hence there would exist a subsequence of $(\mu_{m_{k}})$ converging to some $\tilde{\mu}\in\mathbb{R}$. Our construction ensures $\mu\neq\tilde{\mu}$ and repeating the arguments around (C.5) would give us $\mathrm{P}_{X}={\cal N}(\tilde{\mu},\sigma^{2})$. This, however, is impossible by the uniqueness of the weak limit of probability measures, see discussion below (6). ∎
Finally, we briefly discuss characteristic functions in the Banach space case. To this end, let $E$ be a separable Banach space and $\mathrm{P}$ be a probability measure on $\mathfrak{B}(E)=\sigma(E^{\prime})$. Then the characteristic function $\varphi_{\mathrm{P}}:E^{\prime}\to\mathbb{C}$ of $\mathrm{P}$ is defined by
$\varphi_{\mathrm{P}}(x^{\prime}):=\int_{E}\mathrm{e}^{\mathrm{i}\langle x^{\prime},x\rangle_{E^{\prime},E}}\,\mathrm{d}\mathrm{P}(x)\,,\qquad\qquad x^{\prime}\in E^{\prime},$
see e.g. [46, Chapter IV.2.1]. Like in the finite dimensional case, the characteristic function can be used to identify probability measures. Namely, if $\mathrm{P}$ and $\mathrm{Q}$ are probability measures with $\varphi_{\mathrm{P}}=\varphi_{\mathrm{Q}}$, then we have $\mathrm{P}=\mathrm{Q}$, see e.g. [46, Theorem 2.2 in Chapter IV.2.1].
Similar to the finite dimensional case, a Gaussian measure $\mathrm{P}$ on a separable Banach space $E$ has a characteristic function that is solely described by its mean and covariance. The following result, which can be found in e.g. [46, Proposition 2.8 in Chapter IV.2.4], provides the details.
Proposition C.5.
Let $E$ be a separable Banach space and $\mathrm{P}$ be a probability measure on $\mathfrak{B}(E)$. Then the following statements are equivalent:
  1. i)
    $\mathrm{P}$ is a Gaussian measure.
  2. ii)
    There exist a $\mu\in E$ and an abstract covariance operator $K:E^{\prime}\to E$ such that
    $\varphi_{\mathrm{P}}(x^{\prime})=\mathrm{e}^{\mathrm{i}\langle x^{\prime},\mu\rangle_{E^{\prime},E}}\cdot\exp\biggl{(}-\frac{\langle x^{\prime},Kx^{\prime}\rangle_{E^{\prime},E}}{2}\biggr{)}\,,\qquad\qquad x^{\prime}\in E^{\prime}\,.$ (C.6)
In this case, $\mu$ is the mean of $\mathrm{P}$ and $K$ is the covariance operator of $\mathrm{P}$.
Finally, the following generalization of Theorem C.3 holds true.
Theorem C.6.
Let $E$ be a separable Banach space and $(\mathrm{P}_{n})$ be a sequence of probability measures on $\mathfrak{B}(E)$ such that the following assumptions are satisfied:
  1. i)
    The set $\{\mathrm{P}_{n}:n\geq 1\}$ is uniformly tight.
  2. ii)
    There exists a function $\varphi:E^{\prime}\to\mathbb{C}$ such that $\varphi_{\mathrm{P}_{n}}(x^{\prime})\to\varphi(x^{\prime})$ for all $x^{\prime}\in E^{\prime}$.
Then there exists a probability measure $\mathrm{P}$ on $E$ such $\mathrm{P}_{n}\to\mathrm{P}$ weakly and $\varphi_{\mathrm{P}}=\varphi$.
Proof of Theorem C.6: We will apply [46, Theorem 3.1 in Chapter IV.3.1] for $\Gamma:=E^{\prime}$. Here we first note that Hahn-Banach’s theorem ensures that $E^{\prime}$ is a separating set consisting of continuous functions $E\to\mathbb{R}$. Moreover, every probability measure on $\mathfrak{B}(E)$ is a Radon probability measure, see the references around (44). Finally, [6, Theorem 3.8.4] in combination with i) ensures that $\{\mathrm{P}_{n}:n\geq 1\}$ is relatively weakly compact since the underlying space $E$ is a metric space, and thus a completely regular topological space, where the latter notion can be found in e.g. [25, Chapter 4 on p. 117], see also [6, Appendix A.1 on page 363]. Now the assertion follows from [46, Theorem 3.1 in Chapter IV.3.1]. ∎
D Appendix: Regular Conditional Probabilities
In this appendix we collect some useful properties of regular conditional probabilities and their relationship to Banach space valued conditional expectations.
We begin with a general statement that ensures the existence and uniqueness of regular conditional probabilities. For its formulation recall that that a topological space $(T,{\cal T})$ is called Polish, if there exists a metric $\operatorname{d}:T\times T\to[0,\infty)$ generating the topology ${\cal T}$, such that $(T,\operatorname{d})$ is complete and separable. Obviously, separable Banach spaces are Polish spaces. The following theorem, which can be found in e.g. [3, Theorem 44.3], [14, Theorem 10.2.2], and [27, Theorem 8.37], shows that the existence and uniqueness of regular conditional probabilities can be guaranteed as long as the space $T$ in Definition 2.5 is a Polish space and ${\cal B}$ is the corresponding Borel $\sigma$-algebra.
Theorem D.1.
Let $(T,{\cal T})$ be a Polish space with Borel $\sigma$-algebra $\mathfrak{B}$, $(U,{\cal A})$ be a measurable space, and $\mathrm{P}$ be a probability measure on $(T\times U,\mathfrak{B}\otimes{\cal A})$. Then the following statements hold true:
  1. i)
    There exists a regular conditional probability $\mathrm{P}(\,\cdot\,|\,\cdot\,):\mathfrak{B}\times U\to[0,1]$ of $\mathrm{P}$ given $U$.
  2. ii)
    If $\mathrm{P}_{1}(\,\cdot\,|\,\cdot\,)$ and $\mathrm{P}_{2}(\,\cdot\,|\,\cdot\,)$ are regular conditional probabilities of $\mathrm{P}$ given $U$, then there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ such that for all $B\in\mathfrak{B}$ and $u\in U\setminus N$ we have
    $\mathrm{P}_{1}(B|u)=\mathrm{P}_{2}(B|u)\,.$
Our first result describes how regular conditional probabilities of image measures can be computed with the help of regular conditional probabilities of the initial measure.
Theorem D.2.
For $i=0,1$, let $(T_{i},{\cal T}_{i})$ be Polish spaces with Borel $\sigma$-algebras $\mathfrak{B}_{i}$, $(U_{i},{\cal A}_{i})$ be measurable spaces, and $\mathrm{P}^{(i)}$ be probability measures on $(T_{i}\times U_{i},\mathfrak{B}_{i}\otimes{\cal A}_{i})$. Moreover, let $\Phi:T_{0}\to T_{1}$ and $\Psi:U_{0}\to U_{1}$ be measurable maps, such that we have
$\sigma(\Psi)={\cal A}_{0}\qquad\qquad\mbox{ and }\qquad\qquad\mathrm{P}^{(1)}=\mathrm{P}^{(0)}_{(\Phi,\Psi)}\,.$
Then the following statements hold true:
  1. i)
    We have $\mathrm{P}^{(1)}_{U_{1}}=(\mathrm{P}^{(0)}_{U_{0}})_{\Psi}$.
  2. ii)
    For all regular conditional probabilities $\mathrm{P}^{(0)}(\,\cdot\,|\,\cdot\,)$ of $\mathrm{P}^{(0)}$ given $U_{0}$ and $\mathrm{P}^{(1)}(\,\cdot\,|\,\cdot\,)$ of $\mathrm{P}^{(1)}$ given $U_{1}$ there exists an $N\in{\cal A}_{0}$ with $\mathrm{P}^{(0)}_{U_{0}}(N)=0$ such that for all $B_{1}\in\mathfrak{B}_{1}$ and $u_{0}\in U_{0}\setminus N$ we have
    $\mathrm{P}^{(0)}\bigl{(}(\Phi^{-1}(B_{1})|u_{0}\bigr{)}=\mathrm{P}^{(1)}\bigl{(}B_{1}|\Psi(u_{0})\bigr{)}\,.$ (D.1)
Proof of Theorem D.2: We begin by making some initial observations. To this end, we write $\Upsilon:=(\Phi,\Psi):T_{0}\times U_{0}\to T_{1}\times U_{1}$. A simple calculation then shows
$\pi_{T_{1}}\circ\Upsilon=\Phi\circ\pi_{T_{0}}\qquad\qquad\mbox{ and }\qquad\qquad\pi_{U_{1}}\circ\Upsilon=\Psi\circ\pi_{U_{0}}\,,$
where $\pi_{T_{i}}$ and $\pi_{U_{i}}$ are the coordinate projections for $i=0,1$. Since $\Phi\circ\pi_{T_{0}}$ and $\Psi\circ\pi_{U_{0}}$ are measurable, we then see that $\Upsilon$ is measurable with respect to the product $\sigma$-algebras, see e.g. [27, Corollary 1.82], and in particular our assumption $\mathrm{P}^{(1)}=\mathrm{P}^{(0)}_{(\Phi,\Psi)}=\mathrm{P}^{(0)}_{\Upsilon}$ makes sense. For $B_{1}\in\mathfrak{B}_{1}$ and $A_{1}\in{\cal A}_{1}$ this in turn implies
$\mathrm{P}^{(1)}(B_{1}\times A_{1})=\mathrm{P}^{(0)}(\Upsilon^{-1}(B_{1}\times A_{1}))$ $=\mathrm{P}^{(0)}\bigl{(}\Phi^{-1}(B_{1})\times\Psi^{-1}(A_{1})\bigr{)}\,.$ (D.2)
i). Our initial observations together with the standard formula for image measures of composed transformations imply
$\mathrm{P}^{(1)}_{U_{1}}=\mathrm{P}^{(1)}_{\pi_{U_{1}}}=(\mathrm{P}^{(0)}_{\Upsilon})_{\pi_{U_{1}}}=\mathrm{P}^{(0)}_{\pi_{U_{1}}\circ\Upsilon}=\mathrm{P}^{(0)}_{\Psi\circ\pi_{U_{0}}}=(\mathrm{P}^{(0)}_{\pi_{U_{0}}})_{\Psi}=(\mathrm{P}^{(0)}_{U_{0}})_{\Psi}\,.$
ii). For $A_{0}\in{\cal A}_{0}$ and $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$ and measurable $g:U_{1}\to[0,\infty)$ an application of the transformation formula for integrals of image measures, see e.g. [10, Proposition 2.6.8] yields
$\int_{U_{0}}\!\boldsymbol{1}_{A_{0}}\cdot(g\circ\Psi)\,\mathrm{d}\mathrm{P}^{(0)}_{U_{0}}=\int_{U_{0}}\!(\boldsymbol{1}_{A_{1}}\circ\Psi)\cdot(g\circ\Psi)\,\mathrm{d}(\mathrm{P}^{(0)}_{U_{0}})$ $=\int_{U_{1}}\!\boldsymbol{1}_{A_{1}}\cdot g\,\mathrm{d}(\mathrm{P}^{(0)}_{U_{0}})_{\Psi}$
$=\int_{U_{1}}\!\boldsymbol{1}_{A_{1}}\cdot g\,\mathrm{d}\mathrm{P}^{(1)}_{U_{1}}\,.$ (D.3)
Applying (D.3) to $g:=\mathrm{P}^{(1)}(B_{1}|\,\cdot\,)$ shows
$\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{P}^{(1)}(B_{1}|\Psi(u_{0}))\,\mathrm{d}\mathrm{P}^{(0)}_{U_{0}}(u_{0})$ $=\int_{U_{1}}\boldsymbol{1}_{A_{1}}(u_{1})\mathrm{P}^{(1)}(B_{1}|u_{1})\,\mathrm{d}\mathrm{P}^{(1)}_{U_{1}}(u_{1})$
$=\mathrm{P}^{(1)}(B_{1}\times A_{1})\,.$ (D.4)
Moreover, for a regular conditional probability $\mathrm{P}^{(0)}(\,\cdot\,|\,\cdot\,)$ of $\mathrm{P}^{(0)}$ given $U_{0}$ our observation (D.2) gives
$\mathrm{P}^{(1)}(B_{1}\times A_{1})$ $=\mathrm{P}^{(0)}\bigl{(}\Phi^{-1}(B_{1})\times A_{0}\bigr{)}$
$=\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{P}^{(0)}(\Phi^{-1}(B_{1})|u_{0})\,\mathrm{d}\mathrm{P}^{(0)}_{U_{0}}(u_{0})\,.$ (D.5)
To motivate the next steps we note that combining (D.4) and (D.5) gives the $\mathrm{P}^{(0)}_{U_{0}}$-almost sure identity
$\mathrm{P}^{(0)}(\Phi^{-1}(B_{1})|\,\cdot\,)=\mathrm{P}^{(1)}(B_{1}|\Psi(\,\cdot\,))\,.$
However, the occurring $\mathrm{P}^{(0)}_{U_{0}}$-zero set depends on $B_{1}$, and hence we need to work harder to establish the assertion.
Let us therefore define the probability measure $\mathrm{Q}:=\mathrm{P}^{(0)}_{(\Phi,\operatorname{id}_{U_{0}})}$ on $\mathfrak{B}_{1}\otimes{\cal A}_{0}$. For $B_{1}\in\mathfrak{B}_{1}$, $A_{0}\in{\cal A}_{0}$, and $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$ the calculation (D.2) then shows
$\mathrm{Q}(B_{1}\times A_{0})=\mathrm{P}^{(0)}\bigl{(}\Phi^{-1}(B_{1})\times A_{0}\bigr{)}=\mathrm{P}^{(1)}(B_{1}\times A_{1})\,.$ (D.6)
Moreover, we have $\mathrm{Q}_{U_{0}}(A_{0})=\mathrm{Q}(T_{1}\times A_{0})=\mathrm{P}^{(0)}(T_{0}\times A_{0})=\mathrm{P}^{(0)}_{U_{0}}(A_{0})$, that is $\mathrm{Q}_{U_{0}}=\mathrm{P}^{(0)}_{U_{0}}$.
With these preparations let us first consider the map $\mathrm{Q}_{1}:\mathfrak{B}_{1}\times U_{0}\to[0,1]$ given by
$\mathrm{Q}_{1}(B_{1},u_{0}):=\mathrm{P}^{(1)}(B_{1}|\Psi(u_{0}))\,.$
Obviously, $\mathrm{Q}_{1}(B_{1},\,\cdot\,):U_{0}\to[0,1]$ is ${\cal A}_{0}$-measurable for all $B_{1}\in\mathfrak{B}_{1}$ and $\mathrm{Q}_{1}(\,\cdot\,,u_{0}):\mathfrak{B}_{1}\to[0,1]$ is a probability measure on $\mathfrak{B}_{1}$ for all $u_{0}\in U_{0}$. Let us fix some $B_{1}\in\mathfrak{B}_{1}$ and $A_{0}\in{\cal A}_{0}$. Since we assumed
${\cal A}_{0}=\sigma(\Psi)=\{\Psi^{-1}(A_{1}):A_{1}\in{\cal A}_{1}\}\,,$
where the latter identity is the very definition of $\sigma(\Psi)$, there then exists an $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$. Combining the observed $\mathrm{Q}_{U_{0}}=\mathrm{P}^{(0)}_{U_{0}}$ with (D.4) and (D.6) then yields
$\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{Q}_{1}(B_{1},u_{0})\,\mathrm{d}\mathrm{Q}_{U_{0}}(u_{0})$ $=\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{P}^{(1)}(B_{1}|\Psi(u_{0}))\,\mathrm{d}\mathrm{P}^{(0)}_{U_{0}}(u_{0})$
$=\mathrm{P}^{(1)}(B_{1}\times A_{1})$
$=\mathrm{Q}(B_{1}\times A_{0})\,.$
In other words, $\mathrm{Q}_{1}$ is a regular conditional probability of $\mathrm{Q}$ given $U_{0}$.
In addition to $\mathrm{Q}_{1}$ we also consider the map $\mathrm{Q}_{0}:\mathfrak{B}_{1}\times U_{0}\to[0,1]$ given by
$\mathrm{Q}_{0}(B_{1},u_{0}):=\mathrm{P}^{(0)}(\Phi^{-1}(B_{1})|u_{0})\,.$
Again, $\mathrm{Q}_{0}(B_{1},\,\cdot\,):U_{0}\to[0,1]$ is ${\cal A}_{0}$-measurable for all $B_{1}\in\mathfrak{B}_{1}$ and $\mathrm{Q}_{0}(\,\cdot\,,u_{0}):\mathfrak{B}_{1}\to[0,1]$ is a probability measure on $\mathfrak{B}_{1}$ for all $u_{0}\in U_{0}$, as it is the image measure of $\mathrm{P}^{(0)}(\,\cdot\,|u_{0})$ under $\Phi$. Let us fix some $B_{1}\in\mathfrak{B}_{1}$ and $A_{0}\in{\cal A}_{0}$. Again there then exists an $A_{1}\in{\cal A}_{1}$ with $\Psi^{-1}(A_{1})=A_{0}$ and by combining $\mathrm{Q}_{U_{0}}=\mathrm{P}^{(0)}_{U_{0}}$ with (D.5) and (D.6) we find
$\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{Q}_{0}(B_{1},u_{0})\,\mathrm{d}\mathrm{Q}_{U_{0}}(u_{0})$ $=\int_{U_{0}}\boldsymbol{1}_{A_{0}}(u_{0})\mathrm{P}^{(0)}(\Phi^{-1}(B_{1})|u_{0})\,\mathrm{d}\mathrm{P}^{(0)}_{U_{0}}(u_{0})$
$=\mathrm{P}^{(1)}(B_{1}\times A_{1})$
$=\mathrm{Q}(B_{1}\times A_{0})\,.$
In other words, $\mathrm{Q}_{0}$ is also a regular conditional probability of $\mathrm{Q}$ given $U_{0}$. Now the uniqueness of regular conditional probabilities, see Theorem D.1, yields the assertion, since we have already noted $\mathrm{Q}_{U_{0}}=\mathrm{P}^{(0)}_{U_{0}}$. ∎
Our next goal is to investigate Bochner integrals with respect to regular conditional probabilities and how they relate to Banach space valued conditional expectations. Let us begin by introducing the latter. To this end, let $E$ be a separable Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra. We denote the restriction of $\mathrm{P}$ to ${\cal C}$ by $\mathrm{P}_{|{\cal C}}$ and write accordingly
[…]
With this notation, we can now define conditional expectations.
Definition D.3.
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra, and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. Then $Y\in{{\cal L}_{1}(\mathrm{P}_{|{\cal C}},E)}$ is called a conditional expectation of $X$ given ${\cal C}$, if
$\int_{\Omega}\boldsymbol{1}_{C}Y\,\mathrm{d}\mathrm{P}_{|{\cal C}}=\int_{\Omega}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}\,,\qquad\qquad C\in{\cal C}.$
One can show that for each $X\in{{\cal L}_{1}(\mathrm{P},E)}$ there exists a conditional expectation of $X$ given ${\cal C}$, see e.g. [46, Proposition 4.1 in Chapter II.4.1] or [21, Theorem 2.6.20]. Moreover, if we have two conditional expectations $Y_{1}$ and $Y_{2}$ of $X$ given ${\cal C}$, then $\mathrm{P}_{|{\cal C}}(\{Y_{1}\neq Y_{2}\})=0$, see e.g. [13, Corollary 5 in Chapter II.2] or [21, Theorem 2.6.20]. As in the case $E=\mathbb{R}$, we thus write $\mathbb{E}(X|{\cal C})$ for any version of the conditional expectation of $X$ given ${\cal C}$. Here we note that for $X_{1},X_{2}\in{{\cal L}_{1}(\mathrm{P},E)}$ with $\mathrm{P}(\{X_{1}\neq X_{2}\})$ we obviously have $\mathbb{E}(X_{1}|{\cal C})(\omega)=\mathbb{E}(X_{2}|{\cal C})(\omega)$ for $\mathrm{P}_{|{\cal C}}$-almost all $\omega\in\Omega$. Consequently we can view the conditional expectation as a map ${L_{1}(\mathrm{P},E)}\to{L_{1}(\mathrm{P}_{|{\cal C}},E)}$. As usual, in the case ${\cal C}=\sigma(Y)$, we write $\mathbb{E}(X|Y):=\mathbb{E}(X|{\cal C})$.
The next results collects some useful properties of conditional expectations, where we note that most of them are literal generalizations of the well-known case $E=\mathbb{R}$. For a proof we refer to [46, Chapter II.4.1] or [21, Chapter 2.6.d].
Theorem D.4.
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and ${\cal C}\subset{\cal A}$ be a sub-$\sigma$-algebra. Then the following statements hold true:
  1. i)
    The map $\mathbb{E}(\,\cdot\,|{\cal C}):{L_{1}(\mathrm{P},E)}\to{L_{1}(\mathrm{P}_{|{\cal C}},E)}$ is linear and bounded with $\|\mathbb{E}(\,\cdot\,|{\cal C})\|=1$ and $\mathbb{E}(X|{\cal C})=X$ for all $X\in{L_{1}(\mathrm{P}_{|{\cal C}},E)}$.
  2. ii)
    For all $1\leq p\leq\infty$ the map $\mathbb{E}(\,\cdot\,|{\cal C}):{L_{p}(\mathrm{P},E)}\to{L_{p}(\mathrm{P}_{|{\cal C}},E)}$ is well-defined, linear, and continuous with $\|\mathbb{E}(\,\cdot\,|{\cal C})\|=1$.
  3. iii)
    For all sub-$\sigma$-algebras ${\cal C}_{1}\subset{\cal C}_{2}\subset{\cal A}$ and all $X\in{{\cal L}_{1}(\mathrm{P},E)}$ we have $\mathbb{E}(\mathbb{E}(X|{\cal C}_{2})|{\cal C}_{1})=\mathbb{E}(X|{\cal C}_{1})$.
  4. iv)
    If $T:E\to F$ is bounded and linear, then $\mathbb{E}(T\circ X|{\cal C})=T\circ\mathbb{E}(X|{\cal C})$ for all $X\in{{\cal L}_{1}(\mathrm{P},E)}$.
Our next goal is to collect information about $E$-valued martingales. To this end, let $(\Omega,{\cal A},\mathrm{P})$ be a probability space and $E$ be a Banach space. Then a sequence $({\cal C}_{n})$ of sub-$\sigma$-algebras of ${\cal A}$ is said to be a filtration, if ${\cal C}_{n}\subset{\cal C}_{n+1}$ for all $n\geq 1$. In this case a sequence $(X_{n})$ of maps $X_{n}:\Omega\to E$ is called a $({\cal C}_{n})$-martingale, if for all $n\geq 1$ we have $X_{n}\in{{\cal L}_{1}(\mathrm{P}_{|{\cal C}_{n}},E)}$ and
$\mathbb{E}(X_{n+1}|{\cal C}_{n})=X_{n}\,.$
Recall that $\mathbb{R}$-valued martingales enjoy well-known convergence results. The following theorem, whose proof can be found in e.g. [46, Theorems 4.1 and 4.2 in Chapter II.4.2] or [21, Theorem 3.3.2 in combination with p. 83], collects analogous results for $E$-valued martingales.
Theorem D.5.
Let $E$ be a Banach space, $(\Omega,{\cal A},\mathrm{P})$ be a probability space, and $({\cal C}_{n})$ be a filtration and ${\cal C}_{\infty}:=\sigma({\cal C}_{n}:n\geq 1)$. Moreover, let $1\leq p<\infty$ and $X\in{{\cal L}_{p}(\mathrm{P},E)}$, and for $n\geq 1$ we define
$X_{n}:=\mathbb{E}(X|{\cal C}_{n})\qquad\qquad\mbox{ and }\qquad\qquad X_{\infty}:=\mathbb{E}(X|{\cal C}_{\infty})\,.$
Then $(X_{n})$ is a $({\cal C}_{n})$-martingale and we have both
$\|X_{n}-X_{\infty}\|_{{{\cal L}_{p}(\mathrm{P},E)}}\to 0\qquad\qquad\mbox{ and }\qquad\qquad\mathrm{P}(\lim_{n\to\infty}X_{n}\neq X_{\infty}\})=0.$
Our next result establishes a disintegration formula for Bochner integrals and regular conditional probabilities. In addition, it shows that regular conditional probabilities can be used to construct conditional expectations. In the case $E=\mathbb{R}$, these results are well-known, see e.g. [14, Theorem 10.2.5] and to some extent also [27, Theorem 8.38], where in all cases actually a more general setting that goes beyond the “product case” is treated.
Theorem D.6.
Let $(T,{\cal B})$ and $(U,{\cal A})$ be measurable spaces, $\mathrm{P}$ be a probability measure on $(T\times U,{\cal B}\otimes{\cal A})$, and $\mathrm{P}(\,\cdot\,|\,\cdot\,):{\cal B}\times U\to[0,1]$ be a regular conditional probability of $\mathrm{P}$ given $U$. Moreover, let $E$ be a Banach space and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. Then there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ such that for all $u\in U\setminus N$ we have
$X(\,\cdot\,,u)\in{{\cal L}_{1}(\mathrm{P}(\,\cdot\,|u),E)}\,.$ (D.7)
Moreover, for all $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ and (D.7), the map $Z:U\to E$ defined by
$Z(u):=\int_{T}\boldsymbol{1}_{U\setminus N}(u)X(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,,\qquad\qquad u\in U.$ (D.8)
is strongly measurable and we have both $Z\in{{\cal L}_{1}(\mathrm{P}_{U},E)}$ and
$\int_{T\times U}X\,\mathrm{d}\mathrm{P}=\int_{U}Z\,\mathrm{d}\mathrm{P}_{U}=\int_{U}\int_{T}\boldsymbol{1}_{U\setminus N}(u)X(t,u)\,\mathrm{P}(\mathrm{d}t|u)\mathrm{P}_{U}(u)\,.$ (D.9)
Finally, the function $Z\circ\pi_{U}:T\times U\to E$ is a version of $\mathbb{E}(X|{\cal C})$, where ${\cal C}:=\pi_{U}^{-1}({\cal A})$.
Proof of Theorem D.6: We first note that for $X:T\times U\to[0,\infty)$ the assertions (D.7) and (D.9) are well-known: Indeed, in this case, the measurability of $Z:U\to[0,\infty]$ defined by (D.8) with $N:=\emptyset$ can be found in e.g. [27, Lemma 14.20]. Moreover, (D.9) is shown in e.g. [27, Theorem 14.29]. For $\mathrm{P}$-integrable $X:T\times U\to[0,\infty)$, the just verified (D.9) then shows that there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ and $Z(u)\in[0,\infty)$ for all $u\in U\setminus N$, that is, we have (D.7). Finally, if we fix such an $N$ and apply the previous results to $X_{0}:=\boldsymbol{1}_{U\setminus N}X$, we again have the measurability of $Z$ and (D.9).
Let us now assume that $E$ is a general Banach space. We first consider the case in which $X\in{{\cal L}_{1}(\mathrm{P},E)}$ is a step function. To this end, we pick a representation
$X=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i}}x_{i}$
for some $x_{i}\in E$ and $D_{i}\in{\cal B}\otimes{\cal A}$. For $u\in U$ we have $D_{i,u}:=\{t\in T:(t,u)\in D_{i}\}\in{\cal B}$, see e.g. [27, Lemma 14.13], and hence we see that
$X(\,\cdot\,,u)=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i}}(\,\cdot\,,u)x_{i}=\sum_{i=1}^{m}\boldsymbol{1}_{D_{i,u}}x_{i}$
is a measurable step functions for all fixed $u\in U$. Consequently, (D.7) holds for $N:=\emptyset$. For this $N$ we define $Z$ by (D.8). For $u\in U$ the definition of Bochner integrals for step functions, see (A.3), yields
$Z(u)=\int_{T}\sum_{i=1}^{m}\boldsymbol{1}_{D_{i}}(t,u)x_{i}\,\mathrm{P}(\mathrm{d}t|u)=\int_{T}\sum_{i=1}^{m}\boldsymbol{1}_{D_{i,u}}(t)x_{i}\,\mathrm{P}(\mathrm{d}t|u)$ $=\sum_{i=1}^{m}\biggl{(}\int_{T}\boldsymbol{1}_{D_{i,u}}(t)\,\mathrm{P}(\mathrm{d}t|u)\biggr{)}\cdot x_{i}$
$=\sum_{i=1}^{m}\biggl{(}\int_{T}\boldsymbol{1}_{D_{i}}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\biggr{)}\cdot x_{i}\,.$
Let us define $f_{i}:U\to[0,\infty)$ by
$f_{i}(u):=\int_{T}\boldsymbol{1}_{D_{i}}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,.$
From our initial remarks we know that each $f_{i}$ is measurable, and hence $u\mapsto f_{i}(u)x_{i}$ is a strongly measurable, $E$-valued function. Since $Z$ is the sum of these $m$ functions, it is also strongly measurable. Moreover, the already established (D.9) for non-negative, measurable functions gives
$\int_{U}Z\,\mathrm{d}\mathrm{P}_{U}=\int_{U}\sum_{i=1}^{m}\biggl{(}\int_{T}\boldsymbol{1}_{D_{i}}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\biggr{)}\cdot x_{i}\,\mathrm{d}\mathrm{P}_{U}(u)$ $=\sum_{i=1}^{m}\biggl{(}\int_{U}\int_{T}\boldsymbol{1}_{D_{i}}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,\mathrm{d}\mathrm{P}_{U}(u)\biggr{)}\cdot x_{i}$
$=\sum_{i=1}^{m}\biggl{(}\int_{T\times U}\boldsymbol{1}_{D_{i}}\,\mathrm{d}\mathrm{P}\biggr{)}\cdot x_{i}$
$=\int_{T\times U}X\,\mathrm{d}\mathrm{P}\,.$
In other words, (D.9) holds for measurable $E$-valued step functions and $N=\emptyset$.
Let us now consider the case of general Banach spaces $E$ and $X\in{{\cal L}_{1}(\mathrm{P},E)}$. To prove (D.7), we write $Y:=\|X(\,\cdot\,,\,\cdot\,)\|_{E}:T\times U\to[0,\infty)$. Since $X$ is strongly measurable, $Y$ is measurable. Applying Equation (D.9) in the already established case of $E=\mathbb{R}$ with non-negative functions to $Y$ we then obtain
$\int_{U}\int_{T}\|X(t,u)\|_{E}\,\mathrm{P}(\mathrm{d}t|u)\mathrm{P}_{U}(u)=\int_{T\times U}\|X(\,\cdot\,,\,\cdot\,)\|_{E}\,\mathrm{d}\mathrm{P}<\infty\,.$ (D.10)
Consequently, there exists an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ such that for all $u\in U\setminus N$ we have
$\int_{T}\|X(t,u)\|_{E}\,\mathrm{P}(\mathrm{d}t|u)<\infty.$
This shows (D.7).
For the following steps we fix an $N\in{\cal A}$ with $\mathrm{P}_{U}(N)=0$ such that (D.7) holds. We write $X_{0}:=\boldsymbol{1}_{U\setminus N}X$ and define $g:T\times U\to[0,\infty)$ by $g(t,u):=\|X_{0}(t,u)\|_{E}$ for all $t\in T$ and $u\in U$. By our assumption, $g:T\times U\to[0,\infty)$ is $\mathrm{P}$-integrable.
Let us first verify that $Z$ is strongly measurable. Since $X$ is strongly measurable, so is $X_{0}$, and therefore there exists a sequence $(X_{n})$ of measurable step functions $X_{n}:T\times U\to E$ approximating $X_{0}$ in the sense of (A.1) and (A.2). We define $Z_{n}:U\to E$ in the sense of (D.8), that is
$Z_{n}(u):=\int_{T}X_{n}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,,\qquad\qquad u\in U.$ (D.11)
Since $X_{n}$ is a measurable step function we already know that this definition is possible and that $Z_{n}$ is strongly measurable. Moreover, for each fixed $u$, the sequence $(X_{n}(\,\cdot\,,u))$ consists of measurable step functions approximating $X_{0}(\,\cdot\,,u)$ in the sense of (A.1) and (A.2). By the definition of the Bochner integral (A.4) we thus find
$Z_{n}(u)=\int_{T}X_{n}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\to\int_{T}X_{0}(t,u)\,\mathrm{P}(\mathrm{d}t|u)=Z(u)\,,\qquad\qquad u\in U.$ (D.12)
In other words, $Z$ is the pointwise limit of a sequence of strongly measurable functions, and thus also strongly measurable.
To verify that $Z\in{{\cal L}_{1}(\mathrm{P}_{U},E)}$, we simply note that the norm inequality for Bochner integrals (A.6) together with (D.10) yields
$\int_{U}\|Z\|_{E}\,\mathrm{d}\mathrm{P}_{U}$ $=\int_{U}\Bigl{\|}\,\int_{T}\boldsymbol{1}_{U\setminus N}(u)X(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,\Bigr{\|}_{E}\,\mathrm{d}\mathrm{P}_{U}(u)$
$\leq\int_{U}\int_{T}\|X(t,u)\|_{E}\,\mathrm{P}(\mathrm{d}t|u)\,\mathrm{d}\mathrm{P}_{U}(u)$
[…]
To establish (D.9) for general $X\in{{\cal L}_{1}(\mathrm{P},E)}$ we again fix a sequence $(X_{n})$ of measurable step functions $X_{n}:T\times U\to E$ approximating $X_{0}$ in the sense of (A.1) and (A.2). For the strongly measurable functions $Z_{n}:U\to E$ defined by (D.11) we have already established (D.12). By (A.2) we further find
$\|Z_{n}(u)\|_{E}=\Bigl{\|}\,\int_{T}X_{n}(t,u)\,\mathrm{P}(\mathrm{d}t|u)\,\Bigr{\|}_{E}\leq\int_{T}\|X_{n}(t,u)\|_{E}\,\mathrm{P}(\mathrm{d}t|u)\leq\int_{T}g(t,u)\,\mathrm{P}(\mathrm{d}t|u):=h(u)$
for all $u\in U$. Since $g$ is $\mathrm{P}$-integrable, the initially considered case of non-negative, measurable functions shows that $h:U\to[0,\infty]$ is $\mathrm{P}_{U}$-integrable. In summary, the sequence $(Z_{n})$ together with $Z$ and $h$ satisfy (A.7) and (A.8), and hence an application of the dominated convergence theorem for Bochner integrals (A.9) shows
$\int_{U}Z_{n}\,\mathrm{d}\mathrm{P}_{U}\to\int_{U}Z\,\mathrm{d}\mathrm{P}_{U}\,.$ (D.13)
Since the sequence $(X_{n})$ of measurable step functions $X_{n}$ approximates $X_{0}$ in the sense of (A.1) and (A.2), the definition of the Bochner integral further shows
$\int_{T\times U}X_{n}\,\mathrm{d}\mathrm{P}\to\int_{T\times U}X_{0}\,\mathrm{d}\mathrm{P}$ (D.14)
Since we already know that (D.9) holds for all pairs of $X_{n}$ and $Z_{n}$, that is
$\int_{U}Z_{n}\,\mathrm{d}\mathrm{P}_{U}=\int_{T\times U}X_{n}\,\mathrm{d}\mathrm{P}\,,$
the just established (D.13) and (D.14) then yield (D.9) in the general case.
Finally, we need to verify that $Z\circ\pi_{U}$ is a version of $\mathbb{E}(X|{\cal C})$. Here we first note that $Z\circ\pi_{U}$ is ${\cal C}$-measurable by construction. Let us now pick a $C\in{\cal C}$. Then there exists an $A\in{\cal A}$ with $C=T\times A$. Now, the function $X_{A}:=\boldsymbol{1}_{A}X_{0}$ satisfies $X_{A}\in{{\cal L}_{1}(\mathrm{P},E)}$ and by construction we have (D.7) for $N_{A}:=\emptyset$. In addition, we find
$Z_{A}(u):=\int_{T}X_{A}(t,u)\,\mathrm{P}(\mathrm{d}t|u)=\boldsymbol{1}_{A}(u)\int_{T}X_{0}(t,u)\,\mathrm{P}(\mathrm{d}t|u)=\boldsymbol{1}_{A}(u)Z(u)$
for all $u\in U$. From this we can conclude $Z_{A}\circ\pi_{U}=1_{C}\cdot Z\circ\pi_{U}$, and applying (D.9) to $X_{A}$ thus yields
$\int_{T\times U}\boldsymbol{1}_{C}X\,\mathrm{d}\mathrm{P}=\int_{T\times U}X_{A}\,\mathrm{d}\mathrm{P}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{U}=\int_{U}Z_{A}\,\mathrm{d}\mathrm{P}_{\pi_{U}}=\int_{T\times U}Z_{A}\circ\pi_{U}\,\mathrm{d}\mathrm{P}=\int_{T\times U}\boldsymbol{1}_{C}Z\circ\pi_{U}\,\mathrm{d}\mathrm{P}\,.$
Consequently, $Z\circ\pi_{U}$ is a version of $\mathbb{E}(X|{\cal C})$. ∎